A Complete Bibliography of Publications in AStA. Advances in Statistical Analysis

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

13 March 2024 Version 1.31

Title word cross-reference

[0, ∞)^{k} [488]. ℓ_{∞} [528]. Γ [28]. h [314]. J [3]. k [233, 362, 314]. L [54]. L^{p} [15]. L_{1} [11]. M [132, 605]. Z [369]. NMCpm [431]. n [371]. P [291]. s^{n−m} [390]. t [55, 598, 412].

-depth [15]. -distribution [598]. -estimators [132]. -factor [412]. -group [233]. -median [15]. -quantile [605]. -spline [291]. -Statistics [54]. -transformation [3].

/Book [44, 52, 60, 68]. /Books [8, 24].

19 [574, 600, 575, 552].

2002 [110]. 2005 [104]. 2nd [176].

3rd [207].

8 [143].

95 [156].

abruptly [576]. Absenteeism [595]. Absolute [270]. absolutely [442]. accelerated [373]. Access [47, 49, 193]. account [284]. Accumulation [31]. accuracy [500, 435]. across [506]. Action [567]. actions [567]. actuarial [299, 296, 295]. Adaptive [290]. Additive [291, 508, 399, 441]. Addressing [598]. adjusted [132, 478, 67]. administered [38]. Administrative [18, 130]. Advances [239, 446, 82]. advantage [574]. advice [48, 149]. affair [150]. after [420]. Agency [49]. agreement [223]. agricultural [563, 557]. aircraft [260]. al. [547]. Algebra [183]. algorithm [449, 170, 165, 587, 455]. algorithms [258, 282, 205]. All-pairs [397]. allocation [475, 307]. ALT [246]. älteren [32]. alternating [439]. alternative [380, 310, 246, 244]. Alternatives [207]. am [49, 63]. American [170]. Ämter [59]. amtlichen [32, 108]. analyses [285]. analysing [136]. Analysis [357, 390, 94, 140, 57, 129, 64, 185, 240, 426, 87, 506, 202, 589, 519, 395, 323, 299, 77, 451, 500, 279, 76, 199, 253, 532, 75, 335, 435, 598, 388, 181, 333, 218, 267, 176, 37, 347, 501, 483, 203, 251, 259, 498, 289, 222]. analytics [203]. analyzers [412]. analyzing [485, 340, 415]. animal [421]. Anisotropy [218]. anomie [451]. anonymised [192]. Anonymisierung [5]. anonymization [5]. anonymized [80]. ANOVA [514, 258]. Antwort [23]. Anwendung [115]. any [541]. Application [58, 130, 287, 520, 596, 603, 372, 162, 572, 305, 330, 517, 466, 224, 97, 229, 370, 601, 495]. Applications [185, 183, 430, 195, 528, 306, 417, 419, 354, 362, 379, 309, 569, 176, 444, 206, 203]. applied [105]. approach [292, 476, 481, 485, 487, 42, 131, 284, 287, 321, 322, 26, 193, 180, 486, 563, 473, 296, 137, 604, 272, 516, 507, 570, 222, 557]. Approaches [333, 387]. appropriateness [88]. Approximate [466, 269]. approximation [275]. approximations [21]. Arbeit [49]. Arbeitsmarkt [49, 46]. Arbeitsmarkt- [49]. Arbeitsmarktforschung [107, 46, 48]. Arbeitsmarktstatistik [48]. arbitrary [470]. Archimedean [489, 396]. Archipelago [327]. Ardilly [114, 125]. area [130, 544]. ARIMA [87]. ARIMA-methods [87]. ARL [359]. ASN [220]. aspects [581, 279, 22]. Aspekte [22].

Assessing [141, 284, 181, 321, 577, 106].

Assessment

[426, 563, 581, 500, 269, 387, 557]. asset [351]. Assisted [199]. association [223, 474, 138]. association/correlation [474]. assumptions [361]. assurance [191, 154]. Asymmetric [120, 521, 250]. Asymptotic [367, 235, 271, 349, 353, 386, 275]. asymptotics [520, 384]. athletes [581]. attrition [181]. außeruniversitären [111]. Author [551]. autodependogram [342]. automatic [338]. automobile [63]. autoregression [492]. autoregressions [269]. autoregressions-Monte [269]. Autoregressive [74, 406, 440, 75, 398, 526, 525]. auxiliary [446]. availability [49]. average [16]. axial [226].

B. [184]. Bachelor [115]. Bachelor- [115]. back [502]. back-door [502]. backward [358]. Baden [169]. ball [578]. bands [481, 478, 312]. Bandwidth [338]. based [292, 454, 42, 380, 556, 519, 164, 210, 101, 581, 568, 400, 451, 602, 522, 228, 606, 414, 486, 387, 170, 373, 122, 502, 467, 16, 563, 557, 527, 365, 438, 597, 577, 4, 397, 566, 333, 389, 300, 396, 277, 545, 531, 259, 587, 455, 524, 564, 317]. basic [537]. Basis [67, 1, 438]. basketball [579, 577]. Bates [584]. Bayes [573, 516]. Bayesian [202, 541, 426, 284, 437, 394, 572, 354, 373, 466, 597, 348, 296, 336, 501, 324, 540, 524, 498, 222]. Befragung [7]. behaviour [136]. behind [509]. Beispiel [63]. benefits [51, 105]. Beratung [48]. Berufsforschung [49]. Beta [66, 538, 365, 459, 503]. Betrieben [62]. Betriebswirte [6, 23]. better [193]. between [515, 62, 425, 4, 459, 150]. Bias [164, 240, 141, 241, 242, 577]. bimodality [162]. Binary [595, 391, 222]. binomial [542, 511, 492, 398, 142]. binormal [500]. biodiversity [423]. Biomarker [500]. biomechanics [582]. bivariate [454, 574, 486, 270, 435, 229, 63]. Blinder [172]. body [12, 394]. Book [44, 52, 60, 68, 202, 116, 206, 204, 156, 166, 176, 185, 115, 114, 125, 157, 142, 133, 184, 183, 207, 203, 205, 113, 134, 143]. Books [8, 24]. Boosting [293]. bootstrap [554, 478]. Bootstrapping [123, 585]. bounded [530, 542]. box [556, 307, 417]. box-counting [556]. branching [38]. Brazilian [517]. break [311]. Brownian [119]. Buchbesprechung [112]. Buchbesprechungen [44, 52, 60, 68, 92, 102]. Bundesagentur [49]. Bundesliga [573, 370]. Business [50, 19, 6, 23]. BWL [115].

C [115, 157]. C. [207]. calculating [28]. Calculation [200]. Calculator [576]. calibration [453]. CALT [246]. CAMCR [199]. capability [431, 267]. Capture [199, 195, 424, 197, 196]. capture-recapture [195, 197, 196]. Care [595]. Carlo [269, 170, 57, 165]. cars [123]. cartographic [130]. case [262, 602, 575, 310, 577, 190, 359]. categorical [292, 163]. Cauchy [512]. causal [554, 600]. cell [405]. censoring [461]. census [101, 18]. centers [59]. Centre [49]. chain [165]. challenge [474]. challenges [46, 421, 108]. Chancen [46]. Change [329, 462, 574, 286, 245]. Change-in-mean [462]. changes [401, 359]. characteristic [86, 467]. Characteristics [37]. characterization [442]. characterizations [606]. chart [219]. charts [42, 406, 594, 16, 359]. checking [543]. choice [136, 396]. Christian [207]. CIR [584]. claim [354]. claims [601]. Class [54, 542]. classes [433]. classification [298, 1, 527, 509, 531]. classifier [531]. Clémentine [206]. climate [549]. cliques [348]. Closure [433]. clustered [511]. Clustering [603, 590, 381, 545, 412]. clusters [449]. clusterwise [272]. Cochran [314]. Codependent [332]. coefficient [409, 383, 450]. coefficients [440, 525]. cogent [553]. cognition [285]. cognitive [340]. Coherent [376]. cointegrated [75]. Cointegration [129, 56, 251, 74]. collaboration [536]. collective [301]. collinearity [535]. colored [348]. combination [191]. Combinations [55]. Combining [106]. Comment [547, 549]. common [27, 405, 412]. communicate [29]. companies [62]. comparative [555]. Comparing [257, 12, 506, 517]. Comparison [174, 110, 211, 463, 127, 445, 380, 210, 90, 21, 469, 370, 537, 444, 464]. comparisons [591, 372, 397, 447]. competitions [575]. completely [132]. completion [590]. complex [443, 26, 262, 343, 333, 446]. Component [581, 589, 162, 267, 347, 289]. Component-based [581]. components [222]. composite [435, 159]. comprehension [38]. computational [105]. Computations [183]. Computer [254, 253, 199]. Computer-Assisted [199]. concave [434]. concentration [322]. concepts [298]. Conceptual [31, 105]. Conditional [351, 353, 367, 286, 377, 437, 366, 27]. Confidence [519, 481, 478, 312, 389, 277, 416]. confounding [390, 498]. connectivity [287]. considerations [359]. Consistency [132, 353, 428]. consistent [268, 607]. constant [415]. constraint [173, 214]. constraints [307, 255]. construct [481]. consulting [236, 145, 146, 154, 148]. consumer [136]. Consumption [64]. Contacts [9, 17, 25, 34, 45, 53, 61, 69, 85, 93, 103, 117, 126, 135, 144, 158, 167, 177, 186, 194, 208, 216, 225, 230, 238, 247, 252, 261, 266, 273, 278, 288, 294, 302, 308, 313, 319, 328, 334, 339, 345, 350, 355, 360]. contamination [327]. content [408].

Contextual [568]. contingency [405]. continue [420]. Continuous [579, 282, 479, 442, 233, 488, 470, 270, 584]. Continuous-discrete [282]. Continuous-time [579]. contribution [284, 152]. contributions [100]. Control [406, 594, 42, 193, 16, 464, 463, 187]. controlled [310]. conversion [573]. convex [12]. Conway [357]. copula [485, 484, 597, 566, 374]. copula-based [597, 566]. copulas [515, 303, 396]. copulis [231]. corporate [59]. Correction [599, 557, 464, 164]. correlated [348, 190, 203]. Correlation [606, 515, 351, 474, 14]. Correlation-type [606]. correlations [169]. correspondence [519]. Count [506, 504, 490, 392, 449, 199, 473, 555, 222, 571]. counting [556]. countries [563, 557]. counts [542, 264, 405, 182]. coupled [259]. course [372]. covariance [413, 528, 268, 352, 229]. covariate [458]. covariates [439, 452, 348, 378]. coverage [141, 482]. COVID [600, 575, 552, 574]. COVID-19 [600, 575, 552, 574]. Cox [562, 599, 417, 310, 570]. Cramér [448]. Creation [50]. credibility [354]. criterion [256, 502]. Critical [314]. cross [287, 271]. cross-dependent [271]. cross-sectional [287]. Cucconi [397]. Cumulants [371]. cumulative [519, 486, 138, 495]. Curricular [22]. Curriculare [22]. curve [286, 500]. curves [509, 318]. CUSUM [219]. Cyber [522]. Cycles [50]. cyclical [534].

Danksagung [33]. dare [549]. Data [402, 58, 13, 57, 175, 107, 104, 14, 479, 504, 223, 481, 430, 461, 586, 130, 353, 485, 512, 280, 141, 536, 123, 454, 548, 192, 326, 506, 131, 380, 538, 392, 535, 449, 26, 49, 562, 599, 193, 109, 120, 391, 417, 198, 226, 427, 59, 19, 168, 357, 593, 279, 76, 269, 180, 547, 78, 551, 546, 511, 242, 281, 199, 16, 407, 376, 94, 140, 473, 31, 597, 10, 97, 552, 309, 229, 582, 432, 333, 108, 549, 474, 580, 36, 405, 381, 80, 5, 105, 166, 37, 282, 483, 555, 370, 203]. data [316, 601, 450, 455, 505, 507, 498, 196, 222, 571, 49, 134]. datasets [560]. date [91]. Daten [108]. Datenverfügbarkeit [49]. Datenzugang [49]. David [176, 156]. DDMA [16]. DDMA-charts [16]. deal [555]. dealing [588]. deals [90]. decision [538, 509]. decisions [548, 547, 551, 546, 549]. decomposition [172, 62, 179]. Decompositions [569]. Dedecker [206]. defence [410]. definite [528]. Definition [96]. deformation [496]. Dekomposition [62]. Delampady [202]. delay [552]. delta [309]. delta-sequences [309]. dementia [340]. Denmark [50]. Dennis [204]. Density [540, 353, 227, 414, 338, 155, 604]. dependence [287, 303, 372, 206, 517, 163]. dependencies [342]. dependent [235, 271, 585, 428]. depth [513, 530, 539, 16, 14, 178, 15]. depths [13, 175]. derivation [354]. described [393]. Describing [550]. Design [253, 443, 448, 372]. designs [28, 256, 390, 257, 447, 523]. desired [515]. Destruction [50]. detect [240, 449, 242]. Detecting [342, 243, 263]. detection [420, 241, 329, 524]. deterministic [296]. Deutschland [110, 30]. Developing [1, 156]. development [153]. Developments [70, 195, 109, 462]. deviation [86, 220]. devices [174]. Diagnostic [543, 500, 435]. diagnostics [151, 375]. Diagonal [138]. dialogue [425]. Dieter [185]. difference [411, 140]. differences [62]. different [304, 90, 387, 127]. differential [393, 246]. differentials [537]. dimension [541, 556, 404]. Dimensional [64, 26]. dimensions [476, 70]. Directional [486, 512]. Dirichlet [291]. disagreement [223]. disclosure [193, 191, 19, 187]. discovering [550]. Discrete [384, 134, 276, 357, 170, 376, 211, 282]. Discriminant [388]. Discussion [548, 421, 23]. Discussions [6]. disjoint [589]. Diskussion [23]. Diskussionsbeiträge [6]. disparities [106]. dispersion [490, 384, 260]. disputandum [231]. dissemination [32]. Distance [365, 424, 448, 200, 438, 459, 518]. Distance-based [365, 438]. distributed [74, 607, 371]. Distribution [249, 512, 306, 357, 561, 162, 343, 330, 486, 470, 362, 270, 453, 155, 201, 213, 274, 598, 275, 495]. distribution-free [343]. Distributional [584, 431]. distributions [442, 417, 400, 226, 382, 521, 488, 361, 459, 545, 508, 416]. distributors [517]. disturbances [280]. Divergence [129, 411, 443, 386, 604]. diversification [59]. do [51]. domains [544, 475]. domains-efficient [475]. dominance [586, 122]. door [502]. dose [444]. dose-response [444]. Double [220, 191, 86]. doubly [505]. Doukhan [206]. driven [379, 525]. drought [97]. due [141, 243]. duration [51, 77, 89, 181, 316]. during [574]. Dynamic [72, 227, 536, 600, 393, 285, 166]. dynamics [120, 229].

E-learning [90, 90]. earnings [232]. East [110, 62]. Eastern [89]. Ecological [418, 387]. ecology [419, 425]. Econometric [76, 57]. econometrics [7, 70]. Economic [51, 48, 6, 22, 5, 149, 23, 141]. economics [7]. economists [6, 23]. edition [207]. editor [418]. Editorial [565, 583, 217, 477, 35, 118]. edn. [176]. education [6, 23]. effect [554, 321, 497, 592, 434, 399, 27, 604, 263, 41]. Effects [56, 426, 83, 434, 89, 511, 432]. Efficiency [364]. Efficient [305, 281, 495, 221, 475, 258]. Einkommenslagen [110]. Einsatz [107]. Einzeldaten [5]. elastic [535, 587]. election [482]. electricity [305, 330]. element [259]. elementary [429]. Elliott [142]. elliptical [234]. Emergence [512]. Empirical [411, 383, 505, 499, 46, 19, 387, 467, 407, 335, 516, 601, 564]. empirische [46]. Employment [49, 51]. ending [575]. Endogenous [595, 452, 378]. Energy [64, 520, 517]. Engineers [157]. engines [540]. enhancements [219]. enmity [286]. enrollment [87]. entitlement [51]. environmental [320]. equation [239, 581, 242, 502, 517, 246, 244, 41]. equations [393, 358]. equidispersed [555]. equity [161]. equivalence [518]. Erfahrungen [48]. Ergebnisse [7, 5]. Eric [116]. erroneous [108]. Error [104, 383, 594, 140, 82, 21, 374, 457, 458, 523]. errors [499, 431, 235, 394, 190, 267, 428, 349]. errors-in-variables [499, 394, 349]. Erste [5]. est [231]. establishment [193, 262]. ESTAR [99]. estate [30]. estimate [367]. estimated [221, 502, 21]. Estimates [405, 366]. Estimating [574, 439, 595, 228, 232, 380, 480, 553]. Estimation [223, 192, 520, 325, 575, 387, 436, 453, 98, 188, 472, 461, 554, 493, 363, 422, 426, 287, 544, 83, 71, 290, 491, 341, 299, 578, 391, 603, 198, 191, 393, 227, 168, 420, 414, 338, 215, 301, 305, 79, 475, 532, 577, 249, 489, 127, 309, 229, 190, 446, 604, 318, 378, 41, 540, 316, 458, 495]. estimations [445]. estimator [353, 268, 432]. Estimators [197, 132, 443, 210, 235, 364, 160, 428, 349]. EU [162]. Europe [451]. European [557, 563]. EV [235, 428]. evaluating [578]. Evaluation [310, 141, 4, 580, 149, 317]. event [333, 570]. events [439, 310]. Evidence [57, 129, 269, 331, 184]. EWMA [359]. Exact [361, 344, 518, 444]. Examining [161]. example [63]. examples [206, 529, 176]. exceedances [4]. excited [492]. excitement [578]. exciting [398]. Exercises [114, 125]. existence [160]. Expectation [455]. expectile [312].

Experience [47, 123]. experiences [48]. experimental [259, 331]. experiments [253, 254]. Explanatory [595]. exponential [504, 270, 274]. exponentiated [274]. Extended [370, 306, 412]. extension [172, 519]. extensions [488]. external [559, 502, 516]. extra [542]. extra-binomial [542]. Extrapolation [189]. extremal [508]. Extreme [299, 97, 185, 326, 385, 603, 558, 361, 495].

fachsystematische [22]. Factor [64, 240, 72, 284, 340, 227, 137, 412]. Factors [38, 577]. factual [5]. faculties [7]. failure [382, 373]. fair [576]. faked [37]. fakes [37]. faktischer [5]. Fakultäten [7]. family [607, 374]. family-wise [374]. Fashion [596]. fear [150]. feature [553]. Federal [49]. federalism [91]. Fehlende [108]. fehlerhafte [108]. fiducial [410]. fields [524, 185]. Fieller [277]. Figures [150]. filter [534, 282]. Filtered [57]. filtering [171]. Final [576, 575]. Finance [185, 142]. Financial [116]. Finanzmarktstatistik [112]. findings [19]. finite [292, 453, 275, 259]. First [101, 440, 420, 526, 371, 5]. First-order [440, 526]. fit [510, 606, 330, 243]. fitting [533]. fixed [448, 434, 432, 604]. Flexible [555, 412]. flow [200]. Flows [50, 568]. focus [111, 111]. focused [286]. Föderalismus [91]. Fokus [111, 111]. Football [580, 576, 575, 572, 566, 569, 370, 570]. Forecasting [180, 315, 376, 552]. forecasts [87]. form [332]. formal [429]. forming [259]. formulas [502]. formulation [198]. Forrest [166]. Forschungsdatenzentren [59]. Forschungsdatenzentrum [49]. Fortran [156]. foundations [31]. four [286, 577]. Fourier [415]. fractional [164, 399]. fractions [100]. fragmented [228]. frailty [454]. framework [429, 500, 553, 99]. free [343]. frequencies [519]. frequency [76, 180]. Friedrich [112].

Friendly [166]. frontier [221]. frontiers [232]. fully [338]. function [481, 290, 377, 168, 382, 486, 467, 453, 468, 503, 178, 336, 469, 495]. functional [353, 287, 367, 427, 309, 582, 381, 347, 585, 483, 472, 259, 523, 222]. functions [380, 439, 478, 312, 511, 11, 438]. funds [365]. Further [58]. fusing [506]. future [354, 111]. futures [379]. fuzzy [538].

G [134]. G. [134]. Gabriel [206]. game [578, 577]. Gamma [66]. GARCH [542, 351, 375]. Gauss [171]. Gaussian [485, 562, 599, 341, 561, 388, 524]. Gaussians [373]. GDP [180]. GEE [229]. gemischte [63]. general [510, 395, 191, 269, 390]. generalised [571]. Generalizations [207]. Generalized [171, 442, 449, 270, 441, 2]. generating [257]. generation [402]. generator [489]. genetics [419]. Gentle [183]. genuine [37]. Geoadditive [326]. Geometric [479]. geometry [11, 10]. geostatistical [323, 327, 496]. German [130, 141, 51, 110, 46, 49, 262, 101, 65, 48, 7, 62, 59, 67, 6, 232, 90, 30, 22, 32, 31, 18, 108, 107, 5, 111, 100, 370, 148, 63, 23, 91]. Germany [110, 286, 1, 89, 30, 229]. Ghosh [202]. Gini [363]. given [228]. GLMs [533]. global [413, 480, 453]. GMM [380, 532]. Goodness [510, 330, 606]. Goodness-of-fit [330, 606]. Grammar [133]. graphical [388]. Graphics [133, 166]. graphs [447]. great [368]. Group [587, 239, 286, 414, 233, 563, 557]. group-based [563, 557]. group-focused [286]. grouped [131, 474]. grouping [321]. Growth [283, 286, 100]. Grubbs [314]. Grundlagen [31]. Guest [418].

H [176]. hand [579]. handle [576]. harmonically [534]. Hausman [119]. Having [578]. hazard [439, 316]. heading [537]. health [297, 285, 549, 106, 601]. health-dare [549]. heavy [521, 545]. heavy-tailed [545]. Hedonic [596, 20, 123, 429, 287, 605]. height [394]. Helge [207]. Help [9, 17, 25, 34, 45, 53, 61, 69, 85, 93, 103, 117, 126, 135, 144, 158, 167, 177, 186, 194, 208, 216, 225, 230, 238, 247, 252, 261, 266, 273, 278, 288, 294, 302, 308, 313, 319, 328, 334, 339, 345, 350, 355, 360]. Herausforderungen [46, 108]. here [529]. Hermite [171]. heterogeneity [284, 340, 501, 161]. heterogeneous [430]. Heteroscedastic [521, 409, 349]. heteroscedasticity [514]. heteroskedasticity [284, 268, 27]. heteroskedasticity-consistent [268]. Heumann [207]. hidden [276, 566]. Hierarchical [589, 573, 590, 321, 572]. High [64, 26, 76]. high-dimensional [26]. history [333]. Hoek [142]. home [574]. homogeneous [496, 405]. homoscedastic [523]. hot [579]. housing [287, 20]. HP [57]. Hydrology [185]. Hypercube [257, 256, 255]. hypotheses [280, 265]. hypothesis [518]. hysteretic [492].

i.i.d [464, 463]. IAB [51, 193, 107]. identification [452, 18, 258]. identifying [436]. Ignoring [56]. Illuminate [317]. illustration [241]. im [49, 115]. imaging [466]. immigration [1]. Impact [431, 595, 50, 40, 577, 181, 37]. imperfect [420, 494]. implementable [514]. implementation [31]. implications [489]. implied [161]. Importance [358, 147]. Improved [503]. Improving [554, 425]. impulse [478, 436]. Imputation [564, 193, 262, 40, 593, 165, 543, 317]. Imputation-based [564]. impute [281]. imputed [289]. in-match [578]. inappropriate [88]. INAR [369]. Incidence [40, 229, 324]. income [110, 40, 39, 162, 343]. incomplete [481, 487, 281]. independence [606]. independent [249]. index [353, 363, 445, 123, 431, 591, 500, 569, 537, 458, 65]. indexes [488]. indicators [544, 572, 30]. indices [429, 20, 224]. indirect [290]. Individual [50, 421, 98, 188, 501]. Inducing [515]. Industry [596]. inefficiency [139]. inequality [40, 255]. infections [552]. Inference [292, 221, 560, 410, 437, 400, 393, 343, 387, 386, 582, 588, 540, 505]. Inferences [277, 503, 564]. Inferential [467, 105]. infinite [460]. infinite-order [460]. inflated [487, 362]. Inflation [129]. Influence [459, 375, 395, 532, 38]. Information [408, 506, 106, 446]. informative [461]. infrared [260]. infrequent [263]. initial [181]. injuries [570]. innovations [526]. innovative [511]. input [590]. input-output [590]. inspection [470]. Institut [49]. Institute [49]. institutes [111]. instructions [38]. Insurance [185, 298, 297, 26, 601, 63]. integer [542, 248, 440, 526, 525, 375]. integer-valued [542, 440, 526, 525, 375]. integers [371]. Integrated [198, 422, 539, 514, 467]. Integration [577, 164, 271]. intensity [351, 168]. interactions [553, 146]. interactive [166]. Interdisciplinary [147, 279, 425]. internal [516]. interpolation [130]. Interpretation [115]. interval [461, 402, 277, 289]. intervals [478, 389, 277]. intervention [502]. interventions [506]. interviews [37]. Introducing [571]. Introduction [279, 243, 10, 36, 202, 418, 113]. investigation [169]. investment [304]. IPPS [95]. IRT [451]. issue [323, 295, 565, 279, 418, 327]. Italian [596]. Italy [327, 600, 569]. Item [40, 36, 84, 446, 41]. items [5]. iterative [587].

J [142]. J. [183]. Jahn [547]. Jayanta [202]. Jérôme [206]. Jiahui [116]. Job [50]. joint [471, 464, 463, 571]. José [206]. Joseph [156]. journalist [150]. jump [228]. jumps [263, 275].

K. [113]. kernel [367, 414, 338, 518]. kernels [356]. key [421]. Kingdom [109]. knowledge [516]. KOALA [482]. Kolmogorov [358]. Konzeptionelle [31]. Kraftfahrthaftpflichtversicherung [63]. Kriging [257]. Kullback [256]. Kumulation [31]. künftigen [111]. Kurtosis [3, 560, 96].

L [133]. Labor [47, 46, 48, 1, 107, 48]. laboratory [151]. lag [74]. landscape [550]. Lang [206]. language [147]. Laplace [224]. large [476, 560, 71, 269, 466, 165, 277]. large-scale [466]. largest [201]. Lasso [602, 449, 571]. Lasso-based [602]. LASSO-type [571]. latent [430, 240, 286, 137, 222]. Latin [256, 255, 257]. lattice [341]. League [580]. learning [147, 90, 137, 90]. learning-software [90]. Least [207, 244, 272]. least-squares [272]. Lecture [206]. Left [601, 530, 316]. left-bounded [530]. Left-truncated [601, 316]. Leibler [256]. Lemmon [156]. length [51, 500]. León [206]. Lernsoftwareangebote [90]. Level [286, 56, 29, 537]. levels [451]. Lévy [454, 379]. Lévy-driven [379]. liability [63]. Life [296, 157, 195]. lifetime [455]. Likelihood [400, 461, 499, 341, 383, 391, 533, 337, 215, 301, 162, 514, 407, 335, 234, 441, 435, 336, 505, 564]. Likelihood-based [400]. likelihoods [159]. limitation [191]. Lindley [204]. Linear [55, 510, 192, 497, 499, 268, 393, 434, 173, 214, 502, 335, 2, 405, 604, 98, 188, 472, 457, 375, 207]. linearity [128]. link [511, 503]. linkage [198, 100]. list [196]. Literatur [8, 24]. Lives [157]. LM [164]. Local [562, 395, 532, 449, 325, 539, 599]. location [132, 364, 389, 359]. locations [110]. log [562, 599, 405, 277, 375]. log-Gaussian [562, 599]. log-linear [375]. logarithmic [362, 386]. logistic [443, 213, 525]. logit [136]. Lohndifferenzen [62]. long [491, 436, 534, 585, 462]. long-memory [462]. long-range [585]. long-run [436]. Longitudinal [285, 134, 485, 487, 372, 242, 471, 456, 333, 283]. Lorenz [122]. Loss [214, 300]. Lösungsansätze [108]. Louhichi [206]. love [150]. Löwenbein [237]. lower [390]. lower-order [390]. LS [235, 428]. lumpability [276].

M [166]. macroeconometric [71]. Maddalena [327]. magnetic [466]. Make [191]. malaria [324]. Mallows [494]. management [402, 236]. Mandel [314]. Mantel [427]. mapping [559, 327]. Marginal [215, 276, 301, 330, 159]. Marginal-sum [215, 301]. Marginalized [511]. Mark [112]. marker [527]. Market [47, 46, 48, 1, 20, 107]. marketing [209]. Markov [520, 276, 372, 228, 305, 566, 165]. match [578]. matching [212, 139]. matrices [561]. matrix [528, 268, 352, 590, 183]. matter [322]. Max [115]. maximization [455]. maximizes [403]. maximizing [160]. Maximum [461, 301, 200, 215]. Maximum-likelihood [301, 215]. Maxwell [357]. MCD [12]. McDonald [306]. MDCgo [474]. mean [476, 493, 221, 212, 352, 502, 503, 21, 462]. mean-variance [221]. means [411, 3, 27]. measure [539, 243]. Measurement [241, 104, 240, 131, 431, 383, 594, 242, 94, 140, 190, 267, 82, 457, 458, 523]. measurements [586, 259]. measures [513, 560, 211, 459, 483, 275]. Measuring [423, 537, 163, 174]. Median [233, 15]. medians [389]. Medical [151, 595]. Meeting [104, 239, 153]. Mehr [6, 23]. membership [414]. memoriam [43]. memory [491, 534, 585, 462]. meta [426, 506, 435]. meta-analysis [426, 506, 435]. Metamodeling [260]. method [189, 310, 453, 4, 441, 309, 100, 404, 444]. Methoden [115]. methodology [198, 425, 354]. Methods [114, 125, 87, 195, 109, 153, 191, 602, 421, 381, 82, 518, 469, 537, 415, 202]. Michael [166, 185]. micro [31, 80]. micro-data [31]. microaggregation [98, 188]. Microcensus [141]. Microdata [47, 19, 32]. Microeconometric [80, 83]. Microeconometrics [187]. Mikrodaten [32, 31]. minimax [28, 220]. minimizers [12]. minimizing [528]. Minimum [443, 416, 413, 121, 480, 390, 588, 604]. minimum-variance [480]. misclassified [498]. Mises [448]. Missing [109, 107, 104, 510, 198, 593, 414, 407, 335, 405, 105, 507, 564, 289, 196, 108]. mixed [430, 512, 506, 497, 434, 291, 180, 526, 604, 63]. mixed-effect [497, 434]. Mixture [199, 430, 340, 291, 283]. mixtures [392, 400, 533, 162, 373, 545, 412]. mobility [40]. mode [367]. Model [497, 595, 545, 223, 430, 485, 510, 536, 340, 538, 448, 290, 351, 26, 268, 377, 299, 235, 383, 391, 500, 394, 414, 173, 214, 199, 471, 438, 577, 4, 456, 496, 379, 566, 246, 229, 468, 98, 188, 21, 378, 501, 428, 458, 349]. Model-based [545, 414, 577]. Modeling [538, 394, 430, 353, 239, 600, 326, 287, 322, 581, 572, 517, 466, 244, 282, 203, 324, 182, 116]. Modelling [392, 259, 596, 3, 245, 573, 242, 579, 566, 421, 605, 571]. Models [142, 207, 479, 504, 409, 280, 172, 454, 422, 192, 81, 72, 380, 321, 497, 520, 71, 499, 542, 297, 276, 395, 286, 567, 437, 452, 568, 385, 191, 451, 406, 434, 227, 522, 594, 372, 228, 315, 88, 74, 291, 301, 521, 78, 572, 241, 305, 330, 384, 264, 79, 511, 136, 373, 209, 502, 407, 335, 211, 234, 441, 459, 329, 97, 128, 494, 2, 296, 246, 388, 257, 432, 258, 283, 293, 405, 80, 137, 604, 336, 82, 396, 374, 346, 555, 138, 332, 370, 251, 450, 508, 523, 507, 543, 444, 564, 375, 570]. Models [134]. moderated [240]. moderation [368]. modification [528]. Modified [2, 449, 234]. Mohan [202]. Molenberghs [134]. Moments [155]. momentum [566]. monitoring [530, 120, 121]. monotonicity [356]. Monte [269, 170, 57, 165]. Mora [166]. mortality [315]. motion [119]. movement [568, 421]. movers [141]. moving [16, 495]. Multi [226, 541, 591]. multi-period [591]. Multi-sample [226, 541]. multidimensional [451, 241, 531]. multidimensionality [241]. multilateral [591]. Multilevel [78, 451]. multimethod [245]. Multinomial [396, 443, 136]. Multiple [262, 304, 593, 193, 198, 433, 471, 397, 489, 106, 165, 543, 444, 196]. multiplicative [189, 215]. multistage [475]. Multistate [297]. Multitrait [245]. Multitrait-multimethod [245]. Multivariate [42, 219, 401, 122, 327, 267, 457, 485, 535, 323, 431, 120, 400, 406, 488, 11, 16, 563, 557, 527, 211, 464, 463, 598, 566, 501, 374, 587]. multiway [179]. mutual [365]. MVE [12]. MZE [12].

National [580]. Nature [184]. Nearest [316]. necessary [529]. needs [153]. negative [492]. negatively [428]. neighbor [316]. nested [265]. net [535, 587]. network [536, 572]. Neue [108]. neutral [227]. next [29]. Neyman [475]. Neyman-type [475]. Nichtparametrische [62]. noch [91]. noise [189]. noisy [380]. Non [434, 280, 141, 538, 231, 465, 265, 496, 361, 598, 111, 460, 555, 62]. Non-concave [434]. non-coverage [141]. non-equidispersed [555]. non-homogeneous [496]. non-nested [265]. non-normality [598]. Non-parametric [62]. non-random [538]. non-standard [361]. non-stationary [280, 465, 460]. non-university-sector [111]. nonignorable [510]. noninformative [426]. Nonlinear [244, 404, 172, 120, 352, 521, 78, 243, 293, 99, 282, 564]. nonmonotone [507]. Nonparametric [366, 79, 16, 489, 113, 292, 541, 392, 383, 1, 168, 468, 531]. nonresponse [40, 181, 84, 36, 41]. nonsense [67]. nonseparable [441]. nonstationarity [491]. Nonstationary [368]. Nonstationary-volatility [368]. nonuniform [240]. norm [528]. normal [391, 400, 271, 464, 463, 416]. normality [353, 514, 598, 349]. normalization [445]. normalizing [568]. normally [607]. note [175, 178, 95, 483]. notes [206]. now- [552]. number [537]. Numerical [307]. nutritional [417]. Nutzung [59].

Oaxaca [172]. Obituary [237]. object [18]. objective [11]. observables [592]. observation [228]. observational [484]. obtained [160]. occupancy [420]. odds [405]. Oded [237]. off [300]. offerings [29]. offices [59]. official [109, 32, 108]. Ogive [391]. OK [143]. Ökonometrie [7]. older [32]. Oliver [577]. OLS [520]. omitted [243]. one [385, 514, 378]. one-sided [385]. one-step [378]. one-way [514]. ones [264]. operating [86]. operations [182]. operator [309]. opportunities [46]. Optimal [256, 527, 447, 523, 28, 403, 200, 344, 307]. optimality [233]. optimization [528]. options [170]. Order [460, 442, 362, 390, 440, 526, 246]. Ordered [81, 408, 501, 370]. ordering [522, 96]. ordinal [223, 593, 474]. origin [530]. Orthogonal [179]. Ost [110, 62]. ost- [62]. Other [185, 147, 519]. our [29]. outbreak [600]. outcome [506]. outcomes [487, 572, 471]. outlier [132]. outlier-adjusted [132]. outliers [399]. output [590, 463, 464]. outputs [259]. overall [346]. overdispersed [504]. overqualification [232]. overview [83, 421].

paired [372, 94, 140, 370]. pairs [397].

Pairwise [391]. pandemic [574, 547, 551, 546, 549, 548]. Panel [58, 280, 192, 286, 193, 40, 368, 279, 78, 281, 285, 229, 432, 181, 282, 41, 141, 51]. panel-multiple [193]. panels [271, 331]. paradigm [482]. Parameter [393, 585]. parameters [221, 604, 416]. parametric [487, 62, 503, 588, 555]. Pareto [508]. Pareto-type [508]. Paris [287, 20]. parsimonious [545]. part [383]. partial [393, 244, 472]. partially [499, 408, 335, 457, 564]. particle [323]. particular [301]. particulate [322]. partitioning [577]. Pascal [114, 125]. passing [572]. patent [536]. Paul [206]. Pearson [490, 155]. Pedro [166]. penalisation [571]. penalization [434]. Penalized [499, 336, 337, 373, 453, 441, 432, 161]. penalty [573]. people [147]. Performance [331, 431, 581, 304, 315, 269, 211, 381, 555]. period [591]. periods [51]. Perron [99]. Perron-type [99]. personal [146]. perspective [559, 480, 105]. perspectives [48]. Perspektiven [48]. Perturbation [189]. Peter [203]. phenomenon [479]. phi [411, 443]. phi-divergence [411, 443]. Phillips [99]. piece [521]. piecewise [415]. Pietra [569]. plans [470, 86, 220, 344]. players [573]. Plenary [104]. PLUS [116]. pMST [364]. point [515, 48, 168, 425, 329, 218, 179]. point-scale [515]. point-symmetry [179]. Poisson [63, 504, 574, 357, 496, 526, 63]. Poisson-exponential-Tweedie [504]. Poissonity [607]. polarization [131]. Policy [149, 48, 1]. pollution [323]. polynomial [523, 444]. polynomials [553]. poor [243]. popularity [516]. population [292, 422, 453, 588]. populations [430, 411, 233, 470]. portfolio [403, 413, 528, 121, 480, 127, 160]. Positive [528, 303]. Positive-definite [528]. possibly [280, 460]. posteriori [298].

Power [96, 445, 377, 169, 604]. practical [321, 421]. practice [262]. Precarious [110]. precise [275]. predicted [502]. Predicting [516, 567]. Prediction [380, 323, 414, 173, 157, 570, 354, 214, 365, 438, 300]. Prediction-based [380]. predictions [558]. predictive [292, 212]. predictor [269]. predictors [596]. preference [559]. prehistory [394]. preisbereinigten [67]. Preisindikatoren [30]. Prekäre [110]. premature [575]. premium [354, 161]. presence [439, 592, 267, 275, 457]. Preservation [382]. pressed [218]. previous [100]. price [123, 429, 591, 67, 20, 224, 263, 605, 100, 537, 30]. price-adjusted [67]. price-method [100]. Prices [129, 287, 305, 330]. pricing [596, 356, 170, 379]. Prieur [206]. Principal [289, 589, 267, 347, 222]. prior [597]. priori [298]. priors [426, 291]. Probabilistic [576]. probabilities [138]. probability [342, 561, 577]. Probit [595, 452, 501]. problem [18]. problems [307, 105, 196]. procedure [513, 515, 28, 221, 531]. Procedures [54, 433, 27, 467, 317]. process [369, 236, 291, 425, 496, 526, 509, 267, 259, 63]. processes [454, 402, 530, 562, 599, 325, 341, 228, 465, 440, 534, 398, 584, 218, 525]. processing [535]. Product [66, 361, 155, 59]. Production [65]. productivity [284]. Products [55]. Produktdiversifizierung [59]. professional [569]. profit [229]. profit-sharing [229]. programmes [89]. progress [549]. projection [42, 13, 175, 524]. projection-based [524]. Prokhorov [200]. propensity [554, 139]. Properties [65, 306, 367, 431, 235, 219, 433, 96, 386, 249, 584, 346]. property [13, 175]. Proportions [124]. proposal [31]. protection [549]. Prozesse [63]. pseudo [341, 332]. pseudo-likelihood [341]. pseudo-structural [332]. psycho [581]. psycho-social [581]. Purchase [209]. purchasing [169].

Quality [154, 36]. quantification [374]. Quantile [409, 487, 559, 77, 311, 471, 407, 335, 318, 605, 450, 508, 564]. quantiles [366, 558, 486]. quantitative [466]. quantization [200]. Quarterback [580]. questionnaires [38]. questions [40, 39]. quotas [67]. Quoten [67].

R [206, 176, 142]. R&D [284]. Radhakrishna [207]. radial [438]. radioactive [327]. Rafael [206]. rainfall [326, 379]. Random [55, 66, 525, 426, 538, 593, 511, 440, 407, 224, 249, 371, 524, 507]. random-effects [426, 511]. randomized [363]. randomly [450]. randomness [556]. range [585]. rank [522, 408, 505]. rank-based [522]. rank-ordered [408]. Ranked [553, 494, 495]. ranking [494, 98, 188]. ranks [531]. Rao [207]. Rasch [377, 437]. rate [567, 382, 374]. rates [353, 573]. ratings [538]. Ratio [66, 403, 337, 162, 514, 234, 361, 249, 405, 160, 277]. Ratios [55, 87, 343, 124]. raw [37]. RDELA [364]. reading [38]. real [269, 30]. realised [275]. Realisierung [31]. realized [426]. really [529]. recapture [195, 424, 197, 196, 199]. reconstruction [590]. Record [18, 198]. recovery [587]. recurrent [439, 570]. recursive [577, 258]. redescending [132]. Reduction [64, 577, 404]. regime [401, 305, 330]. regime-switching [305, 330]. Regional [552, 162]. regions [519]. register [101]. Regression [309, 479, 292, 481, 409, 574, 487, 443, 448, 290, 484, 268, 559, 164, 235, 385, 77, 191, 602, 521, 170, 310, 471, 407, 335, 473, 365, 597, 459, 128, 468, 336, 272, 605, 346, 472, 469, 428, 450, 457, 525, 404, 508, 505, 507, 444, 564, 571, 113]. regression-based [292, 164, 170]. regressions [442, 311, 503]. regressor [280, 367]. regularization [535, 373, 553]. regularized [434, 570]. Reiss [185]. Relative [488, 129, 169, 531]. relaxation [28]. reliability [495, 157]. repeated [586, 483]. repetition [94, 140]. Reply [23]. report [506]. reported [552]. representation [561]. represented [447]. reproducibility [342]. Resampling [389, 19]. Resampling-based [389]. Research [47, 147, 46, 48, 59, 107, 283, 111, 49]. residential [141, 30]. residuals [490]. resonance [466]. response [292, 510, 363, 81, 478, 551, 242, 502, 407, 335, 444]. responses [39, 436, 564]. restricted [240]. restrictions [436]. Results [58, 51, 7, 5, 250, 148]. retirement [285]. return [351]. Review [202, 116, 206, 204, 156, 166, 176, 185, 115, 114, 125, 157, 142, 133, 184, 183, 207, 203, 205, 113, 134, 143, 338, 209, 254, 250, 601, 462]. reviews [44, 52, 60, 68]. revisions [180]. ridge [597]. Riesz [561]. risk [298, 227, 522, 480, 161]. risks [299]. risky [403]. Robert [176]. Robust [533, 168, 343, 354, 604, 210, 368, 486, 265]. Robustness [15]. ROC [500, 509]. role [548, 547, 572, 551, 546, 549, 111]. Rolf [185]. Rolf-Dieter [185]. Rolle [111]. Root [58, 399, 368, 99, 331, 73, 587]. Rounded [39]. rounding [250]. row [535]. row-elastic-net [535]. RR [531]. RR-classifier [531]. rules [325, 509]. run [436, 4, 300]. run-off [300].

S [176, 116]. S-PLUS [116]. Sam [157]. Samanta [202]. same [51]. sample [541, 413, 443, 385, 226, 475, 514, 378, 277, 275]. sampled [544]. samples [408]. sampling [424, 343, 475, 86, 220, 344, 307, 494, 255, 95, 358, 495, 114, 125, 205]. Sana [206]. Sardinia [327]. satisfying [13, 175]. Saunders [157]. SAVE [317]. scalar [469]. scalar-on-function [469]. scale [132, 515, 530, 71, 400, 269, 466, 359, 524]. scatter [364]. Schafer [156]. scheme [95]. schemes [464, 463]. Schmid [112]. science [195, 32]. Scientists [157]. SCOMDY [329]. score [139]. scores [554, 527]. Scoring [558, 325, 578]. Seasonal [58]. seasons [576]. Second [246, 600]. Second-order [246]. sectional [287]. sector [111]. seeing [166]. Seeking [152]. seemingly [346]. seems [529]. seismic [562, 599]. selecting [553]. Selection [54, 513, 497, 448, 535, 592, 602, 434, 338, 269, 373, 348, 127, 336, 378, 460, 472]. selectors [338]. Self [492, 398, 310, 38]. self-administered [38]. Self-excited [492]. Self-exciting [398]. SEM [239, 243]. semi [487]. semi-parametric [487]. semicontinuous [479]. Semiparametric [212, 136, 378, 227, 605]. semivarying [383]. Sense [67]. sensitivity [498]. sequences [309]. Sequential [121, 4]. serial [342, 163]. Series [116, 65, 64, 490, 353, 542, 366, 248, 352, 228, 264, 310, 362, 376, 293, 176, 585, 460, 163, 182, 462]. Service [157]. set [408, 494, 495]. sets [416]. setting [448, 296]. several [444]. Shalabh [207]. shall [420]. shape [382]. shapes [456]. share [343]. sharing [229]. Sharpe [403, 160]. sheets [318]. shift [242, 389]. Shifts [56, 493]. short [602, 436, 140]. short-run [436]. short-term [140]. Shumway [176]. Sicht [48]. sided [385, 470]. Sieve [585]. signal [524]. signature [260]. signed [234]. SIMEX [190, 458]. simple [481, 128]. simpler [529]. simplicial [178]. simply [264]. Simulation [58, 240, 12, 189]. Simulations [57]. Simultaneous [530, 312, 582, 318, 132, 71, 389]. single [353, 344, 458]. single-index [458]. single-sampling [344]. singular [413]. Sinn [67]. size [354]. skew [369, 400, 213]. Skewness [478]. Skewness-adjusted [478]. Small [544, 385, 413, 514, 21]. Small-sample [385]. smooth [536]. Smoothed [407, 335]. smoothing [170, 258]. SNR [201]. soccer [574, 567]. social [195, 581, 22]. Socio [51, 141]. Socio-Economic [51, 141]. socioeconomic [544]. SOEP [37]. Software [156, 90]. solution [307]. solutions [108, 114, 125]. Some [560, 109, 419, 57, 82, 346, 301, 250]. Song [203]. sources [506, 197, 569]. Sozialstatistik [22]. space [579, 441, 21, 282, 531, 324, 251, 524]. space-time [441, 324]. sparse [587]. sparsity [553]. Spatial [320, 424, 556, 449, 562, 599, 367, 395, 325, 323, 169, 406, 427, 425, 438, 466, 327, 496, 381, 605, 537]. Spatio [322, 321, 323, 394]. Spatio-temporal [322, 321, 323, 394]. SPDE [322]. Special [295, 323, 565, 279, 418, 327]. Species [420]. Specific [269]. Specific-to-general [269]. specification [71, 468]. spline [291, 170, 432, 258]. splines [337, 161]. split [233]. sports [565, 582, 570]. spot [305, 330]. square [138, 587]. square-root [587]. squared [21]. Squares [207, 244, 272]. Stage [54]. standard [86, 220, 361]. Standing [576]. standings [575]. state [579, 21, 282, 324, 251]. state-space [579, 282]. stationary [280, 341, 465, 376, 585, 460]. statistic [314]. STATISTICA [143]. Statistical [298, 221, 352, 145, 421, 156, 184, 146, 148, 402, 153, 419, 236, 522, 59, 27, 105, 154, 185]. Statistics [153, 565, 54, 418, 411, 151, 548, 442, 46, 109, 320, 48, 6, 206, 547, 425, 551, 546, 90, 29, 22, 2, 108, 549, 5, 166, 150, 23, 91, 295, 7, 63, 183]. Statistik [46, 90, 108, 91, 7, 115, 63]. Statistikausbildung [6, 23]. Statistischen [59]. StatSoft [143]. step [515, 378]. still [91]. stochastic [586, 380, 576, 401, 232, 246, 537]. Stoffer [176]. stopped [576]. strategies [304, 310]. stratified [307]. streaks [578]. Strong [353, 428]. Structural [75, 517, 239, 87, 71, 581, 311, 242, 502, 436, 234, 244, 332].

Structurally [196]. structure

[490, 340, 229, 455]. structured [137]. structures [240]. student [87]. studies [197, 484, 320, 435]. Studium [115]. Study [454, 240, 12, 262, 285, 577, 444]. subject [414]. subjective [410]. subpopulations [569]. subsample [51]. success [59]. Sucre [324]. sum [215, 301, 446]. sums [124]. super [386]. superadditive [428]. superadditive-dependent [428]. Supermigrative [303]. sure [191]. surface [404]. Surveillance [465, 352, 552]. Survey [84, 461, 262, 7, 19, 274, 333, 347, 182, 148, 317]. surveys [363, 40, 420, 181, 106, 446, 37, 165]. survival [454, 597]. sustainability [563, 557]. switching [520, 305, 330, 137]. Symmetric [250, 223, 417, 561]. symmetry [179, 138]. system [455]. system-based [455]. systematic [22, 140]. Systems [56, 198, 393].

tables [590, 296, 405, 179, 138]. tail [445]. tailed [521, 545]. takes [474]. Takezawa [113]. Taking [29]. Tapas [202]. tariff [215]. teams [569]. technical [22]. Techniken [107]. technique [446]. techniques [127, 107, 293]. temporal [321, 423, 322, 323, 394]. term [140]. terminal [310]. terms [243]. Test [280, 448, 58, 541, 510, 411, 119, 413, 556, 164, 427, 399, 162, 514, 265, 234, 397, 435, 468, 314]. Testing [490, 276, 592, 356, 128, 157, 476, 385, 337, 433, 330, 489, 503, 518, 469, 73, 450]. Tests [586, 311, 56, 377, 226, 304, 606, 368, 271, 607, 99, 462, 331]. text [602]. texts [602]. their [560, 384, 79]. theoretical [105, 601]. Theory [183, 202]. thermomechanically [259]. Thinning [182]. Thomas [185]. Thompson [184]. three [405]. three-way [405]. threshold [440, 398, 525]. thresholding [587]. thresholding-based [587]. Tillé [114, 125, 205]. Time [116, 65, 64, 490, 353, 542, 366, 248, 352, 264, 170, 373, 376, 517, 579, 441, 584, 293, 585, 460, 324, 163, 182, 462, 570, 176]. time-to-event [570]. timing [209]. tool [576]. tools [519]. total [284]. Toutenburg [207]. tracking [580]. trading [263]. training [89]. trajectory [563, 557]. transformation [3]. transformations [96, 527]. transformed [188]. transition [485]. transitions [387]. treatment [554, 83, 592, 84, 447]. treatments [506]. Trede [112]. tree [388]. trend [578]. trends [423]. triangles [300]. triangulation [456]. trimming [210]. Trivariate [595]. True [248]. truncated [473, 316, 601, 450, 505]. Tulsa [143]. Tweedie [479, 504, 384]. tweets [516]. Two [595, 470, 54, 411, 515, 197, 439, 210, 191, 162, 521, 514, 569, 381, 41]. two-component [162]. two-panel-waves [41]. two-piece [521]. Two-sided [470]. Two-Stage [54]. two-step [515]. two-way [514]. type [491, 606, 475, 155, 99, 516, 508, 571]. types [210].

U.S. [47]. ultra [504]. ultra-overdispersed [504]. unbalanced [331]. unbiased [160]. Uncertainty [374, 538, 204]. unconditional [286]. underdispersion [555]. understanding [509, 204]. Unemployment [51, 58, 89, 181]. Unified [507, 476]. uniform [223, 353, 240, 138]. Uniformly [514, 371]. Union [557, 563]. Uniqueness [442]. Unit [58, 73, 399, 368, 99, 331]. United [109]. units [271]. universitären [22]. universities [148]. University [87, 22, 111]. unknown [470, 86, 220, 588, 455, 317]. unmeasured [498]. unrelated [346]. unsampled [544]. Unsinn [67]. unstructured [596]. Unternehmenserfolg [59]. usability [425]. Usage [595]. use [520, 59, 529, 107]. used [123]. Useful [264, 152]. user [153]. Using [240, 77, 242, 54, 57, 258, 64, 223, 490, 485, 515, 141, 454, 574, 87, 506, 392, 448, 325, 262, 578, 198, 500, 394, 521, 572, 453, 438, 598, 580, 389, 446, 267, 318, 516, 161, 495, 570].

V [204]. vague [426]. Valero [166].

Valero-Mora [166]. validation [299]. value [515, 299, 385, 248, 361, 97]. valued [542, 440, 526, 525, 375]. Values [185, 603, 67, 314, 289]. VAR [520, 332, 88, 4]. Variable [535, 513, 430, 510, 602, 373, 470, 348, 472, 371, 222]. Variables [595, 55, 66, 515, 499, 394, 86, 220, 344, 75, 335, 224, 249, 188, 349]. Variance [422, 502, 64, 493, 221, 413, 121, 480, 94, 140, 468]. variances [27, 604]. Variation [64, 479, 542, 488, 275]. variational [137]. Varying [451, 409, 450]. varying-coefficient [450]. vector [352, 75]. vectors [476]. Venezuela [324]. Verbeke [134]. Vergleich [110, 90]. Verkettung [100]. verschiedener [90]. Version [143]. versus [436]. via [530, 535, 348, 545, 540, 412, 587]. view [236, 48, 146]. VII [155]. Vine [484, 597]. Visual [166]. Visualizing [509]. vocational [89]. volatilities [415]. volatility [380, 401, 368]. Volks [6, 23]. Volks- [6, 23]. volume [416, 100]. Volumenanteile [100]. Vorjahrespreismethode [100]. Vorschlag [31]. voter [387]. VWL [115].

W [166]. Wachstumsbeiträge [100]. wage [62, 229, 41]. Wald [2]. Wang [116]. Watson [226]. wavelet [290]. waves [286, 41]. way [514, 405]. ways [281]. Weak [452, 428, 206]. web [29]. Weibull [249]. Weighted [450, 533, 382, 534, 272, 15]. weights [480, 160]. Weitergabe [32]. Werten [67]. West [110, 110, 62, 229]. West-Vergleich [110]. westdeutschen [62]. Wetzel [43]. Wewel [115]. where [529]. which [403]. Whittle [491]. Whittle-type [491]. Wilkinson [133]. winning [577]. Wirtschafts [22]. Wirtschafts- [22]. Wirtschaftsforschungsinstitute [111]. wirtschaftspolitischen [48]. Wirtschaftsstatistik [6, 23]. wirtschaftsstatistischer [5]. wirtschaftswissenschaftlichen [7]. wise [374]. Wishart [401]. Wissenschaft [32]. without [94, 140]. Wohnimmobilien [30]. Wolfgang [43]. Work [595]. working [239]. world [590]. wrapped [512]. wrong [264]. Württemberg [169].

X [203]. X.-K [203].

year [100]. yield [51]. Young [166]. Yves [114, 125, 205].

zeitgemäß [91]. zero [487, 362]. zeroinflated [487, 362]. Zivot [116]. zonoids [122]. Zur [111, 31, 23, 32]. zwischen [62].

References

Frolich:2004:DIP

[1] Markus Frölich and Patrick A. Puhani. Developing an immigration policy for Germany on the basis of a nonparametric labor market classification. AStA. Advances in Statistical Analysis, 88(1):1–22, February 2004. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s101820400156.

Oelerich:2004:MWS

[2] Andreas Oelerich and Thorsten Poddig. Modified Wald statistics for generalized linear models. AStA. Advances in Statistical Analysis, 88(1):23–34, February 2004. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s101820400157.

Fischer:2004:KMM

[3] Matthias Fischer and Ingo Klein. Kurtosis modelling by means of the Jtransformation. AStA. Advances in Statistical Analysis, 88(1):35–50, February 2004. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s101820400158.

Mihailescu:2004:SME

[4] Laurentiu Mihailescu. A sequential method for the evaluation of the VaR model based on the run between exceedances. AStA. Advances in Statistical Analysis, 88(1):51–72, February 2004. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s101820400159.

Rosemann:2004:EEF

[5] Martin Rosemann, Daniel Vorgrimler, and Rainer Lenz. Erste Ergebnisse faktischer Anonymisierung wirtschaftsstatistischer Einzeldaten. (German) [First results of factual anonymization of economic statistics data items]. AStA. Advances in Statistical Analysis, 88 (1):73–99, February 2004. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL http:

//link.springer.com/article/10.
1007/s101820400160.

Grunewald:2004:DMW

[6] Werner Grünewald, Hans-Joachim Mittag, Michael Müller, Reiner Stäglin, Peter Lorscheid, and Roland Gnoss. Diskussionsbeiträge zu “Mehr Wirtschaftsstatistik in der Statistikausbildung für Volksund Betriebswirte”. (German) [Discussions on “More economic statistics in statistics education for economists and business economists”]. AStA. Advances in Statistical Analysis, 88 (1):100–117, February 2004. CO-DEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s101820400161. See reply

[23].

Frohn:2004:SOW

[7] Joachim Frohn. Statistik und Ökonometrie an wirtschaftswissenschaftlichen Fakultäten — Ergebnisse einer Befragung. (German) [Statistics and econometrics in economics faculties — results of a survey]. AStA. Advances in Statistical Analysis, 88(1):118–125, February 2004. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s101820400162.

Anonymous:2004:LBa

[8] Anonymous. Literatur /books. AStA. Advances in Statistical Analysis, 88 (1):126–132, February 2004. CO-DEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s101820400163.

Anonymous:2004:HCa

[9] Anonymous. Help & contacts. AStA. Advances in Statistical Analysis, 88(1): ??, February 2004. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

Mosler:2004:IGD

[10] Karl Mosler. Introduction: The geometry of data. AStA. Advances in Statistical Analysis, 88(2):133–135, May 2004. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s101820400164.

Koshevoy:2004:GMO

[11] Gleb A. Koshevoy, Jyrki Möttönen, and Hannu Oja. On the geometry of multivariate L_{1} objective functions. AStA. Advances in Statistical Analysis, 88(2):137–154, May 2004. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s101820400165.

Becker:2004:MMM

[12] Claudia Becker and Sebastian Paris Scholz. MVE, MCD, and MZE: a simulation study comparing convex body minimizers. AStA. Advances in Statistical Analysis, 88(2):155–162, May 2004. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s101820400166.

Dyckerhoff:2004:DDS

[13] Rainer Dyckerhoff. Data depths satisfying the projection property. AStA. Advances in Statistical Analysis, 88 (2):163–190, May 2004. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s101820400167. See note [175].

Romanazzi:2004:DDC

[14] Mario Romanazzi. Data depth and correlation. AStA. Advances in Statistical Analysis, 88(2):191–214, May 2004. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s101820400168.

Zuo:2004:RWD

[15] Yijun Zuo. Robustness of weighted L^{p}-depth and L^{p}-median. AStA. Advances in Statistical Analysis, 88 (2):215–234, May 2004. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL http:

//link.springer.com/article/10.
1007/s101820400169.

Liu:2004:DCN

[16] Regina Y. Liu, Kesar Singh, and Julie H. Teng. DDMA-charts: Nonparametric multivariate moving average control charts based on data depth. AStA. Advances in Statistical Analysis, 88(2):235–258, May 2004. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s101820400170.

Anonymous:2004:HCb

[17] Anonymous. Help & contacts. AStA. Advances in Statistical Analysis, 88(2): ??, May 2004. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

Neiling:2004:GAR

[18] Mattis Neiling and Hans-J. Lenz. The German Administrative Record Census — an object identification problem. AStA. Advances in Statistical Analysis, 88(3):259–277, August 2004. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s101820400171.

Gottschalk:2004:MDR

[19] Sandra Gottschalk. Microdata disclosure by resampling — empirical findings for business survey data. AStA. Advances in Statistical Analysis, 88(3):279–302, August 2004. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s101820400172.

Maurer:2004:HPI

[20] Raimond Maurer, Martin Pitzer, and Steffen Sebastian. Hedonic price indices for the Paris housing market. AStA. Advances in Statistical Analysis, 88(3):303–326, August 2004. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s101820400173.

Schumacher:2004:CMS

[21] Christian Schumacher. A comparison of mean squared error approximations for a small estimated state space model. AStA. Advances in Statistical Analysis, 88(3):327–345, August 2004. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s101820400174.

Litz:2004:CFA

[22] Hans Peter Litz. Curriculare und fachsystematische Aspekte einer universitären Wirtschafts- und Sozialstatistik. (German) [Curricular and systematic technical aspects of university economic and social statistics]. AStA. Advances in Statistical Analysis, 88(3):347–361, August 2004. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s101820400175.

vonderLippe:2004:ADM

[23] Peter von der Lippe and Sibylle Schmerbach. Antwort zur Diskussion um “Mehr Wirtschaftsstatistik in der Statistikausbildung für Volks- und Betriebswirte”. (German) [Reply to the discussion on “More economic statistics in statistics education for economists and business economists”]. AStA. Advances in Statistical Analysis, 88(3):362–367, August 2004. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s101820400176.

Anonymous:2004:LBb

[24] Anonymous. Literatur /books. AStA. Advances in Statistical Analysis, 88 (3):368–374, August 2004. CO-DEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s101820400177.

Anonymous:2004:HCc

[25] Anonymous. Help & contacts. AStA. Advances in Statistical Analysis, 88(3): ??, August 2004. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

Christmann:2004:AMC

[26] Andreas Christmann. An approach to model complex high-dimensional insurance data. AStA. Advances in Statistical Analysis, 88(4):375–396, December 2004. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s101820400178.

Klaver:2004:ECH

[27] Hendrik Kläver and Friedrich Schmid. The effect of conditional heteroskedasticity on common statistical procedures for means and variances. AStA. Advances in Statistical Analysis, 88(4):397–407, December 2004. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s101820400179.

Begun:2004:RPC

[28] Alexander Begun and Wilfried Seidel. A relaxation procedure for calculating (Γ-)minimax optimal designs. AStA. Advances in Statistical Analysis, 88(4):409–425, December 2004. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s101820400180.

Kopsch:2004:HCS

[29] Günter Kopsch, Michael Neutze, and Annette Pfeiffer. How to communicate statistics — taking our web offerings to the next level. AStA. Advances in Statistical Analysis, 88(4):427–434, December 2004. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s101820400181.

Leifer:2004:PWD

[30] Hans-Albert Leifer. Preisindikatoren für Wohnimmobilien in Deutschland.

(German) [Price indicators for residential real estate in Germany]. AStA. Advances in Statistical Analysis, 88(4):435–450, December 2004. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s101820400182.

Merz:2004:KMK

[31] Joachim Merz. Kumulation von Mikrodaten — Konzeptionelle Grundlagen und ein Vorschlag zur Realisierung. (German) [Accumulation of micro-data — conceptual foundations and a proposal for implementation]. AStA. Advances in Statistical Analysis, 88(4):451–472, December 2004. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s101820400183.

Luttinger:2004:WAA

[32] Paul Lüttinger and Heike Wirth. Zur Weitergabe von älteren amtlichen Mikrodaten an die Wissenschaft. (German) [On the dissemination of older official microdata in science]. AStA. Advances in Statistical Analysis, 88 (4):473–486, December 2004. CO-DEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s101820400184.

Anonymous:2004:D

[33] Anonymous. Danksagung. AStA. Advances in Statistical Analysis, 88(4): 487–488, December 2004. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s101820400185.

Anonymous:2004:HCd

[34] Anonymous. Help & contacts. AStA. Advances in Statistical Analysis, 88(4): ??, December 2004. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

Seidel:2005:E

[35] Wilfried Seidel. Editorial. AStA. Advances in Statistical Analysis, 89 (1):1, February 2005. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/s101820500186.

Riphahn:2005:IIN

[36] Regina T. Riphahn. Introduction: Item nonresponse and data quality. AStA. Advances in Statistical Analysis, 89(1):3–5, February 2005. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s101820500187.

Schraepler:2005:CIF

[37] Joerg-Peter Schraepler and Gert G. Wagner. Characteristics and impact of faked interviews in surveys — an analysis of genuine fakes in the raw data of SOEP. AStA. Advances in Statistical Analysis, 89(1):7–20, February 2005. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s101820500188.

Redline:2005:FIR

[38] Cleo D. Redline, Don A. Dillman, Lisa Carley-Baxter, and Robert H. Creecy. Factors that influence reading and comprehension of branching instructions in self-administered questionnaires. AStA. Advances in Statistical Analysis, 89(1):21–38, February 2005. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s101820500189.

Hanisch:2005:RRI

[39] Jens U. Hanisch. Rounded responses to income questions. AStA. Advances in Statistical Analysis, 89(1):39–48, February 2005. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s101820500190.

Frick:2005:INI

[40] Joachim R. Frick and Markus M. Grabka. Item nonresponse on income questions in panel surveys: Incidence, imputation and the impact on inequality and mobility. AStA. Advances in Statistical Analysis, 89(1):49–61, February 2005. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s101820500191.

Spiess:2005:EIN

[41] Martin Spiess and Jan Goebel. On the effect of item nonresponse on the estimation of a two-panel-waves wage equation. AStA. Advances in Statistical Analysis, 89(1):63–74, February 2005. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s101820500192.

Bodnar:2005:MCC

[42] Olha Bodnar and Wolfgang Schmid. Multivariate control charts based on a projection approach. AStA. Advances in Statistical Analysis, 89(1):75–93, February 2005. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s101820500193.

Hansen:2005:WWM

[43] Gerd Hansen. Wolfgang Wetzel in memoriam. AStA. Advances in Statistical Analysis, 89(1):95–97, February 2005. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s101820500194.

Anonymous:2005:BBRa

[44] Anonymous. Buchbesprechungen /book reviews. AStA. Advances in Statistical Analysis, 89(1):98–105, February 2005. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s101820500195.

Anonymous:2005:HCa

[45] Anonymous. Help & contacts. AStA. Advances in Statistical Analysis, 89(1): ??, February 2005. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

D:2005:ASC

[46] Prof. Bernd Fitzenberger Ph. D. and Prof. Dr. Joachim Möller. Arbeitsmarkt und Statistik: Chancen und Herausforderungen für die empirische Arbeitsmarktforschung. (German) [labor market and statistics: opportunities and challenges for empirical labor market research]. AStA. Advances in Statistical Analysis, 89(2):109–119, June 2005. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-005-0196-7.

Abraham:2005:MAL

[47] Katharine G. Abraham. Microdata access and labor market research: The U.S. experience. AStA. Advances in Statistical Analysis, 89(2):121–139, June 2005. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-005-0197-6.

Franz:2005:AAS

[48] Prof. Dr. Dr. h. c. mult. Wolfgang Franz. Arbeitsmarktforschung und Arbeitsmarktstatistik aus der Sicht der wirtschaftspolitischen Beratung: Erfahrungen und Perspektiven. (German) [Labor market research and labor market statistics from the point of view of economic policy advice: experiences and perspectives]. AStA. Advances in Statistical Analysis, 89(2):141–158, June 2005. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-005-0198-5.

D:2005:DDF

[49] Prof. Jutta Allmendinger Ph. D. and Dr. Annette Kohlmann. Datenverfügbarkeit und Datenzugang am Forschungsdatenzentrum der Bundesagentur für Arbeit im Institut für Arbeitsmarkt- und Berufsforschung. (German) [Data availability and data access at the Research Data Centre of the Federal Employment Agency at the Institute for Employment Research]. AStA. Advances in Statistical Analysis, 89(2):159–182, June 2005. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-005-0199-4.

Ibsen:2005:JCD

[50] Prof. Dr. Rikke Ibsen and Prof. Dr. Niels Westergaard-Nielsen. Job creation and destruction over the business cycles and the impact on individual job flows in Denmark 1980–2001. AStA. Advances in Statistical Analysis, 89(2):183–207, June 2005. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-005-0200-2.

Biewen:2005:UDL

[51] Martin Biewen and Ralf A. Wilke. Unemployment duration and the length of entitlement periods for unemployment benefits: do the IAB employment subsample and the German Socio-Economic Panel yield the same results? AStA. Advances in Statistical Analysis, 89(2):209–236, June 2005. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-005-0201-1.

Anonymous:2005:BBRb

[52] Anonymous. Buchbesprechungen / book reviews. AStA. Advances in Statistical Analysis, 89(2):237–240, June 2005. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-005-0202-0.

Anonymous:2005:HCb

[53] Anonymous. Help & contacts. AStA. Advances in Statistical Analysis, 89(2): ??, June 2005. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

Kumar:2005:CTS

[54] Narinder Kumar, Gobind P. Mehta, and Virender Kumar. A class of twostage selection procedures using Lstatistics. AStA. Advances in Statistical Analysis, 89(3):241–261, August 2005. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-005-0203-z.

Nadarajah:2005:LCP

[55] Saralees Nadarajah and Samuel Kotz. Linear combinations, products and ratios of t random variables. AStA. Advances in Statistical Analysis, 89 (3):263–280, August 2005. CO-DEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-005-0204-y.

Trenkler:2005:EIL

[56] Carsten Trenkler. The effects of ignoring level shifts on systems cointegration tests. AStA. Advances in Statistical Analysis, 89(3):281–301, August 2005. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-005-0205-x.

Meyer:2005:UHF

[57] Mark Meyer and Peter Winker. Using HP filtered data for econometric analysis: Some evidence from Monte Carlo simulations. AStA. Advances in Statistical Analysis, 89(3):303–320, August 2005. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-005-0206-9.

Dreger:2005:PSU

[58] Christian Dreger and Hans-Eggert Reimers. Panel seasonal unit root test: Further simulation results and an application to unemployment data. AStA. Advances in Statistical Analysis, 89(3):321–337, August 2005. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-005-0207-8.

Gorzig:2005:PUN

[59] Bernd Görzig, Hartmut Bömermann, and Ramona Pohl. Produktdiversifizierung und Unternehmenserfolg: Nutzung der Forschungsdatenzentren der Statistischen Ämter. (German) [Product diversification and corporate success: use of research data centers of the statistical offices]. AStA. Advances in Statistical Analysis, 89(3):339–354, August 2005. CODEN ???? ISSN 1863-8171 (print), 1863-818X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10182-005-
0208-7.

Anonymous:2005:BBRc

[60] Anonymous. Buchbesprechungen /book reviews. AStA. Advances in Statistical Analysis, 89(3):355–363, August 2005. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-005-0209-6.

Anonymous:2005:HCc

[61] Anonymous. Help & contacts. AStA. Advances in Statistical Analysis, 89(3): ??, August 2005. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

Gorzig:2005:NDL

[62] Bernd Görzig, Martin Gornig, and Axel Werwatz. Nichtparametrische Dekomposition der Lohndifferenzen zwischen ost- und westdeutschen Betrieben. (German) [Non-parametric decomposition of wage differences between East and West German companies]. AStA. Advances in Statistical Analysis, 89(4):365–381, November 2005. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-005-0210-0.

Zocher:2005:SBG

[63] Mathias Zocher. Statistik für bivariate gemischte Poisson–Prozesse am Beispiel der Kraftfahrthaftpflichtversicherung. (German) [Statistics for bivariate mixed Poisson process: the example of automobile liability insurance]. AStA. Advances in Statistical Analysis, 89(4):383–402, November 2005. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-005-0211-z.

Schneider:2005:UAV

[64] Carsten Schneider, Gerhard Arminger, and Alexandra Schwarz. Using analysis of variance and factor analysis for the reduction of high dimensional variation in time series of energy consumption. AStA. Advances in Statistical Analysis, 89(4):403–418, November 2005. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-005-0212-y.

Flaig:2005:TSP

[65] Gebhard Flaig. Time series properties of the German Production Index. AStA. Advances in Statistical Analysis, 89(4):419–434, November 2005. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-005-0213-x.

Nadarajah:2005:PRG

[66] Saralees Nadarajah and Samuel Kotz. On the product and ratio of gamma and beta random variables. AStA. Advances in Statistical Analysis, 89 (4):435–449, November 2005. CO-DEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-005-0214-9.

Gromling:2005:SUQ

[67] Michael Grömling. Sinn und Unsinn von Quoten auf Basis von preisbereinigten Werten. (German) [Sense and nonsense of quotas on the basis of price-adjusted values]. AStA. Advances in Statistical Analysis, 89(4):451–468, November 2005. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-005-0215-8.

Anonymous:2005:BBRd

[68] Anonymous. Buchbesprechungen /book reviews. AStA. Advances in Statistical Analysis, 89(4):469–477, November 2005. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-005-0216-7.

Anonymous:2005:HCd

[69] Anonymous. Help & contacts. AStA. Advances in Statistical Analysis, 89(4): ??, November 2005. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

Hubler:2006:DND

[70] Olaf Hübler and Joachim Frohn. Developments and new dimensions in econometrics. AStA. Advances in Statistical Analysis, 90(1):1–7, March 2006. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-006-0217-1.

Chen:2006:SEL

[71] Pu Chen and Joachim Frohn. On the specification and estimation of large scale simultaneous structural macroeconometric models. AStA. Advances in Statistical Analysis, 90(1):9–25, March 2006. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-006-0218-0.

Breitung:2006:DFM

[72] Jörg Breitung and Sandra Eickmeier. Dynamic factor models. AStA. Advances in Statistical Analysis, 90(1):27– 42, March 2006. CODEN ???? ISSN 1863-8171 (print), 1863-818X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10182-006-
0219-z.

Wolters:2006:URT

[73] Jürgen Wolters and Uwe Hassler. Unit root testing. AStA. Advances in Statistical Analysis, 90(1):43–58, March 2006. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-006-0220-6.

Hassler:2006:ADL

[74] Uwe Hassler and Jürgen Wolters. Autoregressive distributed lag models and cointegration. AStA. Advances in Statistical Analysis, 90(1):59–74, March 2006. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-006-0221-5.

Lutkepohl:2006:SVA

[75] Helmut Lütkepohl. Structural vector autoregressive analysis for cointegrated variables. AStA. Advances in Statistical Analysis, 90(1):75–88, March 2006. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-006-0222-4.

Herwartz:2006:EAH

[76] Helmut Herwartz. Econometric analysis of high frequency data. AStA. Advances in Statistical Analysis, 90(1):89– 104, March 2006. CODEN ???? ISSN 1863-8171 (print), 1863-818X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10182-006-
0223-3.

Fitzenberger:2006:UQR

[77] Bernd Fitzenberger and Ralf A. Wilke. Using quantile regression for duration analysis. AStA. Advances in Statistical Analysis, 90(1):105–120, March 2006. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-006-0224-2.

Hubler:2006:MNP

[78] Olaf Hübler. Multilevel and nonlinear panel data models. AStA. Advances in Statistical Analysis, 90(1):121–136, March 2006. CODEN ???? ISSN 1863-8171 (print), 1863-818X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10182-006-
0225-1.

Kauermann:2006:NMT

[79] Göran Kauermann. Nonparametric models and their estimation. AStA. Advances in Statistical Analysis, 90 (1):137–152, March 2006. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL http:

//link.springer.com/article/10.
1007/s10182-006-0226-0.

Ronning:2006:MMA

[80] Gerd Ronning. Microeconometric models and anonymized micro data. AStA. Advances in Statistical Analysis, 90(1):153–166, March 2006. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-006-0227-z.

Boes:2006:ORM

[81] Stefan Boes and Rainer Winkelmann. Ordered response models. AStA. Advances in Statistical Analysis, 90(1):167–181, March 2006. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-006-0228-y.

Schneeweiss:2006:SRA

[82] Hans Schneeweiss and Thomas Augustin. Some recent advances in measurement error models and methods. AStA. Advances in Statistical Analysis, 90(1):183–197, March 2006. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-006-0229-x.

Caliendo:2006:MET

[83] Marco Caliendo and Reinhard Hujer. The microeconometric estimation of treatment effects — an overview. AStA. Advances in Statistical Analysis, 90(1):199–215, March 2006. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-006-0230-4.

Rassler:2006:SIN

[84] Susanne Rässler and Regina T. Riphahn. Survey item nonresponse and its treatment. AStA. Advances in Statistical Analysis, 90(1):217–232, March 2006. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-006-0231-3.

Anonymous:2006:HCa

[85] Anonymous. Help & contacts. AStA. Advances in Statistical Analysis, 90(1): ??, March 2006. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

Krumbholz:2006:OCD

[86] Wolf Krumbholz and Andreas Rohr. The operating characteristic of double sampling plans by variables when the standard deviation is unknown. AStA. Advances in Statistical Analysis, 90(2):233–251, June 2006. CO-DEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-006-0232-2.

Boes:2006:USE

[87] Stephan Boes and Peter Pflaumer. University student enrollment forecasts by analysis structural ratios using ARIMAmethods. AStA. Advances in Statistical Analysis, 90(2):253–271, June 2006. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-006-0233-1.

Hardle:2006:AIV

[88] Wolfgang Härdle, Zdenek Hlávka, and Gerhard Stahl. On the appropriateness of inappropriate VaR models. AStA. Advances in Statistical Analysis, 90(2):273–297, June 2006. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-006-0234-0.

Hujer:2006:EVT

[89] Reinhard Hujer, Stephan L. Thomsen, and Christopher Zeiss. The effects of vocational training programmes on the duration of unemployment in Eastern Germany. AStA. Advances in Statistical Analysis, 90(2):299–321, June 2006. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-006-0235-z.

Kladroba:2006:LSV

[90] Andreas Kladroba. E-learning in der Statistik — Ein Vergleich verschiedener Lernsoftwareangebote. (German) [E-learning in statistics — a comparison of different learning-software deals]. AStA. Advances in Statistical Analysis, 90(2):323–340, June 2006. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-006-0236-y.

vonderLippe:2006:FSN

[91] Peter von der Lippe. Ist der Föderalismus in der Statistik noch zeitgemäß?. (German) [Is federalism in statistics still up to date?]. AStA. Advances in Statistical Analysis, 90(2): 341–355, June 2006. CODEN ???? ISSN 1863-8171 (print), 1863-818X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10182-006-
                                        0237-x.

Fickel:2006:B

[92] Norman Fickel, Götz Uebe, Karsten Webel, and Benedikt Köhler. Buchbesprechungen. AStA. Advances in Statistical Analysis, 90(2):357–363, June 2006. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-006-0238-9.

Anonymous:2006:HCb

[93] Anonymous. Help & contacts. AStA. Advances in Statistical Analysis, 90(2): ??, June 2006. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

Martin:2006:AVP

[94] Klaus Martin and Annette Böckenhoff. Analysis of variance of paired data without repetition of measurement. AStA. Advances in Statistical Analysis, 90(3):365–384, September 2006. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-006-0239-8.

Sahoo:2006:NIS

[95] L. N. Sahoo, G. N. Singh, and B. C. Das. A note on an IPPS sampling scheme. AStA. Advances in Statistical Analysis, 90(3):385–393, September 2006. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-006-0240-2.

Klein:2006:PKT

[96] Ingo Klein and Matthias Fischer. Power kurtosis transformations: Definition, properties and ordering. AStA. Advances in Statistical Analysis, 90 (3):395–401, September 2006. CO-DEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-006-0241-1.

Nadarjah:2006:EVM

[97] Saralees Nadarjah. Extreme value models with application to drought data. AStA. Advances in Statistical Analysis, 90(3):403–418, September 2006. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-006-0242-0.

Schmid:2006:ELM

[98] Matthias Schmid. Estimation of a linear model under microaggregation by individual ranking. AStA. Advances in Statistical Analysis, 90(3):419–438, September 2006. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-006-0243-z.

Rothe:2006:PPT

[99] Christoph Rothe and Philipp Sibbertsen. Phillips–Perron-type unit root tests in the nonlinear ESTAR framework. AStA. Advances in Statistical Analysis, 90(3):439–456, September 2006. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-006-0244-y.

Todter:2006:VWV

[100] Karl-Heinz Tödter. Volumenanteile und Wachstumsbeiträge bei der Vorjahrespreismethode mit Verkettung. (German) [Volume fractions and contributions to growth in the previous year’s price-method linkage]. AStA. Advances in Statistical Analysis, 90(3):457–464, September 2006. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-006-0245-x.

Eppmann:2006:FGR

[101] Helmut Eppmann, Sonja Krügener, and Josef Schäfer. First German register based census in 2011. AStA. Advances in Statistical Analysis, 90(3):465–482, September 2006. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-006-0246-9.

Uebe:2006:B

[102] Götz Uebe, Norman Fickel, Karsten Webel, Rainer Schlittgen, and Joachim Merz. Buchbesprechungen. AStA. Advances in Statistical Analysis, 90 (3):483–490, September 2006. CO-DEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-006-0247-8.

Anonymous:2006:HCc

[103] Anonymous. Help & contacts. AStA.

Advances in Statistical Analysis, 90 (3):??, September 2006. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

Rendtel:2006:PMM

[104] Ulrich Rendtel. The 2005 Plenary Meeting on “Missing Data and Measurement Error”. AStA. Advances in Statistical Analysis, 90(4):493–499, December 2006. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-006-0011-0.

Rubin:2006:CCI

[105] Donald B. Rubin. Conceptual, computational and inferential benefits of the missing data perspective in applied and theoretical statistical problems. AStA. Advances in Statistical Analysis, 90(4):501–513, December 2006. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-006-0004-z.

Raghunathan:2006:CIM

[106] Trivellore E. Raghunathan. Combining information from multiple surveys for assessing health disparities. AStA. Advances in Statistical Analysis, 90(4):515–526, December 2006. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-006-0003-0.

Rassler:2006:EMD

[107] Susanne Rässler. Der Einsatz von Missing Data Techniken in der Arbeitsmarktforschung des IAB. (German) [The use of missing data techniques in labor market research of the IAB]. AStA. Advances in Statistical Analysis, 90(4):527–552, December 2006. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-006-0010-1.

Radermacher:2006:FFD

[108] Walter Radermacher and Thomas Körner. Fehlende und fehlerhafte Daten in der amtlichen Statistik. Neue Herausforderungen und Lösungsansätze. (German) [Missing and erroneous data in official statistics. New challenges and solutions]. AStA. Advances in Statistical Analysis, 90(4):553–576, December 2006. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-006-0001-2.

Durrant:2006:MDM

[109] Gabriele B. Durrant. Missing data methods in official statistics in the United Kingdom: Some recent developments. AStA. Advances in Statistical Analysis, 90(4):577–593, December 2006. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-006-0002-1.

Buscher:2006:PED

[110] Herbert S. Buscher and Juliane Parys. Prekäre Einkommenslagen in Deutschland. Ein Ost–West-Vergleich 1996– 2002. (German) [Precarious income locations in Germany. an East–West comparison 1996–2002]. AStA. Advances in Statistical Analysis, 90(4):595–615, December 2006. CODEN ???? ISSN 1863-8171 (print), 1863-818X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10182-006-
0005-y.

Schmidt:2006:FFF

[111] Christoph M. Schmidt. Fokus, Fokus, Fokus? Zur künftigen Rolle der außeruniversitären Wirtschaftsforschungsinstitute. (German) [Focus, focus, focus? The future role of the non-university-sector research institutes]. AStA. Advances in Statistical Analysis, 90(4):617–622, December 2006. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-006-0012-z.

Fischer:2006:BSF

[112] Matthias Fischer. Buchbesprechung:

Schmid Friedrich, Mark Trede: Finanzmarktstatistik. AStA. Advances in Statistical Analysis, 90(4):623– 624, December 2006. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s10182-006-0006-x.

Webel:2006:BRK

[113] Karsten Webel. Book review: K.

Takezawa: Introduction to Nonparametric Regression. AStA. Advances in Statistical Analysis, 90(4): 625–626, December 2006. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s10182-006-0007-9.

Uebe:2006:BRP

[114] Götz Uebe. Book review: Pascal Ardilly, Yves Tillé: Sampling Methods: Exercises and Solutions. AStA. Advances in Statistical Analysis, 90(4): 627–628, December 2006. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/s10182-006-0008-8.

Todter:2006:BRM

[115] Karl-Heinz Tödter. Book review: Max C. Wewel, Statistik im Bachelor-Studium der BWL und VWL, Methoden, Anwendung und Interpretation. AStA. Advances in Statistical Analysis, 90(4):629–630, December 2006. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/s10182-006-0009-7.

Fischer:2006:BRZ

[116] Matthias Fischer. Book review: Zivot Eric and Jiahui Wang, Modeling Financial Time Series with S-PLUS. AStA. Advances in Statistical Analysis, 90(4): 631–632, December 2006. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/s10182-006-0013-y.

Anonymous:2006:HCd

[117] Anonymous. Help & contacts. AStA. Advances in Statistical Analysis, 90(4): ??, December 2006. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

Seidel:2007:E

[118] Wilfried Seidel. Editorial. AStA.

Advances in Statistical Analysis, 91 (1):1–2, March 2007. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s10182-007-0021-6.

Becker:2007:HTB

[119] Martin Becker, Ralph Friedmann, Stefan Klößner, and Walter Sanddorf-Köhle. A Hausman test for Brownian motion. AStA. Advances in Statistical Analysis, 91(1):3–21, March 2007. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-006-0019-5.

Fasso:2007:AMM

[120] Alessandro Fass‘o and Samuele Locatelli. Asymmetric monitoring of multivariate data with nonlinear dynamics. AStA. Advances in Statistical Analysis, 91(1):23–37, March 2007. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-006-0014-x.

Golosnoy:2007:SMM

[121] Vasyl Golosnoy. Sequential monitoring of minimum variance portfolio. AStA. Advances in Statistical Analysis, 91(1):39–55, March 2007. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-006-0016-8.

Koshevoy:2007:MLD

[122] Gleb A. Koshevoy and Karl Mosler. Multivariate Lorenz dominance based on zonoids. AStA. Advances in Statistical Analysis, 91(1):57–76, March 2007. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-006-0017-7.

Beer:2007:BHP

[123] Michael Beer. Bootstrapping a hedonic price index: experience from used cars data. AStA. Advances in Statistical Analysis, 91(1):77–92, March 2007. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-006-0015-9.

Nadarajah:2007:PSR

[124] Saralees Nadarajah and Samuel Kotz. Proportions, sums and ratios. AStA. Advances in Statistical Analysis, 91(1):93– 106, March 2007. CODEN ???? ISSN 1863-8171 (print), 1863-818X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10182-006-
                                        0018-6.

Uebe:2007:BRP

[125] Götz Uebe. Book review: Pascal Ardilly, Yves Tillé: Sampling Methods: Exercises and Solutions. AStA. Advances in Statistical Analysis, 91 (1):107–108, March 2007. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/s10182-007-0020-7.

Anonymous:2007:HCa

[126] Anonymous. Help & contacts. AStA.

Advances in Statistical Analysis, 91(1): ??, March 2007. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

Okhrin:2007:CDE

[127] Yarema Okhrin and Wolfgang Schmid.

Comparison of different estimation techniques for portfolio selection. AStA. Advances in Statistical Analysis, 91 (2):109–127, August 2007. CO-DEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-007-0026-1.

Niermann:2007:TLS

[128] Stefan Niermann. Testing for linearity in simple regression models. AStA. Advances in Statistical Analysis, 91(2):129–139, August 2007. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-007-0025-2.

Scharff:2007:IDR

[129] Juliane Scharff. Inflation and the divergence of relative prices: Evidence from a cointegration analysis. AStA. Advances in Statistical Analysis, 91(2):141–158, August 2007. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-007-0024-3.

Arntz:2007:ACA

[130] Melanie Arntz and Ralf A. Wilke.

An application of cartographic area interpolation to German administrative data. AStA. Advances in Statistical Analysis, 91(2):159–180, August 2007. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-007-0022-5.

Bomsdorf:2007:NAM

[131] Eckart Bomsdorf and Clemens Otto. A new approach to the measurement of polarization for grouped data. AStA. Advances in Statistical Analysis, 91(2):181–196, August 2007. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-007-0027-0.

Bachmaier:2007:CCO

[132] Martin Bachmaier. Consistency of completely outlier-adjusted simultaneous redescending M-estimators of location and scale. AStA. Advances in Statistical Analysis, 91(2):197–219, August 2007. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-007-0023-4.

Uebe:2007:BRW

[133] Götz Uebe. Book review: Wilkinson, L.: The Grammar of Graphics. AStA. Advances in Statistical Analysis, 91(2):221–222, August 2007. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s10182-007-0028-z.

Webel:2007:BRG

[134] Karsten Webel. Book review: G. Molenberghs and G. Verbeke: Models for Discrete Longitudinal Data. AStA. Advances in Statistical Analysis, 91 (2):223–224, August 2007. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s10182-007-0029-y.

Anonymous:2007:HCb

[135] Anonymous. Help & contacts. AStA.

Advances in Statistical Analysis, 91(2): ??, August 2007. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

Kneib:2007:SML

[136] Thomas Kneib, Bernhard Baumgartner, and Winfried J. Steiner. Semiparametric multinomial logit models for analysing consumer choice behaviour. AStA. Advances in Statistical Analysis, 91(3):225–244, October 2007. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-007-0033-2.

Saidane:2007:SVL

[137] Mohamed Saidane and Christian Lavergne. A structured variational learning approach for switching latent factor models. AStA. Advances in Statistical Analysis, 91(3):245–268, October 2007. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-007-0031-4.

Tomizawa:2007:DUA

[138] Sadao Tomizawa and Nobuko Miyamoto.

Diagonal uniform association symmetry models for cumulative probabilities in square tables. AStA. Advances in Statistical Analysis, 91(3):269–278, October 2007. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-007-0034-1.

Frolich:2007:IPS

[139] Markus Frölich. On the inefficiency of propensity score matching. AStA. Advances in Statistical Analysis, 91(3):279–290, October 2007. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-007-0035-0.

Martin:2007:AST

[140] Klaus Martin and Annette Böckenhoff. Analysis of short-term systematic measurement error variance for the difference of paired data without repetition of measurement. AStA. Advances in Statistical Analysis, 91(3):291–310, October 2007. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-007-0036-z.

Basic:2007:ABD

[141] Edin Basic and Ulrich Rendtel. Assessing the bias due to non-coverage of residential movers in the German Microcensus Panel: an evaluation using data from the Socio-Economic Panel. AStA. Advances in Statistical Analysis, 91(3):311–334, October 2007. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-007-0030-5.

Uebe:2007:BRV

[142] Götz Uebe. Book review: van der Hoek, J. and Elliott, R. J.: Binomial Models in Finance. AStA. Advances in Statistical Analysis, 91(3):335–337, October 2007. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-007-0037-y.

Weiss:2007:BRS

[143] Christian H. Weiß. Book review: Stat-Soft, Inc., Tulsa, OK: STATISTICA, Version 8. AStA. Advances in Statistical Analysis, 91(3):339–341, October 2007. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-007-0038-x.

Anonymous:2007:HCc

[144] Anonymous. Help & contacts. AStA.

Advances in Statistical Analysis, 91(3): ??, October 2007. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

Kauermann:2007:SC

[145] Göran Kauermann and Claus Weihs.

Statistical consulting. AStA. Advances in Statistical Analysis, 91(4):343–347, December 2007. CODEN ???? ISSN 1863-8171 (print), 1863-818X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10182-007-
0047-9.

Unwin:2007:SCI

[146] Antony Unwin. Statistical consulting interactions: a personal view. AStA. Advances in Statistical Analysis, 91(4):349–359, December 2007. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-007-0044-z.

Bowman:2007:IRI

[147] Adrian W. Bowman. Interdisciplinary research: the importance of learning other people’s language. AStA. Advances in Statistical Analysis, 91(4):361–365, December 2007. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-007-0039-9.

Windmann:2007:SCG

[148] Michael Windmann and Göran Kauermann. Statistical consulting at German universities: Results of a survey. AStA. Advances in Statistical Analysis, 91(4):367–378, December 2007. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-007-0045-y.

Schmidt:2007:PEE

[149] Christoph M. Schmidt. Policy evaluation and economic policy advice. AStA. Advances in Statistical Analysis, 91(4):379–389, December 2007. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-007-0040-3.

Wormer:2007:FSJ

[150] Holger Wormer. Figures, statistics and the journalist: an affair between love and fear. AStA. Advances in Statistical Analysis, 91(4):391–397, December 2007. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-007-0041-2.

Berding:2007:MLD

[151] Christoph Berding and Wilhelm Kleider. Medical laboratory diagnostics and statistics. AStA. Advances in Statistical Analysis, 91(4):399–406, December 2007. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-007-0043-0.

Ritter:2007:SUC

[152] Christian Ritter. Seeking useful contribution. AStA. Advances in Statistical Analysis, 91(4):407–411, December 2007. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-007-0042-1.

Engel:2007:SDS

[153] Jan Engel and Henriette (Jettie) C. M.

Hoonhout. Statistics development: statistical methods meeting the user’s needs. AStA. Advances in Statistical Analysis, 91(4):413–427, December 2007. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-007-0046-x.

Weihs:2007:QAS

[154] Claus Weihs. Quality assurance for statistical consulting. AStA. Advances in Statistical Analysis, 91(4):429–440, December 2007. CODEN ???? ISSN 1863-8171 (print), 1863-818X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10182-007-
0048-8.

Nadarajah:2007:MPP

[155] Saralees Nadarajah and Samuel Kotz.

Moments of a product Pearson type VII density distribution. AStA. Advances in Statistical Analysis, 91(4):441–448, December 2007. CODEN ???? ISSN 1863-8171 (print), 1863-818X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10182-007-
0032-3.

Schlittgen:2007:BRD

[156] Rainer Schlittgen. Book review: David Lemmon and Joseph Schafer: Developing Statistical Software in Fortran 95. AStA. Advances in Statistical Analysis, 91(4):449–450, December 2007. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s10182-007-0050-1.

Uebe:2007:BRS

[157] Goetz Uebe. Book review: Sam C. Saunders: Reliability, Life Testing and the Prediction of Service Lives for Engineers and Scientists. AStA. Advances in Statistical Analysis, 91(4): 451–452, December 2007. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s10182-007-0051-0.

Anonymous:2007:HCd

[158] Anonymous. Help & contacts. AStA. Advances in Statistical Analysis, 91(4): ??, December 2007. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

Varin:2008:CML

[159] Cristiano Varin. On composite marginal likelihoods. AStA. Advances in Statistical Analysis, 92(1):1–28, February 2008. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-008-0060-7.

Schmid:2008:EUE

[160] Wolfgang Schmid and Taras Zabolotskyy. On the existence of unbiased estimators for the portfolio weights obtained by maximizing the Sharpe ratio. AStA. Advances in Statistical Analysis, 92(1):29–34, February 2008. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-008-0054-5.

Wegener:2008:EHI

[161] Michael Wegener and Göran Kauermann. Examining heterogeneity in implied equity risk premium using penalized splines. AStA. Advances in Statistical Analysis, 92(1):35–56, February 2008. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-007-0052-z.

Holzmann:2008:LRT

[162] Hajo Holzmann and Sebastian Vollmer.

A likelihood ratio test for bimodality in two-component mixtures with application to regional income distribution in the EU. AStA. Advances in Statistical Analysis, 92(1):57–69, February 2008. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-008-0057-2.

Weiss:2008:MSD

[163] Christian H. Weiß and Rainer Göb.

Measuring serial dependence in categorical time series. AStA. Advances in Statistical Analysis, 92(1):71–89, February 2008. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-008-0055-4.

Demetrescu:2008:BCR

[164] Matei Demetrescu and Adina I. Tarcolea. Bias correction for the regressionbased LM fractional integration test. AStA. Advances in Statistical Analysis, 92(1):91–99, February 2008. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-008-0058-1.

Schunk:2008:MCM

[165] Daniel Schunk. A Markov chain Monte Carlo algorithm for multiple imputation in large surveys. AStA. Advances in Statistical Analysis, 92(1):101–114, February 2008. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-008-0053-6.

Schlittgen:2008:BRF

[166] Rainer Schlittgen. Book review: Forrest W. Young, Pedro M. Valero-Mora and Michael Friendly: Visual statistics: seeing data with dynamic interactive graphics. AStA. Advances in Statistical Analysis, 92(1):115, February 2008. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/s10182-008-0059-0.

Anonymous:2008:HCa

[167] Anonymous. Help & contacts. AStA. Advances in Statistical Analysis, 92(1): ??, February 2008. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

Grillenzoni:2008:RNE

[168] Carlo Grillenzoni. Robust nonparametric estimation of the intensity function of point data. AStA. Advances in Statistical Analysis, 92(2):117–134, May 2008. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-008-0065-2.

Eckel:2008:ISC

[169] S. Eckel, F. Fleischer, P. Grabarnik, and V. Schmidt. An investigation of the spatial correlations for relative purchasing power in Baden– Württemberg. AStA. Advances in Statistical Analysis, 92(2):135–152, May 2008. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-008-0066-1.

Kohler:2008:RBS

[170] Michael Kohler. A regression-based smoothing spline Monte Carlo algorithm for pricing American options in discrete time. AStA. Advances in Statistical Analysis, 92(2):153–178, May 2008. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-008-0067-0.

Singer:2008:GGH

[171] Hermann Singer. Generalized Gauss– Hermite filtering. AStA. Advances in Statistical Analysis, 92(2):179–195, May 2008. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-008-0068-z.

Bauer:2008:EBO

[172] Thomas K. Bauer and Mathias Sinning. An extension of the Blinder– Oaxaca decomposition to nonlinear models. AStA. Advances in Statistical Analysis, 92(2):197–206, May 2008. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-008-0056-3.

Kloberdanz:2008:PLM

[173] Kathrin Kloberdanz and Klaus D. Schmidt. Prediction in the linear model under a linear constraint. AStA. Advances in Statistical Analysis, 92(2): 207–215, May 2008. CODEN ???? ISSN 1863-8171 (print), 1863-818X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10182-008-
                                        0062-5.

Bukac:2008:CMD

[174] Josef Bukac. Comparison of measuring devices. AStA. Advances in Statistical Analysis, 92(2):217–227, May 2008. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-008-0063-4.

Qiang:2008:NDD

[175] Li Qiang and Wu Yi. A note on “Data depths satisfying the projection property”. AStA. Advances in Statistical Analysis, 92(2):229–232, May 2008. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-008-0061-6. See [13].

Schlittgen:2008:BRR

[176] Rainer Schlittgen. Book review: Robert H. Shumway and David S. Stoffer: Time series analysis and its applications with R examples, 2nd edn. AStA. Advances in Statistical Analysis, 92(2):233–234, May 2008. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s10182-008-0064-3.

Anonymous:2008:HCb

[177] Anonymous. Help & contacts. AStA.

Advances in Statistical Analysis, 92(2): ??, May 2008. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

Romanazzi:2008:NSD

[178] Mario Romanazzi. A note on simplicial depth function. AStA. Advances in Statistical Analysis, 92(3):235–253, August 2008. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-008-0073-2.

Tahata:2008:ODP

[179] Kouji Tahata and Sadao Tomizawa.

Orthogonal decomposition of pointsymmetry for multiway tables. AStA. Advances in Statistical Analysis, 92 (3):255–269, August 2008. CO-DEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-008-0070-5.

Hogrefe:2008:FDR

[180] Jens Hogrefe. Forecasting data revisions of GDP: a mixed frequency approach. AStA. Advances in Statistical Analysis, 92(3):271–296, August 2008. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-008-0071-4.

Pyy-Martikainen:2008:AII

[181] Marjo Pyy-Martikainen and Ulrich Rendtel. Assessing the impact of initial nonresponse and attrition in the analysis of unemployment duration with panel surveys. AStA. Advances in Statistical Analysis, 92(3):297–318, August 2008. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-008-0069-y.

Weiss:2008:TOM

[182] Christian H. Weiß. Thinning operations for modeling time series of counts — a survey. AStA. Advances in Statistical Analysis, 92(3):319–341, August 2008. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-008-0072-3.

Uebe:2008:BRJ

[183] Götz Uebe. Book review: J. E. Gentle: Matrix Algebra — Theory, Computations, and Applications in Statistics. AStA. Advances in Statistical Analysis, 92(3):343–344, August 2008. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s10182-008-0074-1.

Uebe:2008:BRB

[184] Götz Uebe. Book review: B. Thompson: The Nature of Statistical Evidence. AStA. Advances in Statistical Analysis, 92(3):345–347, August 2008. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-008-0075-0.

Stoimenova:2008:BRR

[185] Eugenia Stoimenova. Book review:

Rolf-Dieter Reiss and Michael Thomas: Statistical Analysis of Extreme Values. With Applications to Insurance, Finance, Hydrology and Other Fields. AStA. Advances in Statistical Analysis, 92(3):349–350, August 2008. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s10182-008-0076-z.

Anonymous:2008:HCc

[186] Anonymous. Help & contacts. AStA.

Advances in Statistical Analysis, 92(3): ??, August 2008. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

Pohlmeier:2008:MDC

[187] Winfried Pohlmeier and Gerd Ronning. Microeconometrics and disclosure control. AStA. Advances in Statistical Analysis, 92(4):351–357, December 2008. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-008-0093-y.

Schmid:2008:ELM

[188] Matthias Schmid and Hans Schneeweiss.

Estimation of a linear model in transformed variables under microaggregation by individual ranking. AStA. Advances in Statistical Analysis, 92 (4):359–374, December 2008. CO-DEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-008-0087-9.

Biewen:2008:PMN

[189] Elena Biewen, Sandra Nolte, and Martin Rosemann. Perturbation by multiplicative noise and the Simulation Extrapolation method. AStA. Advances in Statistical Analysis, 92(4):375–389, December 2008. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-008-0089-7.

Ronning:2008:SEC

[190] Gerd Ronning and Martin Rosemann. SIMEX estimation in case of correlated measurement errors. AStA. Advances in Statistical Analysis, 92 (4):391–404, December 2008. CO-DEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-008-0086-x.

Flossmann:2008:MAD

[191] Anton Flossmann and Sandra Nolte. Make assurance double sure: combination of two disclosure limitation methods and estimation of general regression models. AStA. Advances in Statistical Analysis, 92(4):405–422, December 2008. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-008-0094-x.

Biewen:2008:ELM

[192] Elena Biewen and Gerd Ronning. Estimation of linear models with anonymised panel data. AStA. Advances in Statistical Analysis, 92(4):423–438, December 2008. CODEN ???? ISSN 1863-8171 (print), 1863-818X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10182-008-
0088-8.

Drechsler:2008:NAD

[193] Jörg Drechsler, Agnes Dundler, Stefan Bender, Susanne Rässler, and Thomas Zwick. A new approach for disclosure control in the IAB establishment panelmultiple imputation for a better data access. AStA. Advances in Statistical Analysis, 92(4):439–458, December 2008. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-008-0090-1.

Anonymous:2008:HCd

[194] Anonymous. Help & contacts. AStA.

Advances in Statistical Analysis, 92(4): ??, December 2008. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

Bohning:2009:RDL

[195] Dankmar Böhning and Peter van der Heijden. Recent developments in life and social science applications of capture-recapture methods. AStA. Advances in Statistical Analysis, 93 (1):1–3, March 2009. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL http://link.

springer.com/content/pdf/10.1007/
s10182-008-0097-7.pdf.

vanderHeijden:2009:SMD

[196] Peter G. M. van der Heijden, Eugene Zwane, and David Hessen. Structurally missing data problems in multiple list capture-recapture data. AStA.

Advances in Statistical Analysis, 93 (1):5–21, March 2009. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL http://link.

                                        springer.com/content/pdf/10.1007/
                                        s10182-008-0098-6.pdf.

Brittain:2009:ECR

[197] Sarah Brittain and Dankmar Böhning. Estimators in capture-recapture studies with two sources. AStA. Advances in Statistical Analysis, 93(1):23–47, March 2009. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-008-0085-y.

Fienberg:2009:IMM

[198] Stephen E. Fienberg and Daniel Manrique-Vallier. Integrated methodology for multiple systems estimation and record linkage using a missing data formulation. AStA. Advances in Statistical Analysis, 93(1):49–60, March 2009. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-008-0084-z.

Kuhnert:2009:CCA

[199] Ronny Kuhnert and Dankmar Böhning. CAMCR: Computer-Assisted Mixture model analysis for Capture–Recapture count data. AStA. Advances in Statistical Analysis, 93(1):61–71, March 2009. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-008-0092-z.

Garel:2009:CPD

[200] Bernard Garel and Jean-Claude Massé. Calculation of the Prokhorov distance by optimal quantization and maximum flow. AStA. Advances in Statistical Analysis, 93(1):73–88, March 2009. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-008-0082-1.

Nadarajah:2009:LSD

[201] Saralees Nadarajah. The largest SNR distribution. AStA. Advances in Statistical Analysis, 93(1):89–107, March 2009. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-008-0080-3.

Bornkamp:2009:BRJ

[202] Björn Bornkamp. Book review: Jayanta K. Ghosh, Mohan Delampady and Tapas Samanta: An introduction to Bayesian analysis — theory and methods. AStA. Advances in Statistical Analysis, 93 (1):109–110, March 2009. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s10182-008-0100-3.

Uebe:2009:BRP

[203] Götz Uebe. Book review: Peter X.-K. Song: Correlated data analysis: modeling, analytics, and applications. AStA. Advances in Statistical Analysis, 93(1):111–113, March 2009. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-008-0095-9.

Kischka:2009:BRD

[204] Peter Kischka. Book review: Dennis V. Lindley: Understanding uncertainty.

AStA. Advances in Statistical Analysis, 93(1):115–116, March 2009. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s10182-008-0103-0.

Uebe:2009:BRY

[205] Götz Uebe. Book review: Yves Tillé: Sampling algorithms. AStA. Advances in Statistical Analysis, 93 (1):117–118, March 2009. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s10182-008-0104-z.

Haupt:2009:BRJ

[206] Harry Haupt. Book review: Jérôme Dedecker, Paul Doukhan, Gabriel Lang, José Rafael León R., Sana Louhichi and Clémentine Prieur: Weak dependence: with examples and applications. (Lecture notes in statistics). AStA. Advances in Statistical Analysis, 93 (1):119–120, March 2009. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s10182-008-0102-1.

Uebe:2009:BRC

[207] Götz Uebe. Book review: C. Radhakrishna Rao, Helge Toutenburg, Shalabh and Christian Heumann. Linear Models and Generalizations, Least Squares and Alternatives, 3rd edition. AStA. Advances in Statistical Analysis, 93 (1):121–122, March 2009. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s10182-008-0101-2.

Anonymous:2009:HCa

[208] Anonymous. Help & contacts. AStA.

Advances in Statistical Analysis, 93(1): ??, March 2009. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

Kopperschmidt:2009:PTM

[209] Kai Kopperschmidt and Winfried Stute.

Purchase timing models in marketing: a review. AStA. Advances in Statistical Analysis, 93(2):123–149, June 2009. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-008-0096-8.

Dhar:2009:CRE

[210] Subhra Sankar Dhar and Probal Chaudhuri. A comparison of robust estimators based on two types of trimming. AStA. Advances in Statistical Analysis, 93(2):151–158, June 2009. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-008-0099-5.

Meinel:2009:CPM

[211] Nina Meinel. Comparison of performance measures for multivariate discrete models. AStA. Advances in Statistical Analysis, 93(2):159–174, June 2009. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-008-0078-x.

DiZio:2009:SPM

[212] Marco Di Zio and Ugo Guarnera.

Semiparametric predictive mean matching. AStA. Advances in Statistical Analysis, 93(2):175–186, June 2009.

CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-008-0081-2.

Nadarajah:2009:SLD

[213] Saralees Nadarajah. The skew logistic distribution. AStA. Advances in Statistical Analysis, 93(2):187–203, June 2009. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-009-0105-6.

Kloberdanz:2009:LPL

[214] Kathrin Kloberdanz and Klaus D. Schmidt. Loss prediction in a linear model under a linear constraint. AStA. Advances in Statistical Analysis, 93(2):205–220, June 2009. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-008-0077-y.

Hess:2009:MSM

[215] Klaus T. Hess. Marginal-sum and maximum-likelihood estimation in a multiplicative tariff. AStA. Advances in Statistical Analysis, 93(2):221–233, June 2009. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-008-0079-9.

Anonymous:2009:HCb

[216] Anonymous. Help & contacts. AStA. Advances in Statistical Analysis, 93(2): ??, June 2009. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

Kauermann:2009:E

[217] Goeran Kauermann and Stefan Lang. Editorial. AStA. Advances in Statistical Analysis, 93(3):235–236, September 2009. CODEN ???? ISSN 1863-8171 (print), 1863-818X (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s10182-
009-0112-7.pdf.

Redenbach:2009:AAP

[218] Claudia Redenbach, Aila Särkkä, Johannes Freitag, and Katja Schladitz. Anisotropy analysis of pressed point processes. AStA. Advances in Statistical Analysis, 93(3):237–261, September 2009. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-009-0106-5.

Golosnoy:2009:MCC

[219] Vasyl Golosnoy, Sergiy Ragulin, and Wolfgang Schmid. Multivariate CUSUM chart: properties and enhancements. AStA. Advances in Statistical Analysis, 93(3):263–279, September 2009. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-009-0107-4.

Krumbholz:2009:DAM

[220] Wolf Krumbholz and Andreas Rohr.

Double ASN Minimax sampling plans by variables when the standard deviation is unknown. AStA. Advances in Statistical Analysis, 93(3):281–294, September 2009. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-009-0111-8.

Bodnar:2009:SIP

[221] Olha Bodnar and Taras Bodnar. Statistical inference procedure for the meanvariance efficient frontier with estimated parameters. AStA. Advances in Statistical Analysis, 93(3):295–306, September 2009. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-009-0109-2.

vanderLinde:2009:BLV

[222] Angelika van der Linde. A Bayesian latent variable approach to functional principal components analysis with binary and count data. AStA. Advances in Statistical Analysis, 93(3):307–333, September 2009. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-009-0113-6.

Aktas:2009:ESD

[223] Serpil Aktas and Tülay Saraçbasi. Estimation of symmetric disagreement using a uniform association model for ordinal agreement data. AStA. Advances in Statistical Analysis, 93(3):335–343, September 2009. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-008-0083-0.

Nadarajah:2009:LRV

[224] Saralees Nadarajah. Laplace random variables with application to price indices. AStA. Advances in Statistical Analysis, 93(3):345–369, September 2009. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-009-0108-3.

Anonymous:2009:HCc

[225] Anonymous. Help & contacts. AStA. Advances in Statistical Analysis, 93 (3):??, September 2009. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

Figueiredo:2009:MST

[226] Adelaide Figueiredo. Multi-sample tests for axial data from Watson distributions. AStA. Advances in Statistical Analysis, 93(4):371–386, December 2009. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-009-0114-5.

Giacomini:2009:DSF

[227] Enzo Giacomini, Wolfgang Härdle, and Volker Krätschmer. Dynamic semiparametric factor models in risk neutral density estimation. AStA. Advances in Statistical Analysis, 93(4):387–402, December 2009. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-009-0115-4.

Hahn:2009:EMB

[228] Markus Hahn, Sylvia Frühwirth-Schnatter, and Jörn Sass. Estimating models based on Markov jump processes given fragmented observation series. AStA. Advances in Statistical Analysis, 93(4):403–425, December 2009. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-009-0116-3.

Pannenberg:2009:GEC

[229] Markus Pannenberg and Martin Spiess.

GEE estimation of the covariance structure of a bivariate panel data model with an application to wage dynamics and the incidence of profit-sharing in West Germany. AStA. Advances in Statistical Analysis, 93(4):427–447, December 2009. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-009-0117-2.

Anonymous:2009:HCd

[230] Anonymous. Help & contacts. AStA. Advances in Statistical Analysis, 93(4): ??, December 2009. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

Hardle:2010:CNE

[231] Wolfgang Karl Härdle and Ostap Okhrin. De copulis non est disputandum. AStA. Advances in Statistical Analysis, 94(1):1–31, March 2010. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-009-0118-1.

Jensen:2010:EGO

[232] Uwe Jensen, Hermann Gartner, and Susanne Rässler. Estimating German overqualification with stochastic earnings frontiers. AStA. Advances in Statistical Analysis, 94(1):33–51, March 2010. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-009-0121-6.

Knuppel:2010:MSG

[233] Lothar Knüppel and Oliver Hermsen. Median split, k-group split, and optimality in continuous populations. AStA. Advances in Statistical Analysis, 94(1):53–74, March 2010. CO-DEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-010-0122-5.

Melo:2010:MSL

[234] Tatiane F. N. Melo and Silvia L. P.

Ferrari. A modified signed likelihood ratio test in elliptical structural models. AStA. Advances in Statistical Analysis, 94(1):75–87, March 2010. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-010-0123-4.

Fan:2010:APL

[235] Guo-Liang Fan, Han-Ying Liang, Jiang-Feng Wang, and Hong-Xia Xu. Asymptotic properties for LS estimators in EV regression model with dependent errors. AStA. Advances in Statistical Analysis, 94(1):89–103, March 2010. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-010-0124-3.

Fink:2010:MCV

[236] Dietmar Fink and Oded Löwenbein.

A management consulting view on the statistical consulting process. AStA. Advances in Statistical Analysis, 94 (1):105–109, March 2010. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL http:

//link.springer.com/article/10.
1007/s10182-009-0119-0.

Fink:2010:OOL

[237] Dietmar Fink and Marc Ant. Obituary: Oded Löwenbein (1953–2009). AStA. Advances in Statistical Analysis, 94(1):111, March 2010. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL http://link.

                                        springer.com/content/pdf/10.1007/
                                        s10182-009-0120-7.pdf.

Anonymous:2010:HCa

[238] Anonymous. Help & contacts. AStA. Advances in Statistical Analysis, 94(1): ??, March 2010. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

Betzin:2010:ASE

[239] Jörg Betzin and Dirk Temme. Advances in structural equation modeling — the 2009 meeting of the working group SEM. AStA. Advances in Statistical Analysis, 94(2):113–116, June 2010. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL http://link.

                                        springer.com/content/pdf/10.1007/
                                        s10182-010-0133-2.pdf.

Barendse:2010:URF

[240] M. T. Barendse, F. J. Oort, and G. J. A. Garst. Using restricted factor analysis with latent moderated structures to detect uniform and nonuniform measurement bias; a simulation study. AStA. Advances in Statistical Analysis, 94(2):117–127, June 2010. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL http://link.

                                        springer.com/content/pdf/10.1007/
                                        s10182-010-0126-1.pdf.

Jak:2010:MBM

[241] Suzanne Jak, Frans J. Oort, and Conor V. Dolan. Measurement bias and multidimensionality; an illustration of bias detection in multidimensional measurement models. AStA. Advances in Statistical Analysis, 94 (2):129–137, June 2010. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL http://link.

springer.com/content/pdf/10.1007/
s10182-010-0128-z.pdf.

King-Kallimanis:2010:USE

[242] B. L. King-Kallimanis, F. J. Oort, and G. J. A. Garst. Using structural equation modelling to detect measurement bias and response shift in longitudinal data. AStA. Advances in Statistical Analysis, 94(2):139–156, June 2010. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL http://

link.springer.com/content/pdf/10.
1007/s10182-010-0129-y.pdf.

Klein:2010:INM

[243] Andreas G. Klein and Karin Schermelleh-Engel. Introduction of a new measure for detecting poor fit due to omitted nonlinear terms in SEM. AStA. Advances in Statistical Analysis, 94(2):157–166, June 2010. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-010-0130-5.

Schermelleh-Engel:2010:NSE

[244] Karin Schermelleh-Engel, Christina S.

Werner, Andreas G. Klein, and Helfried Moosbrugger. Nonlinear structural equation modeling: is partial least squares an alternative? AStA. Advances in Statistical Analysis, 94(2): 167–184, June 2010. CODEN ???? ISSN 1863-8171 (print), 1863-818X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10182-010-
0132-3.

Geiser:2010:MMC

[245] Christian Geiser, Michael Eid, Fridtjof W.

Nussbeck, Delphine S. Courvoisier, and David A. Cole. Multitrait-multimethod change modelling. AStA. Advances in Statistical Analysis, 94(2):185–201, June 2010. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-010-0127-0.

Oud:2010:SOS

[246] J. H. L. Oud. Second-order stochastic differential equation model as an alternative for the ALT and CALT models. AStA. Advances in Statistical Analysis, 94(2):203–215, June 2010. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL http://link.

                                        springer.com/content/pdf/10.1007/
                                        s10182-010-0131-4.pdf.

Anonymous:2010:HCb

[247] Anonymous. Help & contacts. AStA. Advances in Statistical Analysis, 94(2): ??, June 2010. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

Freeland:2010:TIV

[248] R. Keith Freeland. True integer value time series. AStA. Advances in Statistical Analysis, 94(3):217–229, September 2010. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-010-0135-0.

Nadarajah:2010:DPE

[249] Saralees Nadarajah. Distribution properties and estimation of the ratio of independent Weibull random variables. AStA. Advances in Statistical Analysis, 94(3):231–246, September 2010. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-010-0134-1.

Schneeweiss:2010:SAR

[250] H. Schneeweiss, J. Komlos, and A. S.

Ahmad. Symmetric and asymmetric rounding: a review and some new results. AStA. Advances in Statistical Analysis, 94(3):247–271, September 2010. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-010-0125-2.

Wagner:2010:CAS

[251] Martin Wagner. Cointegration analysis with state space models. AStA. Advances in Statistical Analysis, 94 (3):273–305, September 2010. CO-DEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-010-0138-x.

Anonymous:2010:HCc

[252] Anonymous. Help & contacts. AStA.

Advances in Statistical Analysis, 94 (3):??, September 2010. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

Kuhnt:2010:DAC

[253] Sonja Kuhnt and David M. Steinberg. Design and analysis of computer experiments. AStA. Advances in Statistical Analysis, 94(4):307– 309, December 2010. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL http://link.

springer.com/content/pdf/10.1007/
s10182-010-0143-0.pdf.

Levy:2010:CER

[254] Sigal Levy and David M. Steinberg. Computer experiments: a review. AStA. Advances in Statistical Analysis, 94(4):311–324, December 2010. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-010-0147-9.

Petelet:2010:LHS

[255] Matthieu Petelet, Bertrand Iooss, Olivier Asserin, and Alexandre Loredo. Latin hypercube sampling with inequality constraints. AStA. Advances in Statistical Analysis, 94(4):325–339, December 2010. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-010-0144-z.

Jourdan:2010:OLH

[256] A. Jourdan and J. Franco. Optimal latin hypercube designs for the Kullback– Leibler criterion. AStA. Advances in Statistical Analysis, 94(4):341–351, December 2010. CODEN ???? ISSN 1863-8171 (print), 1863-818X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10182-010-
0145-y.

Pistone:2010:CGL

[257] Giovanni Pistone and Grazia Vicario. Comparing and generating latin hypercube designs in kriging models. AStA. Advances in Statistical Analysis, 94(4):353–366, December 2010. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-010-0142-1.

Ratto:2010:URA

[258] Marco Ratto and Andrea Pagano. Using recursive algorithms for the efficient identification of smoothing spline ANOVA models. AStA. Advances in Statistical Analysis, 94(4):367–388, December 2010. CODEN ???? ISSN 1863-8171 (print), 1863-818X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10182-010-
0148-8.

Wagner:2010:MTC

[259] Tobias Wagner, Christoph Bröcker, Nicolas Saba, Dirk Biermann, Anton Matzenmiller, and Kurt Steinhoff. Modelling of a thermomechanically coupled forming process based on functional outputs from a finite element analysis and from experimental measurements. AStA. Advances in Statistical Analysis, 94(4):389–404, December 2010. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-010-0149-7.

Lefebvre:2010:MAI

[260] Sidonie Lefebvre, Antoine Roblin, Suzanne Varet, and Gérard Durand. Metamodeling of aircraft infrared signature dispersion. AStA. Advances in Statistical Analysis, 94(4):405–422, December 2010. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-010-0146-x.

Anonymous:2010:HCd

[261] Anonymous. Help & contacts. AStA.

Advances in Statistical Analysis, 94(4): ??, December 2010. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

Drechsler:2011:MIP

[262] Jörg Drechsler. Multiple imputation in practice — a case study using a complex German establishment survey. AStA. Advances in Statistical Analysis, 95(1):1–26, March 2011. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-010-0136-z.

Schulz:2011:EIT

[263] Frowin C. Schulz and Karl Mosler. The effect of infrequent trading on detecting price jumps. AStA. Advances in Statistical Analysis, 95(1):27–58, March 2011. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-010-0137-y.

Jung:2011:UMT

[264] Robert C. Jung and A. R. Tremayne. Useful models for time series of counts or simply wrong ones? AStA. Advances in Statistical Analysis, 95(1):59–91, March 2011. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-010-0139-9.

Lin:2011:RTN

[265] Hsin-Yi Lin. A robust test for nonnested hypotheses. AStA. Advances in Statistical Analysis, 95(1):93–111, March 2011. CODEN ???? ISSN 1863-8171 (print), 1863-818X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10182-010-
                                        0140-3.

Anonymous:2011:HCa

[266] Anonymous. Help & contacts. AStA. Advances in Statistical Analysis, 95(1): ??, March 2011. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

Scagliarini:2011:MPC

[267] Michele Scagliarini. Multivariate process capability using principal component analysis in the presence of measurement errors. AStA. Advances in Statistical Analysis, 95(2):113–128, June 2011. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-011-0156-3.

Cribari-Neto:2011:NHC

[268] Francisco Cribari-Neto and Wilton Bernardino da Silva. A new heteroskedasticityconsistent covariance matrix estimator for the linear regression model. AStA. Advances in Statistical Analysis, 95(2):129–146, June 2011. CO-DEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-010-0141-2.

Herwartz:2011:SGP

[269] Helmut Herwartz. Specific-to-general predictor selection in approximate autoregressions-Monte Carlo evidence and a large scale performance assessment with real data. AStA. Advances in Statistical Analysis, 95(2):147–168, June 2011. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-010-0150-1.

Kundu:2011:ACB

[270] Debasis Kundu and Rameshwar D.

Gupta. Absolute continuous bivariate generalized exponential distribution. AStA. Advances in Statistical Analysis, 95(2):169–185, June 2011. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-010-0151-0.

Hassler:2011:ANT

[271] Uwe Hassler, Matei Demetrescu, and Adina I. Tarcolea. Asymptotic normal tests for integration in panels with crossdependent units. AStA. Advances in Statistical Analysis, 95(2):187–204, June 2011. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-011-0153-6.

Schlittgen:2011:WLS

[272] Rainer Schlittgen. A weighted leastsquares approach to clusterwise regression. AStA. Advances in Statistical Analysis, 95(2):205–217, June 2011. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-011-0155-4.

Anonymous:2011:HCb

[273] Anonymous. Help & contacts. AStA. Advances in Statistical Analysis, 95(2): ??, June 2011. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

Nadarajah:2011:EED

[274] Saralees Nadarajah. The exponentiated exponential distribution: a survey. AStA. Advances in Statistical Analysis, 95(3):219–251, September 2011. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-011-0154-5.

Veraart:2011:HPF

[275] Almut E. D. Veraart. How precise is the finite sample approximation of the asymptotic distribution of realised variation measures in the presence of jumps? AStA. Advances in Statistical Analysis, 95(3):253–291, September 2011. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-011-0158-1.

Colombi:2011:TLM

[276] Roberto Colombi and Sabrina Giordano.

Testing lumpability for marginal discrete hidden Markov models. AStA. Advances in Statistical Analysis, 95 (3):293–311, September 2011. CO-DEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-011-0160-7.

Sherman:2011:IRF

[277] Michael Sherman, Arnab Maity, and Suojin Wang. Inferences for the ratio: Fieller’s interval, log ratio, and large sample based confidence intervals. AStA. Advances in Statistical Analysis, 95(3):313–323, September 2011. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-011-0162-5.

Anonymous:2011:HCc

[278] Anonymous. Help & contacts. AStA.

Advances in Statistical Analysis, 95 (3):??, September 2011. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

Haupt:2011:ISI

[279] Harry Haupt and Cheng Hsiao. Introduction to the special issue: interdisciplinary aspects of panel data analysis. AStA. Advances in Statistical Analysis, 95(4):325–327, December 2011. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-011-0167-0.

Baltagi:2011:THP

[280] Badi H. Baltagi, Chihwa Kao, and Sanggon Na. Test of hypotheses in panel data models when the regressor and disturbances are possibly nonstationary. AStA. Advances in Statistical Analysis, 95(4):329–350, December 2011. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-011-0170-5.

Kleinke:2011:EWI

[281] Kristian Kleinke, Mark Stemmler, Jost Reinecke, and Friedrich Lösel. Efficient ways to impute incomplete panel data. AStA. Advances in Statistical Analysis, 95(4):351–373, December 2011. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-011-0179-9.

Singer:2011:CDS

[282] Hermann Singer. Continuous-discrete state-space modeling of panel data with nonlinear filter algorithms. AStA. Advances in Statistical Analysis, 95 (4):375–413, December 2011. CO-DEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-011-0172-3.

Reinecke:2011:GMM

[283] Jost Reinecke and Daniel Seddig. Growth mixture models in longitudinal research. AStA. Advances in Statistical Analysis, 95(4):415–434, December 2011. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-011-0171-4.

Bresson:2011:ACR

[284] Georges Bresson, Cheng Hsiao, and Alain Pirotte. Assessing the contribution of R&D to total factor productivity — a Bayesian approach to account for heterogeneity and heteroskedasticity. AStA. Advances in Statistical Analysis, 95(4):435–452, December 2011. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-011-0169-y.

McArdle:2011:LDA

[285] John J. McArdle. Longitudinal dynamic analyses of cognition in the health and retirement study panel. AStA. Advances in Statistical Analysis, 95(4):453–480, December 2011. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-011-0168-z.

Davidov:2011:LCG

[286] Eldad Davidov, Stefan Thörner, Peter Schmidt, Stefanie Gosen, and Carina Wolf. Level and change of groupfocused enmity in Germany: unconditional and conditional latent growth curve models with four panel waves. AStA. Advances in Statistical Analysis, 95(4):481–500, December 2011. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-011-0174-1.

Bresson:2011:FCA

[287] Georges Bresson and Cheng Hsiao. A functional connectivity approach for modeling cross-sectional dependence with an application to the estimation of hedonic housing prices in Paris. AStA. Advances in Statistical Analysis, 95(4):501–529, December 2011. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-011-0173-2.

Anonymous:2011:HCd

[288] Anonymous. Help & contacts. AStA. Advances in Statistical Analysis, 95(4): ??, December 2011. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

Zuccolotto:2012:PCA

[289] Paola Zuccolotto. Principal component analysis with interval imputed missing values. AStA. Advances in Statistical Analysis, 96(1):1–23, January 2012. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-011-0164-3.

Chesneau:2012:AWE

[290] Christophe Chesneau and Jalal Fadili. Adaptive wavelet estimation of a function in an indirect regression model. AStA. Advances in Statistical Analysis, 96(1):25–46, January 2012. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-011-0157-2.

Heinzl:2012:AMM

[291] Felix Heinzl, Ludwig Fahrmeir, and Thomas Kneib. Additive mixed models with Dirichlet process mixture and P -spline priors. AStA. Advances in Statistical Analysis, 96(1):47–68, January 2012. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-011-0161-6.

Adhya:2012:IFP

[292] Sumanta Adhya, Tathagata Banerjee, and Gaurangadeb Chattopadhyay. Inference on finite population categorical response: nonparametric regression-based predictive approach. AStA. Advances in Statistical Analysis, 96(1):69–98, January 2012. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-011-0159-0.

Robinzonov:2012:BTN

[293] Nikolay Robinzonov, Gerhard Tutz, and Torsten Hothorn. Boosting techniques for nonlinear time series models. AStA. Advances in Statistical Analysis, 96(1):99–122, January 2012. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-011-0163-4.

Anonymous:2012:HCa

[294] Anonymous. Help & contacts. AStA.

Advances in Statistical Analysis, 96(1): ??, January 2012. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

Denuit:2012:SIA

[295] Michel Denuit. Special issue on Actuarial Statistics. AStA. Advances in Statistical Analysis, 96(2):123–125, June 2012. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL http://

                                        link.springer.com/content/pdf/10.
                                        1007/s10182-012-0195-4.pdf.

Olivieri:2012:LTA

[296] Annamaria Olivieri and Ermanno Pitacco. Life tables in actuarial models: from the deterministic setting to a Bayesian approach. AStA. Advances in Statistical Analysis, 96(2):127–153, June 2012. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-011-0177-y.

Christiansen:2012:MMH

[297] Marcus C. Christiansen. Multistate models in health insurance. AStA. Advances in Statistical Analysis, 96(2): 155–186, June 2012. CODEN ???? ISSN 1863-8171 (print), 1863-818X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10182-012-
                                        0189-2.

Antonio:2012:SCP

[298] Katrien Antonio and Emiliano A. Valdez. Statistical concepts of a priori and a posteriori risk classification in insurance. AStA. Advances in Statistical Analysis, 96(2):187–224, June 2012. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL http://

                                        link.springer.com/content/pdf/10.
                                        1007/s10182-011-0152-7.pdf.

Drees:2012:EVA

[299] Holger Drees. Extreme value analysis of actuarial risks: estimation and model validation. AStA. Advances in Statistical Analysis, 96(2):225–264, June 2012. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-011-0184-z.

Schmidt:2012:LPB

[300] Klaus D. Schmidt. Loss prediction based on run-off triangles. AStA. Advances in Statistical Analysis, 96(2):265–310, June 2012. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-011-0183-0.

Hess:2012:MLM

[301] Klaus T. Hess. Maximum-likelihood and marginal-sum estimation in some particular collective models. AStA. Advances in Statistical Analysis, 96(2): 311–326, June 2012. CODEN ???? ISSN 1863-8171 (print), 1863-818X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10182-011-
0178-x.

Anonymous:2012:HCb

[302] Anonymous. Help & contacts. AStA.

Advances in Statistical Analysis, 96(2): ??, June 2012. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

Durante:2012:SCP

[303] Fabrizio Durante and Roberto Ghiselli-Ricci. Supermigrative copulas and positive dependence. AStA. Advances in Statistical Analysis, 96(3):327–342, July 2012. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-011-0165-2.

Frahm:2012:MTP

[304] Gabriel Frahm, Tobias Wickern, and Christof Wiechers. Multiple tests for the performance of different investment strategies. AStA. Advances in Statistical Analysis, 96(3):343–383, July 2012. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-011-0166-1.

Janczura:2012:EEM

[305] Joanna Janczura and Rafal Weron. Efficient estimation of Markov regimeswitching models: an application to electricity spot prices. AStA. Advances in Statistical Analysis, 96(3): 385–407, July 2012. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL http://link.

springer.com/content/pdf/10.1007/
s10182-011-0181-2.pdf.

Cordeiro:2012:MED

[306] Gauss M. Cordeiro, Elizabeth M. Hashimoto, Edwin M. M. Ortega, and Marcelino A. R. Pascoa. The McDonald extended distribution: properties and applications. AStA. Advances in Statistical Analysis, 96(3):409–433, July 2012. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-011-0180-3.

Munnich:2012:NSO

[307] Ralf T. Münnich, Ekkehard W. Sachs, and Matthias Wagner. Numerical solution of optimal allocation problems in stratified sampling under box constraints. AStA. Advances in Statistical Analysis, 96(3):435–450, July 2012. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-011-0176-z.

Anonymous:2012:HCc

[308] Anonymous. Help & contacts. AStA.

Advances in Statistical Analysis, 96(3): ??, July 2012. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

Ouassou:2012:ROE

[309] Idir Ouassou and Mustapha Rachdi. Regression operator estimation by deltasequences method for functional data and its applications. AStA. Advances in Statistical Analysis, 96(4):451–465, October 2012. CODEN ???? ISSN 1863-8171 (print), 1863-818X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10182-011-
0175-0.

Kuhnert:2012:ESC

[310] Ronny Kuhnert, Martin Schlaud, and Hartmut Hecker. Evaluation strategies for case series: is Cox regression an alternative to the self controlled case series method for terminal events? AStA. Advances in Statistical Analysis, 96(4):467–492, October 2012. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-011-0187-9.

Furno:2012:TSB

[311] Marilena Furno. Tests for structural break in quantile regressions. AStA. Advances in Statistical Analysis, 96(4):493–515, October 2012.

CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-012-0188-3.

Guo:2012:SCB

[312] Mengmeng Guo and Wolfgang Karl Härdle. Simultaneous confidence bands for expectile functions. AStA. Advances in Statistical Analysis, 96(4):517–541, October 2012. CODEN ???? ISSN 1863-8171 (print), 1863-818X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10182-011-
                                        0182-1.

Anonymous:2012:HCd

[313] Anonymous. Help & contacts. AStA. Advances in Statistical Analysis, 96(4): ??, October 2012. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

Wilrich:2013:CVM

[314] Peter-T. Wilrich. Critical values of Mandel’s h and k, the Grubbs and the Cochran test statistic. AStA. Advances in Statistical Analysis, 97(1):1– 10, January 2013. CODEN ???? ISSN 1863-8171 (print), 1863-818X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10182-011-
                                        0185-y.

Hansen:2013:FPM

[315] Hendrik Hansen. The forecasting performance of mortality models. AStA. Advances in Statistical Analysis, 97(1):11–31, January 2013. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-011-0186-x.

Weissbach:2013:NNH

[316] Rafael Weißbach, Wladislaw Poniatowski, and Walter Krämer. Nearest neighbor hazard estimation with lefttruncated duration data. AStA. Advances in Statistical Analysis, 97(1):33– 47, January 2013. CODEN ???? ISSN 1863-8171 (print), 1863-818X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10182-012-
0194-5.

Ziegelmeyer:2013:IUE

[317] Michael Ziegelmeyer. Illuminate the unknown: evaluation of imputation procedures based on the SAVE survey. AStA. Advances in Statistical Analysis, 97(1):49–76, January 2013. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-012-0197-2.

Schnabel:2013:SEQ

[318] Sabine K. Schnabel and Paul H. C.

Eilers. Simultaneous estimation of quantile curves using quantile sheets. AStA. Advances in Statistical Analysis, 97(1):77–87, January 2013. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL http://link.

springer.com/content/pdf/10.1007/
s10182-012-0198-1.pdf.

Anonymous:2013:HCa

[319] Anonymous. Help & contacts. AStA.

Advances in Statistical Analysis, 97(1): ??, January 2013. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

Fasso:2013:SSE

[320] Alessandro Fassò, Alessio Pollice, and Barbara Cafarelli. Spatial statistics for environmental studies. AStA. Advances in Statistical Analysis, 97(2): 89–91, April 2013. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL http://link.

springer.com/content/pdf/10.1007/
s10182-013-0209-x.pdf.

Bruno:2013:PAA

[321] Francesca Bruno, Daniela Cocchi, and Lucia Paci. A practical approach for assessing the effect of grouping in hierarchical spatio-temporal models. AStA. Advances in Statistical Analysis, 97(2):93–108, April 2013. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-012-0193-6.

Cameletti:2013:STM

[322] Michela Cameletti, Finn Lindgren, Daniel Simpson, and Håvard Rue. Spatio-temporal modeling of particulate matter concentration through the SPDE approach. AStA. Advances in Statistical Analysis, 97(2):109–131, April 2013. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-012-0196-3.

DeIaco:2013:PPP

[323] S. De Iaco, M. Palma, and D. Posa. Prediction of particle pollution through spatio-temporal multivariate geostatistical analysis: spatial special issue. AStA. Advances in Statistical Analysis, 97(2):133–150, April 2013. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-012-0199-0.

Villalta:2013:BST

[324] Desirée Villalta, Lelys Guenni, Yasmin Rubio-Palis, and Raúl Ramírez Arbeláez. Bayesian space-time modeling of malaria incidence in Sucre state, Venezuela. AStA. Advances in Statistical Analysis, 97(2):151–171, April 2013. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-012-0190-9.

Dawid:2013:ESP

[325] A. Philip Dawid and Monica Musio. Estimation of spatial processes using local scoring rules. AStA. Advances in Statistical Analysis, 97(2):173–179, April 2013. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-012-0191-8.

Bocci:2013:GME

[326] Chiara Bocci, Enrica Caporali, and Alessandra Petrucci. Geoadditive modeling for extreme rainfall data. AStA. Advances in Statistical Analysis, 97(2):181–193, April 2013. CO-DEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-012-0192-7.

Minozzo:2013:MGM

[327] Marco Minozzo and Clarissa Ferrari.

Multivariate geostatistical mapping of radioactive contamination in the Maddalena Archipelago (Sardinia, Italy): spatial special issue. AStA. Advances in Statistical Analysis, 97(2):195–213, April 2013. CODEN ???? ISSN 1863-8171 (print), 1863-818X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10182-012-
                                        0201-x.

Anonymous:2013:HCb

[328] Anonymous. Help & contacts. AStA. Advances in Statistical Analysis, 97(2): ??, April 2013. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

Na:2013:CPD

[329] Okyoung Na, Jiyeon Lee, and Sangyeol Lee. Change point detection in SCOMDY models. AStA. Advances in Statistical Analysis, 97(3):215–238, July 2013. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-012-0200-y.

Janczura:2013:GFT

[330] Joanna Janczura and Rafal Weron. Goodness-of-fit testing for the marginal distribution of regime-switching models with an application to electricity spot prices. AStA. Advances in Statistical Analysis, 97(3):239–270, July 2013. CO-DEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL http://

                                        link.springer.com/content/pdf/10.
                                        1007/s10182-012-0202-9.pdf.

Werkmann:2013:PUR

[331] Verena Werkmann. Performance of unit root tests in unbalanced panels: experimental evidence. AStA. Advances in Statistical Analysis, 97(3):271–285, July 2013. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-012-0203-8.

Trenkler:2013:CVM

[332] Carsten Trenkler and Enzo Weber.

Codependent VAR models and the pseudo-structural form. AStA. Advances in Statistical Analysis, 97(3):287–295, July 2013. CODEN ???? ISSN 1863-8171 (print), 1863-818X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10182-012-
0204-7.

Pyy-Martikainen:2013:AEH

[333] Marjo Pyy-Martikainen. Approaches for event history analysis based on complex longitudinal survey data. AStA. Advances in Statistical Analysis, 97(3): 297–315, July 2013. CODEN ???? ISSN 1863-8171 (print), 1863-818X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10182-012-
0205-6.

Anonymous:2013:HCc

[334] Anonymous. Help & contacts. AStA.

Advances in Statistical Analysis, 97(3): ??, July 2013. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

Lv:2013:SEL

[335] Xiaofeng Lv and Rui Li. Smoothed empirical likelihood analysis of partially linear quantile regression models with missing response variables. AStA. Advances in Statistical Analysis, 97(4):317–347, October 2013. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-013-0210-4.

Scheipl:2013:PLB

[336] Fabian Scheipl, Thomas Kneib, and Ludwig Fahrmeir. Penalized likelihood and Bayesian function selection in regression models. AStA. Advances in Statistical Analysis, 97(4):349–385, October 2013. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-013-0211-3.

Greven:2013:LRT

[337] Sonja Greven and Ciprian M. Crainiceanu. On likelihood ratio testing for penalized splines. AStA. Advances in Statistical Analysis, 97(4):387–402, October 2013. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-013-0214-0.

Heidenreich:2013:BSK

[338] Nils-Bastian Heidenreich, Anja Schindler, and Stefan Sperlich. Bandwidth selection for kernel density estimation: a review of fully automatic selectors. AStA. Advances in Statistical Analysis, 97(4):403–433, October 2013. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-013-0216-y.

Anonymous:2013:HCd

[339] Anonymous. Help & contacts. AStA. Advances in Statistical Analysis, 97(4): ??, October 2013. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

Cagnone:2014:FMM

[340] Silvia Cagnone and Cinzia Viroli. A factor mixture model for analyzing heterogeneity and cognitive structure of dementia. AStA. Advances in Statistical Analysis, 98(1):1–20, January 2014. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-012-0206-5.

Dimitriou-Fakalou:2014:GPL

[341] Chrysoula Dimitriou-Fakalou. Gaussian pseudo-likelihood estimation for stationary processes on a lattice. AStA. Advances in Statistical Analysis, 98(1):21– 34, January 2014. CODEN ???? ISSN 1863-8171 (print), 1863-818X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10182-013-
0207-z.

Bagnato:2014:DSD

[342] Luca Bagnato, Lucio De Capitani, and Antonio Punzo. Detecting serial dependencies with the reproducibility probability autodependogram. AStA. Advances in Statistical Analysis, 98(1):35–61, January 2014. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-013-0208-y.

Hulliger:2014:RDF

[343] Beat Hulliger and Tobias Schoch. Robust, distribution-free inference for income share ratios under complex sampling. AStA. Advances in Statistical Analysis, 98(1):63–85, January 2014. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-013-0215-z.

Krumbholz:2014:EOS

[344] Wolf Krumbholz and Detlef Steuer. On exact and optimal single-sampling plans by variables. AStA. Advances in Statistical Analysis, 98(1):87–101, January 2014. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-013-0217-x.

Anonymous:2014:HCa

[345] Anonymous. Help & contacts. AStA. Advances in Statistical Analysis, 98(1): ??, January 2014. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

Sun:2014:SOP

[346] Yuqin Sun, Rong Ke, and Yongge Tian. Some overall properties of seemingly unrelated regression models. AStA. Advances in Statistical Analysis, 98(2):103–120, April 2014. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-013-0212-2.

Shang:2014:SFP

[347] Han Lin Shang. A survey of functional principal component analysis. AStA. Advances in Statistical Analysis, 98(2):121–142, April 2014. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-013-0213-1.

Monni:2014:BVS

[348] Stefano Monni. Bayesian variable selection for correlated covariates via colored cliques. AStA. Advances in Statistical Analysis, 98(2):143–163, April 2014. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-013-0218-9.

Zhang:2014:ANE

[349] Jing-Jing Zhang, Han-Ying Liang, and Amei Amei. Asymptotic normality of estimators in heteroscedastic errors-in-variables model. AStA. Advances in Statistical Analysis, 98(2):165–195, April 2014. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-013-0224-y.

Anonymous:2014:HCb

[350] Anonymous. Help & contacts. AStA.

Advances in Statistical Analysis, 98(2): ??, April 2014. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

Choe:2014:CCA

[351] Geon Ho Choe and Kyungsub Lee.

Conditional correlation in asset return and GARCH intensity model. AStA. Advances in Statistical Analysis, 98(3):197–224, July 2014. CO-DEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-013-0219-8.

Garthoff:2014:SSM

[352] Robert Garthoff, Iryna Okhrin, and Wolfgang Schmid. Statistical surveillance of the mean vector and the covariance matrix of nonlinear time series. AStA. Advances in Statistical Analysis, 98(3):225–255, July 2014. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-013-0220-2.

Attaoui:2014:SUC

[353] Said Attaoui. Strong uniform consistency rates and asymptotic normality of conditional density estimator in the single functional index modeling for time series data. AStA. Advances in Statistical Analysis, 98(3):257–286, July 2014. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-014-0227-3.

Karimnezhad:2014:RBM

[354] Ali Karimnezhad and Ahmad Parsian. Robust Bayesian methodology with applications in credibility premium derivation and future claim size prediction. AStA. Advances in Statistical Analysis, 98(3):287–303, July 2014. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-013-0222-0.

Anonymous:2014:HCc

[355] Anonymous. Help & contacts. AStA. Advances in Statistical Analysis, 98(3): ??, July 2014. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

Golubev:2014:TMP

[356] Yuri Golubev, Wolfgang K. Härdle, and Roman Timofeev. Testing monotonicity of pricing kernels. AStA. Advances in Statistical Analysis, 98(4):305–326, October 2014. CODEN ???? ISSN 1863-8171 (print), 1863-818X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10182-014-
                                        0225-5.

Gupta:2014:ADD

[357] Ramesh C. Gupta, S. Z. Sim, and S. H. Ong. Analysis of discrete data by Conway–Maxwell Poisson distribution. AStA. Advances in Statistical Analysis, 98(4):327–343, October 2014. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-014-0226-4.

Singer:2014:ISK

[358] Hermann Singer. Importance sampling for Kolmogorov backward equations. AStA. Advances in Statistical Analysis, 98(4):345–369, October 2014. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-013-0223-z.

Steinmetz:2014:ECA

[359] Sebastian Steinmetz. EWMA charts:

ARL considerations in case of changes in location and scale. AStA. Advances in Statistical Analysis, 98(4):371–387, October 2014. CODEN ???? ISSN 1863-8171 (print), 1863-818X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10182-013-
0221-1.

Anonymous:2014:HCd

[360] Anonymous. Help & contacts. AStA.

Advances in Statistical Analysis, 98(4): ??, October 2014. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

Muller:2015:EEV

[361] Klaus Müller and Wolf-Dieter Richter.

Exact extreme value, product, and ratio distributions under non-standard assumptions. AStA. Advances in Statistical Analysis, 99(1):1–30, January 2015. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-014-0228-2.

Kumar:2015:ZIL

[362] C. Satheesh Kumar and A. Riyaz. A zero-inflated logarithmic series distribution of order k and its applications. AStA. Advances in Statistical Analysis, 99(1):31–43, January 2015. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-014-0229-1.

Barabesi:2015:GIE

[363] Lucio Barabesi, Giancarlo Diana, and Pier Francesco Perri. Gini index estimation in randomized response surveys. AStA. Advances in Statistical Analysis, 99(1):45–62, January 2015. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-014-0230-8.

Liebscher:2015:EPR

[364] Steffen Liebscher and Thomas Kirschstein. Efficiency of the pMST and RDELA location and scatter estimators. AStA. Advances in Statistical Analysis, 99 (1):63–82, January 2015. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL http:

//link.springer.com/article/10.
1007/s10182-014-0231-7.

Melo:2015:DBB

[365] Oscar O. Melo, Carlos E. Melo, and Jorge Mateu. Distance-based beta regression for prediction of mutual funds. AStA. Advances in Statistical Analysis, 99(1):83–106, January 2015. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-014-0232-6.

Franke:2015:NEC

[366] Jürgen Franke, Peter Mwita, and Weining Wang. Nonparametric estimates for conditional quantiles of time series. AStA. Advances in Statistical Analysis, 99(1):107–130, January 2015. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-014-0234-4.

Dabo-Niang:2015:APK

[367] Sophie Dabo-Niang, Zoulikha Kaid, and Ali Laksaci. Asymptotic properties of the kernel estimate of spatial conditional mode when the regressor is functional. AStA. Advances in Statistical Analysis, 99(2):131–160, April 2015. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-014-0233-5.

Hanck:2015:NVR

[368] Christoph Hanck and Robert Czudaj. Nonstationary-volatility robust panel unit root tests and the great moderation. AStA. Advances in Statistical Analysis, 99(2):161–187, April 2015. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-014-0235-3.

Barreto-Souza:2015:SIP

[369] Wagner Barreto-Souza and Marcelo Bourguignon. A skew INAR(1) process on Z. AStA. Advances in Statistical Analysis, 99(2):189–208, April 2015. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-014-0236-2.

Tutz:2015:EOP

[370] Gerhard Tutz and Gunther Schauberger. Extended ordered paired comparison models with application to football data from German Bundesliga. AStA. Advances in Statistical Analysis, 99(2):209–227, April 2015. CO-DEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-014-0237-1.

Withers:2015:CRV

[371] Christopher S. Withers and Saralees Nadarajah. Cumulants of a random variable distributed uniformly on the first n integers. AStA. Advances in Statistical Analysis, 99(2):229–236, April 2015. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-014-0238-0.

Grand:2015:MMD

[372] Alexandra Grand, Regina Dittrich, and Brian Francis. Markov models of dependence in longitudinal paired comparisons: an application to course design. AStA. Advances in Statistical Analysis, 99(2):237–257, April 2015. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-014-0239-z.

Konrath:2015:BAF

[373] Susanne Konrath, Ludwig Fahrmeir, and Thomas Kneib. Bayesian accelerated failure time models based on penalized mixtures of Gaussians: regularization and variable selection. AStA. Advances in Statistical Analysis, 99(3):259–280, July 2015. CO-DEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-014-0240-6.

Stange:2015:UQF

[374] Jens Stange, Taras Bodnar, and Thorsten Dickhaus. Uncertainty quantification for the family-wise error rate in multivariate copula models. AStA. Advances in Statistical Analysis, 99(3):281–310, July 2015. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-014-0241-5.

Zhu:2015:IDL

[375] Fukang Zhu, Lei Shi, and Shuangzhe Liu. Influence diagnostics in loglinear integer-valued GARCH models. AStA. Advances in Statistical Analysis, 99(3):311–335, July 2015. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-014-0242-4.

Maiti:2015:CFS

[376] Raju Maiti and Atanu Biswas. Coherent forecasting for stationary time series of discrete data. AStA. Advances in Statistical Analysis, 99(3):337–365, July 2015. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-014-0243-3.

Draxler:2015:PFC

[377] Clemens Draxler and Johannes Zessin. The power function of conditional tests of the Rasch model. AStA. Advances in Statistical Analysis, 99(3):367–378, July 2015. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-015-0249-5.

Schwiebert:2015:SOS

[378] Jörg Schwiebert. Semiparametric onestep estimation of a sample selection model with endogenous covariates. AStA. Advances in Statistical Analysis, 99(4):379–402, October 2015. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-015-0245-9.

Noven:2015:LDR

[379] Ragnhild C. Noven, Almut E. D. Veraart, and Axel Gandy. A Lévy-driven rainfall model with applications to futures pricing. AStA. Advances in Statistical Analysis, 99(4):403–432, October 2015. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-015-0246-8.

Brix:2015:PBE

[380] Anne Floor Brix and Asger Lunde. Prediction-based estimating functions for stochastic volatility models with noisy data: comparison with a GMM alternative. AStA. Advances in Statistical Analysis, 99(4):433–465, October 2015. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-015-0248-6.

Romano:2015:PTC

[381] Elvira Romano, Jorge Mateu, and Ramon Giraldo. On the performance of two clustering methods for spatial functional data. AStA. Advances in Statistical Analysis, 99(4):467–492, October 2015. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-015-0253-9.

Gupta:2016:PFR

[382] Ramesh C. Gupta and Barry C.

Arnold. Preservation of failure rate function shape in weighted distributions. AStA. Advances in Statistical Analysis, 100(1):1–20, January 2016. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-015-0244-x.

Fan:2016:ELS

[383] Guo-Liang Fan, Hong-Xia Xu, and Zhen-Sheng Huang. Empirical likelihood for semivarying coefficient model with measurement error in the nonparametric part. AStA. Advances in Statistical Analysis, 100(1):21–41, January 2016. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-015-0247-7.

Jorgensen:2016:DDM

[384] Bent Jørgensen and Célestin C. Kokonendji. Discrete dispersion models and their Tweedie asymptotics. AStA. Advances in Statistical Analysis, 100(1):43–78, January 2016. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-015-0250-z.

Ferrari:2016:SSO

[385] Silvia L. P. Ferrari and Eliane C. Pinheiro. Small-sample one-sided testing in extreme value regression models. AStA. Advances in Statistical Analysis, 100(1):79–97, January 2016. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-015-0251-y.

Maji:2016:LSD

[386] Avijit Maji, Abhik Ghosh, and Ayanendranath Basu. The logarithmic super divergence and asymptotic inference properties. AStA. Advances in Statistical Analysis, 100(1):99–131, January 2016. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-015-0252-x.

Klima:2016:EVT

[387] André Klima, Paul W. Thurner, Christoph Molnar, Thomas Schlesinger, and Helmut Küchenhoff. Estimation of voter transitions based on ecological inference: an empirical assessment of different approaches. AStA. Advances in Statistical Analysis, 100(2):133–159, April 2016. CODEN ???? ISSN 1863-8171 (print), 1863-818X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10182-015-
0254-8.

Perez-de-la-Cruz:2016:DAG

[388] Gonzalo Perez de-la Cruz and Guillermina Eslava-Gomez. Discriminant analysis with Gaussian graphical tree models. AStA. Advances in Statistical Analysis, 100(2):161–187, April 2016. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-015-0256-6.

Richter:2016:RBS

[389] Scott J. Richter and Melinda H. Mc-Cann. Resampling-based simultaneous confidence intervals for location shift using medians. AStA. Advances in Statistical Analysis, 100(2):189–205, April 2016. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-015-0258-4.

Li:2016:ASM

[390] Zhiming Li, Zhidong Teng, Tianfang Zhang, and Runchu Zhang. Analysis on s^{n−m} designs with general minimum lower-order confounding. AStA. Advances in Statistical Analysis, 100(2):207–222, April 2016. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-015-0259-3.

Feddag:2016:PLE

[391] M.-L. Feddag. Pairwise likelihood estimation for the normal Ogive model with binary data. AStA. Advances in Statistical Analysis, 100(2):223–237, April 2016. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-015-0263-7.

Chee:2016:MCD

[392] Chew-Seng Chee. Modelling of count data using nonparametric mixtures. AStA. Advances in Statistical Analysis, 100(3):239–257, July 2016. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/s10182-015-0255-7.

Frasso:2016:PEI

[393] Gianluca Frasso, Jonathan Jaeger, and Philippe Lambert. Parameter estimation and inference in dynamic systems described by linear partial differential equations. AStA. Advances in Statistical Analysis, 100 (3):259–287, July 2016. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/s10182-015-0257-5.

Gross:2016:MBH

[394] Marcus Groß. Modeling body height in prehistory using a spatio-temporal Bayesian errors-in-variables model. AStA. Advances in Statistical Analysis, 100(3):289–311, July 2016. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/s10182-015-0260-x.

Dai:2016:LIA

[395] Xiaowen Dai, Libin Jin, Lei Shi, Cuiping Yang, and Shuangzhe Liu. Local influence analysis in general spatial models. AStA. Advances in Statistical Analysis, 100(3):313–331, July 2016. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s10182-015-0261-9.

Schwiebert:2016:MCM

[396] Jörg Schwiebert. Multinomial choice models based on Archimedean copulas. AStA. Advances in Statistical Analysis, 100(3):333–354, July 2016. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s10182-015-0262-8.

Murakami:2016:APM

[397] Hidetoshi Murakami. All-pairs multiple comparisons based on the Cucconi test. AStA. Advances in Statistical Analysis, 100(3):355–368, July 2016. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s10182-016-0268-x.

Moller:2016:SET

[398] Tobias A. Möller, Maria Eduarda Silva, Christian H. Weiß, Manuel G. Scotto, and Isabel Pereira. Self-exciting threshold binomial autoregressive processes. AStA. Advances in Statistical Analysis, 100(4):369–400, October 2016. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s10182-015-0264-6.

Hafner:2016:EAO

[399] Christian M. Hafner and Arie Preminger. The effect of additive outliers on a fractional unit root test. AStA. Advances in Statistical Analysis, 100 (4):401–420, October 2016. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/s10182-015-0265-5.

Ferreira:2016:LBI

[400] Clécio S. Ferreira, Víctor H. Lachos, and Heleno Bolfarine. Likelihood-based inference for multivariate skew scale mixtures of normal distributions. AStA. Advances in Statistical Analysis, 100 (4):421–441, October 2016. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/s10182-016-0266-z.

Gribisch:2016:MWS

[401] Bastian Gribisch. Multivariate Wishart stochastic volatility and changes in regime. AStA. Advances in Statistical Analysis, 100(4):443–473, October 2016. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s10182-016-0269-
                                        9.

Blanco-Fernandez:2016:DGP

[402] Angela Blanco-Fernández and Peter Winker. Data generation processes and statistical management of interval data. AStA. Advances in Statistical Analysis, 100(4):475–494, October 2016. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/s10182-016-0274-z.

Bodnar:2017:HRO

[403] Taras Bodnar and Taras Zabolotskyy. How risky is the optimal portfolio which maximizes the Sharpe ratio? AStA. Advances in Statistical Analysis, 101(1):1–28, January 2017. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s10182-016-0270-3.

Yoshida:2017:NSR

[404] Takuma Yoshida. Nonlinear surface regression with dimension reduction method. AStA. Advances in Statistical Analysis, 101(1):29–50, January 2017. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s10182-016-0271-
2.

Rochani:2017:ECC

[405] Haresh D. Rochani, Robert L. Vogel, Hani M. Samawi, and Daniel F. Linder. Estimates for cell counts and common odds ratio in three-way contingency tables by homogeneous log– linear models with missing data. AStA. Advances in Statistical Analysis, 101 (1):51–65, January 2017. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s10182-016-0275-y.

Garthoff:2017:CCM

[406] Robert Garthoff and Philipp Otto.

Control charts for multivariate spatial autoregressive models. AStA. Advances in Statistical Analysis, 101 (1):67–94, January 2017. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s10182-016-0276-x.

Luo:2017:SEL

[407] Shuanghua Luo, Changlin Mei, and Cheng yi Zhang. Smoothed empirical likelihood for quantile regression models with response data missing at random. AStA. Advances in Statistical Analysis, 101(1):95–116, January 2017. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/s10182-016-0278-8.

Hatefi:2017:ICP

[408] Armin Hatefi and Mohammad Jafari Jozani. Information content of partially rank-ordered set samples. AStA. Advances in Statistical Analysis, 101 (2):117–149, April 2017. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

Andriyana:2017:QRH

[409] Y. Andriyana and I. Gijbels. Quantile regression in heteroscedastic varying coefficient models. AStA. Advances in Statistical Analysis, 101(2):151–176, April 2017. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

Bowater:2017:DSF

[410] Russell J. Bowater. A defence of subjective fiducial inference. AStA. Advances in Statistical Analysis, 101(2):177–197, April 2017. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

Balakrishnan:2017:EPD

[411] N. Balakrishnan, N. Martín, and L. Pardo. Empirical phi-divergence test statistics for the difference of means of two populations. AStA. Advances in Statistical Analysis, 101(2):199–226, April 2017. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

Wang:2017:FCE

[412] Wan-Lun Wang and Tsung-I Lin.

Flexible clustering via extended mixtures of common t-factor analyzers. AStA. Advances in Statistical Analysis, 101(3):227–252, July 2017. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-016-0281-0.

Bodnar:2017:TGM

[413] Taras Bodnar, Stepan Mazur, and Krzysztof Podgórski. A test for the global minimum variance portfolio for small sample and singular covariance. AStA. Advances in Statistical Analysis, 101(3):253–265, July 2017. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-016-0282-z.

He:2017:PMB

[414] Hua He, Wenjuan Wang, and Wan Tang. Prediction model-based kernel density estimation when group membership is subject to missing. AStA. Advances in Statistical Analysis, 101(3): 267–288, July 2017. CODEN ???? ISSN 1863-8171 (print), 1863-818X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10182-016-
0283-y.

Wornowizki:2017:FMA

[415] Max Wornowizki, Roland Fried, and Simos G. Meintanis. Fourier methods for analyzing piecewise constant volatilities. AStA. Advances in Statistical Analysis, 101(3):289–308, July 2017. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-017-0288-1.

Zhang:2017:MVC

[416] Jin Zhang. Minimum volume confidence sets for parameters of normal distributions. AStA. Advances in Statistical Analysis, 101(3):309–320, July 2017. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-017-0290-7.

Ferrari:2017:BCS

[417] Silvia L. P. Ferrari and Giovana Fumes. Box–Cox symmetric distributions and applications to nutritional data. AStA. Advances in Statistical Analysis, 101(3):321–344, July 2017. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-017-0291-6.

Langrock:2017:GEI

[418] Roland Langrock and David L. Borchers. Guest editors’ introduction to the special issue on “ecological statistics”. AStA. Advances in Statistical Analysis, 101(4):345–347, October 2017. CO-DEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10182-017-0307-2;   http://
                                        link.springer.com/content/pdf/10.
                                        1007/s10182-017-0307-2.pdf.

Fewster:2017:SAG

[419] R. M. Fewster. Some applications of genetics in statistical ecology. AStA. Advances in Statistical Analysis, 101(4):349–379, October 2017. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-016-0273-0.

Guillera-Arroita:2017:SOE

[420] Gurutzeta Guillera-Arroita and José J.

Lahoz-Monfort. Species occupancy estimation and imperfect detection: shall surveys continue after the first detection? AStA. Advances in Statistical Analysis, 101(4):381–398, October 2017. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-017-0292-5.

Patterson:2017:SMI

[421] Toby A. Patterson, Alison Parton, Roland Langrock, Paul G. Blackwell, Len Thomas, and Ruth King. Statistical modelling of individual animal movement: an overview of key methods and a discussion of practical challenges. AStA. Advances in Statistical Analysis, 101(4):399–438, October 2017. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-017-0302-7.

Besbeas:2017:VEI

[422] Panagiotis Besbeas and Byron J. T.

Morgan. Variance estimation for integrated population models. AStA. Advances in Statistical Analysis, 101 (4):439–460, October 2017. CO-DEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-017-0304-5.

Buckland:2017:MTT

[423] S. T. Buckland, Y. Yuan, and E. Marcon. Measuring temporal trends in biodiversity. AStA. Advances in Statistical Analysis, 101(4):461–474, October 2017. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL http:

//link.springer.com/article/10.
1007/s10182-017-0308-1;   http://
link.springer.com/content/pdf/10.
1007/s10182-017-0308-1.pdf.

Borchers:2017:DSS

[424] David L. Borchers and Tiago A. Marques. From distance sampling to spatial capture–recapture. AStA. Advances in Statistical Analysis, 101 (4):475–494, October 2017. CO-DEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL http:

//link.springer.com/article/10.
1007/s10182-016-0287-7;   http://
link.springer.com/content/pdf/10.
1007/s10182-016-0287-7.pdf.

Illian:2017:IUS

[425] Janine B. Illian and David F. R. P. Burslem. Improving the usability of spatial point process methodology: an interdisciplinary dialogue between statistics and ecology. AStA. Advances in Statistical Analysis, 101 (4):495–520, October 2017. CO-DEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL http:

//link.springer.com/article/10.
1007/s10182-017-0301-8;   http://
link.springer.com/content/pdf/10.
1007/s10182-017-0301-8.pdf.

Bodnar:2018:AVN

[426] Olha Bodnar and Clemens Elster. Assessment of vague and noninformative priors for Bayesian estimation of the realized random effects in random-effects meta-analysis. AStA. Advances in Statistical Analysis, 102(1):1–20, January 2018. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-016-0279-7.

Giraldo:2018:MTS

[427] Ramón Giraldo, William Caballero, and Jesús Camacho-Tamayo. Mantel test for spatial functional data. AStA. Advances in Statistical Analysis, 102(1):21–39, January 2018. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-016-0280-1.

Wang:2018:SWC

[428] Xuejun Wang, Yi Wu, and Shuhe Hu.

Strong and weak consistency of LS estimators in the EV regression model with negatively superadditive-dependent errors. AStA. Advances in Statistical Analysis, 102(1):41–65, January 2018. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-016-0286-8.

Brachinger:2018:FFH

[429] Hans Wolfgang Brachinger, Michael Beer, and Olivier Schöni. A formal framework for hedonic elementary price indices. AStA. Advances in Statistical Analysis, 102(1):67–93, January 2018. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-017-0293-4.

Amiri:2018:MLV

[430] Leila Amiri, Mojtaba Khazaei, and Mojtaba Ganjali. A mixture latent variable model for modeling mixed data in heterogeneous populations and its applications. AStA. Advances in Statistical Analysis, 102(1):95–115, January 2018. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-017-0294-3.

Dianda:2018:IME

[431] Daniela F. Dianda, Marta B. Quaglino, and José A. Pagura. Impact of measurement errors on the performance and distributional properties of the multivariate capability index NMCpm. AStA. Advances in Statistical Analysis, 102(1):117–143, January 2018. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-017-0295-2.

Putz:2018:PSE

[432] Peter Pütz and Thomas Kneib. A penalized spline estimator for fixed effects panel data models. AStA. Advances in Statistical Analysis, 102(2):145–166, April 2018. CODEN ???? ISSN 1863-8171 (print), 1863-818X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10182-017-
0296-1.

Hahn:2018:CPC

[433] Georg Hahn. Closure properties of classes of multiple testing procedures. AStA. Advances in Statistical Analysis, 102(2):167–178, April 2018. CO-DEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL http:

//link.springer.com/article/10.
1007/s10182-017-0297-0;   http://
link.springer.com/content/pdf/10.
1007/s10182-017-0297-0.pdf.

Ghosh:2018:NCP

[434] Abhik Ghosh and Magne Thoresen.

Non-concave penalization in linear mixed-effect models and regularized selection of fixed effects. AStA. Advances in Statistical Analysis, 102(2):179–210, April 2018. CODEN ???? ISSN 1863-8171 (print), 1863-818X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10182-017-
0298-z.

Nikoloulopoulos:2018:CLB

[435] Aristidis K. Nikoloulopoulos. On composite likelihood in bivariate metaanalysis of diagnostic test accuracy studies. AStA. Advances in Statistical Analysis, 102(2):211–227, April 2018. CO-DEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL http:

//link.springer.com/article/10.
1007/s10182-017-0299-y;   http://
link.springer.com/content/pdf/10.
1007/s10182-017-0299-y.pdf.

Lutkepohl:2018:ESI

[436] Helmut Lütkepohl, Anna Staszewska-Bystrova, and Peter Winker. Estimation of structural impulse responses: short-run versus long-run identifying restrictions. AStA. Advances in Statistical Analysis, 102(2):229–244, April 2018. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-017-0300-9.

Draxler:2018:BCI

[437] Clemens Draxler. Bayesian conditional inference for Rasch models. AStA. Advances in Statistical Analysis, 102(2):245–262, April 2018. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-017-0303-6.

Melo:2018:DBM

[438] Carlos E. Melo, Oscar O. Melo, and Jorge Mateu. A distance-based model for spatial prediction using radial basis functions. AStA. Advances in Statistical Analysis, 102(2):263–288, April 2018. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-017-0305-4.

Chatterjee:2018:EHF

[439] Moumita Chatterjee and Sugata Sen Roy. Estimating the hazard functions of two alternating recurrent events in the presence of covariates. AStA. Advances in Statistical Analysis, 102(2):289–304, April 2018. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-017-0316-1.

Li:2018:FOR

[440] Han Li, Kai Yang, Shishun Zhao, and Dehui Wang. First-order random coefficients integer-valued threshold autoregressive processes. AStA. Advances in Statistical Analysis, 102(3):305–331, July 2018. CODEN ???? ISSN 1863-8171 (print), 1863-818X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10182-017-
                                        0306-3.

Mosammam:2018:PLM

[441] Ali M. Mosammam and Jorge Mateu. A penalized likelihood method for nonseparable space-time generalized additive models. AStA. Advances in Statistical Analysis, 102(3):333–357, July 2018. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-017-0309-0.

Bieniek:2018:UCA

[442] Mariusz Bieniek and Krystyna Macia̧g.

Uniqueness of characterization of absolutely continuous distributions by regressions of generalized order statistics. AStA. Advances in Statistical Analysis, 102(3):359–380, July 2018. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-017-0310-7.

Castilla:2018:MPD

[443] Elena Castilla, Nirian Martín, and Leandro Pardo. Minimum phi-divergence estimators for multinomial logistic regression with complex sample design. AStA. Advances in Statistical Analysis, 102(3):381–411, July 2018. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-017-0311-6.

Zhou:2018:EMM

[444] Sanyu Zhou. An exact method for the multiple comparison of several polynomial regression models with applications in dose-response study. AStA. Advances in Statistical Analysis, 102(3):413–429, July 2018. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-017-0313-4.

Barakat:2018:EUP

[445] H. M. Barakat, E. M. Nigm, O. M. Khaled, and H. A. Alaswed. The estimations under power normalization for the tail index, with comparison. AStA. Advances in Statistical Analysis, 102(3):431–454, July 2018. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-017-0314-3.

Rueda:2018:AEI

[446] María del Mar García Rueda, Pier Francesco Perri, and Beatriz Rodríguez Cobo. Advances in estimation by the item sum technique using auxiliary information in complex surveys. AStA. Advances in Statistical Analysis, 102(3):455–478, July 2018. CODEN ???? ISSN 1863-8171 (print), 1863-818X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10182-017-
                                        0315-2.

Rosa:2018:ODT

[447] Samuel Rosa. Optimal designs for treatment comparisons represented by graphs. AStA. Advances in Statistical Analysis, 102(4):479–503, October 2018. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-017-0312-5.

Chen:2018:TMS

[448] Hong Chen, Maik Döring, and Uwe Jensen. Test for model selection using Cramér–von Mises distance in a fixed design regression setting. AStA. Advances in Statistical Analysis, 102(4):505–535, October 2018. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-017-0317-0.

Choi:2018:MGL

[449] Hosik Choi, Eunjung Song, Seung sik Hwang, and Woojoo Lee. A modified generalized lasso algorithm to detect local spatial clusters for count data. AStA. Advances in Statistical Analysis, 102(4):537–563, October 2018. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-018-0318-7.

Xu:2018:WQR

[450] Hong-Xia Xu, Guo-Liang Fan, Zhen-Long Chen, and Jiang-Feng Wang. Weighted quantile regression and testing for varying-coefficient models with randomly truncated data. AStA. Advances in Statistical Analysis, 102(4):565–588, October 2018. CODEN ???? ISSN 1863-8171 (print), 1863-818X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10182-018-
0319-6.

Fontanella:2018:VLA

[451] Lara Fontanella, Annalina Sarra, Pasquale Valentini, Simone Di Zio, and Sara Fontanella. Varying levels of anomie in Europe: a multilevel analysis based on multidimensional IRT models. AStA. Advances in Statistical Analysis, 102(4):589–610, October 2018. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-018-0320-0.

Dufour:2018:WIP

[452] Jean-Marie Dufour and Joachim Wilde. Weak identification in probit models with endogenous covariates. AStA. Advances in Statistical Analysis, 102 (4):611–631, October 2018. CO-DEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-018-0325-8.

Mayor-Gallego:2019:EFP

[453] J. A. Mayor-Gallego, J. L. Moreno-Rebollo, and M. D. Jiménez-Gamero. Estimation of the finite population distribution function using a global penalized calibration method. AStA. Advances in Statistical Analysis, 103(1):1– 35, March 2019. CODEN ???? ISSN 1863-8171 (print), 1863-818X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10182-018-
                                        0321-z.

Begun:2019:SBS

[454] Alexander Begun and Anatoli Yashin. Study of the bivariate survival data using frailty models based on Lévy processes. AStA. Advances in Statistical Analysis, 103(1):37–67, March 2019. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10182-018-0322-y;   http://
                                        link.springer.com/content/pdf/10.
                                        1007/s10182-018-0322-y.pdf.

Yang:2019:EMA

[455] Yandan Yang, Hon Keung Tony Ng, and Narayanaswamy Balakrishnan. Expectation–maximization algorithm for system-based lifetime data with unknown system structure. AStA. Advances in Statistical Analysis, 103(1):69–98, March 2019. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-018-0323-x.

Moghimbeygi:2019:LMS

[456] Meisam Moghimbeygi and Mousa Golalizadeh. A longitudinal model for shapes through triangulation. AStA. Advances in Statistical Analysis, 103(1):99–121, March 2019. CODEN ???? ISSN 1863-8171 (print), 1863-818X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10182-018-
0324-9.

Yalaz:2019:MPL

[457] Seçil Yalaz. Multivariate partially linear regression in the presence of measurement error. AStA. Advances in Statistical Analysis, 103(1):123–135, March 2019. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-018-0326-7.

Yang:2019:SES

[458] Yiping Yang, Tiejun Tong, and Gaorong Li. SIMEX estimation for singleindex model with covariate measurement error. AStA. Advances in Statistical Analysis, 103(1):137–161, March 2019. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-018-0327-6.

Munoz-Pichardo:2019:IMB

[459] J. M. Muñoz-Pichardo, J. L. Moreno-Rebollo, R. Pino-Mejías, and M. D. Cubiles de la Vega. Influence measures in beta regression models through distance between distributions. AStA. Advances in Statistical Analysis, 103(2):163–185, June 2019. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-018-00332-2.

Sin:2019:OSP

[460] Chor yiu Sin and Shu-Hui Yu. Order selection for possibly infinite-order non-stationary time series. AStA. Advances in Statistical Analysis, 103(2): 187–216, June 2019. CODEN ???? ISSN 1863-8171 (print), 1863-818X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10182-018-
00333-1.

Angelov:2019:MLE

[461] Angel G. Angelov and Magnus Ekström. Maximum likelihood estimation for survey data with informative interval censoring. AStA. Advances in Statistical Analysis, 103(2):217–236, June 2019. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL http:

//link.springer.com/article/10.
1007/s10182-018-00329-x;   http:
//link.springer.com/content/pdf/
10.1007/s10182-018-00329-x.pdf.

Wenger:2019:CMT

[462] Kai Wenger, Christian Leschinski, and Philipp Sibbertsen. Change-in-mean tests in long-memory time series: a review of recent developments. AStA. Advances in Statistical Analysis, 103(2):237–256, June 2019. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-018-0328-5.

Morais:2019:CJC

[463] Manuel Cabral Morais, Wolfgang Schmid, Patrícia Ferreira Ramos, Taras Lazariv, and António Pacheco. Comparison of joint control schemes for multivariate normal i.i.d. output. AStA. Advances in Statistical Analysis, 103(2):257–287, June 2019. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-018-00331-3.  See

correction [464].

Morais:2019:CCJ

[464] Manuel Cabral Morais, Wolfgang Schmid, Patrícia Ferreira Ramos, Taras Lazariv, António Pacheco, and Ivan Semeniuk. Correction to: Comparison of joint control schemes for multivariate normal i.i.d. output. AStA. Advances in Statistical Analysis, 103(2): 289–303, June 2019. CODEN ???? ISSN 1863-8171 (print), 1863-818X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10182-018-
00344-y;   http://link.springer.
com/content/pdf/10.1007/s10182-
018-00344-y.pdf. See [463].

Lazariv:2019:SNS

[465] Taras Lazariv and Wolfgang Schmid.

Surveillance of non-stationary processes. AStA. Advances in Statistical Analysis, 103(3):305–331, September 2019. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-018-00330-4.

Metzner:2019:ALS

[466] Selma Metzner, Gerd Wübbeler, and Clemens Elster. Approximate large-scale Bayesian spatial modeling with application to quantitative magnetic resonance imaging. AStA. Advances in Statistical Analysis, 103(3):333–355, September 2019. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-018-00334-0.

Lee:2019:IPB

[467] Sangyeol Lee, Simos G. Meintanis, and Minyoung Jo. Inferential procedures based on the integrated empirical characteristic function. AStA. Advances in Statistical Analysis, 103(3):357–386, September 2019. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-018-00335-z.

Pardo-Fernandez:2019:MST

[468] Juan Carlos Pardo-Fernández and M. Dolores Jiménez-Gamero. A model specification test for the variance function in nonparametric regression. AStA. Advances in Statistical Analysis, 103(3):387–410, September 2019. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-018-00336-y.

Tekbudak:2019:CTM

[469] Merve Yasemin Tekbudak, Marcela Alfaro-Córdoba, Arnab Maity, and Ana-Maria Staicu. A comparison of testing methods in scalar-on-function regression. AStA. Advances in Statistical Analysis, 103(3):411–436, September 2019. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-018-00337-x.

Kossler:2019:TSV

[470] Wolfgang Kössler and Janine Ott. Twosided variable inspection plans for arbitrary continuous populations with unknown distribution. AStA. Advances in Statistical Analysis, 103(3):437–452, September 2019. CODEN ???? ISSN 1863-8171 (print), 1863-818X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10182-018-
00338-w.

Kulkarni:2019:JQR

[471] Hemant Kulkarni, Jayabrata Biswas, and Kiranmoy Das. A joint quantile regression model for multiple longitudinal outcomes. AStA. Advances in Statistical Analysis, 103(4):453–473, December 2019. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-018-00339-9.

Tang:2019:EVS

[472] Qingguo Tang and Peng Jin. Estimation and variable selection for partial functional linear regression. AStA. Advances in Statistical Analysis, 103(4):475–501, December 2019. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-018-00342-0.

Martinez-Rodriguez:2019:NAT

[473] Ana María Martínez-Rodríguez, Antonio Conde-Sánchez, and María José Olmo-Jiménez. A new approach to truncated regression for count data. AStA. Advances in Statistical Analysis, 103(4):503–526, December 2019. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-018-00345-x.

Raffinetti:2019:MTA

[474] Emanuela Raffinetti and Fabio Aimar. MDCgo takes up the association/ correlation challenge for grouped ordinal data. AStA. Advances in Statistical Analysis, 103(4):527–561, December 2019. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-018-00341-1.

Khan:2019:NTS

[475] M. G. M. Khan and Jacek Wesołowski. Neyman-type sample allocation for domains-efficient estimation in multistage sampling. AStA. Advances in Statistical Analysis, 103(4):563–592, December 2019. CODEN ???? ISSN 1863-8171 (print), 1863-818X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10182-018-
                                        00340-2;   http://link.springer.
                                        com/content/pdf/10.1007/s10182-
                                        018-00340-2.pdf.

Ahmad:2019:UAT

[476] M. Rauf Ahmad. A unified approach to testing mean vectors with large dimensions. AStA. Advances in Statistical Analysis, 103(4):593–618, December 2019. CODEN ???? ISSN 1863-8171 (print), 1863-818X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10182-018-
                                        00343-z;   http://link.springer.
                                        com/content/pdf/10.1007/s10182-
                                        018-00343-z.pdf.

Kauermann:2020:E

[477] Göran Kauermann, Thomas Kneib, and Yarema Okhrin. Editorial.

AStA. Advances in Statistical Analysis, 104(1):1–3, March 2020. CO-DEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL http:

//link.springer.com/article/10.
1007/s10182-020-00361-w;   http:
//link.springer.com/content/pdf/
10.1007/s10182-020-00361-w.pdf.

Grabowski:2020:SAB

[478] Daniel Grabowski, Anna Staszewska-Bystrova, and Peter Winker. Skewnessadjusted bootstrap confidence intervals and confidence bands for impulse response functions. AStA. Advances in Statistical Analysis, 104(1):5–32, March 2020. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-018-00347-9.

Abid:2020:GTR

[479] Rahma Abid, Célestin C. Kokonendji, and Afif Masmoudi. Geometric Tweedie regression models for continuous and semicontinuous data with variation phenomenon. AStA. Advances in Statistical Analysis, 104(1):33–58, March 2020. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-019-00350-8.

Holgersson:2020:RPE

[480] Thomas Holgersson, Peter Karlsson, and Andreas Stephan. A risk perspective of estimating portfolio weights of the global minimum-variance portfolio. AStA. Advances in Statistical Analysis, 104(1):59–80, March 2020. CO-DEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL http:

//link.springer.com/article/10.

                                        1007/s10182-018-00349-7;   http:
                                        //link.springer.com/content/pdf/
                                        10.1007/s10182-018-00349-7.pdf.

Al-Sharadqah:2020:SAC

[481] Ali Al-Sharadqah and Majid Mojirsheibani. A simple approach to construct confidence bands for a regression function with incomplete data. AStA. Advances in Statistical Analysis, 104(1):81–99, March 2020. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-019-00351-7.

Bauer:2020:KNP

[482] Alexander Bauer, Andreas Bender, André Klima, and Helmut Küchenhoff. KOALA: a new paradigm for election coverage. AStA. Advances in Statistical Analysis, 104(1):101–115, March 2020. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-019-00352-6.

Smaga:2020:NRM

[483] Lukasz Smaga. A note on repeated measures analysis for functional data. AStA. Advances in Statistical Analysis, 104(1):117–139, March 2020. CO-DEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10182-018-00348-8;   http:
                                        //link.springer.com/content/pdf/
                                        10.1007/s10182-018-00348-8.pdf.

Cooke:2020:VCR

[484] Roger M. Cooke, Harry Joe, and Bo Chang. Vine copula regression for observational studies. AStA. Advances in Statistical Analysis, 104(2):141–167, June 2020. CODEN ???? ISSN 1863-8171 (print), 1863-818X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10182-019-
00353-5.

Baghfalaki:2020:TMA

[485] Taban Baghfalaki and Mojtaba Ganjali. A transition model for analyzing multivariate longitudinal data using Gaussian copula approach. AStA. Advances in Statistical Analysis, 104(2): 169–223, June 2020. CODEN ???? ISSN 1863-8171 (print), 1863-818X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10182-018-
00346-w.

Klein:2020:DBQ

[486] Nadja Klein and Thomas Kneib. Directional bivariate quantiles: a robust approach based on the cumulative distribution function. AStA. Advances in Statistical Analysis, 104(2):225–260, June 2020. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-019-00355-3.

Biswas:2020:SPQ

[487] Jayabrata Biswas, Pulak Ghosh, and Kiranmoy Das. A semi-parametric quantile regression approach to zeroinflated and incomplete longitudinal outcomes. AStA. Advances in Statistical Analysis, 104(2):261–283, June 2020. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

http://link.springer.com/article/
10.1007/s10182-020-00362-9.

Kokonendji:2020:RVI

[488] Célestin C. Kokonendji, Aboubacar Y. Touré, and Amadou Sawadogo. Relative variation indexes for multivariate continuous distributions on [0, ∞)^{k} and extensions. AStA. Advances in Statistical Analysis, 104(2):285–307, June 2020. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-020-00364-7.

Neumann:2020:NAG

[489] André Neumann and Thorsten Dickhaus. Nonparametric Archimedean generator estimation with implications for multiple testing. AStA. Advances in Statistical Analysis, 104(2):309–323, June 2020. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10182-020-00363-8.

Aleksandrov:2020:TDS

[490] Boris Aleksandrov and Christian H. Weiß. Testing the dispersion structure of count time series using Pearson residuals. AStA. Advances in Statistical Analysis, 104(3):325–361, September 2020. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10182-019-00356-
                                        2.

Cheung:2020:WTE

[491] Ying Lun Cheung and Uwe Hassler. Whittle-type estimation under long memory and nonstationarity. AStA. Advances in Statistical Analysis, 104(3):363–383, September 2020. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

URL https://link.springer.com/
article/10.1007/s10182-019-00358-
0.

Liu:2020:SEH

[492] Mengya Liu, Qi Li, and Fukang Zhu.

Self-excited hysteretic negative binomial autoregression. AStA. Advances in Statistical Analysis, 104(3):385–415, September 2020. CODEN ???? ISSN 1863-8171 (print), 1863-818X (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10182-019-
00360-6.

Axt:2020:VEU

[493] Ieva Axt and Roland Fried. On variance estimation under shifts in the mean. AStA. Advances in Statistical Analysis, 104(3):417–457, September 2020. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

URL https://link.springer.com/
article/10.1007/s10182-020-00366-
5.

Nikolov:2020:MMI

[494] Nikolay I. Nikolov and Eugenia Stoimenova. Mallows’ models for imperfect ranking in ranked set sampling. AStA. Advances in Statistical Analysis, 104(3):459–484, September 2020. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

URL https://link.springer.com/
article/10.1007/s10182-019-00354-
4.

Zamanzade:2020:EEC

[495] Ehsan Zamanzade, M. Mahdizadeh, and Hani M. Samawi. Efficient estimation of cumulative distribution function using moving extreme ranked set sampling with application to reliability. AStA. Advances in Statistical Analysis, 104(3):485–502, September 2020. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

URL https://link.springer.com/
article/10.1007/s10182-020-00368-
3.

Morales:2020:NHP

[496] Fidel Ernesto Castro Morales and Lorena Vicini. A non-homogeneous Poisson process geostatistical model with spatial deformation. AStA. Advances in Statistical Analysis, 104(3):503–527, September 2020. CODEN ???? ISSN 1863-8171 (print), 1863-818X (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10182-020-
00373-6.

Buscemi:2020:MSL

[497] Simona Buscemi and Antonella Plaia. Model selection in linear mixed-effect models. AStA. Advances in Statistical Analysis, 104(4):529–575, December 2020. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

URL https://link.springer.com/
article/10.1007/s10182-019-00359-
z.

Zhou:2020:BSA

[498] Qi Zhou, Yoo-Mi Chin, and Joon Jin Song. Bayesian sensitivity analysis to unmeasured confounding for misclassified data. AStA. Advances in Statistical Analysis, 104(4):577–596, December 2020. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

URL https://link.springer.com/
article/10.1007/s10182-019-00357-
1.

Chen:2020:PEL

[499] Xia Chen and Liyue Mao. Penalized empirical likelihood for partially linear errors-in-variables models. AStA. Advances in Statistical Analysis, 104(4):597–623, December 2020. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

URL https://link.springer.com/
article/10.1007/s10182-020-00365-
6.

Franco-Pereira:2020:BAR

[500] Alba M. Franco-Pereira, Christos T.

Nakas, and M. Carmen Pardo. Biomarker assessment in ROC curve analysis using the length of the curve as an index of diagnostic accuracy: the binormal model framework. AStA. Advances in Statistical Analysis, 104(4):625–647, December 2020. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

URL https://link.springer.com/
article/10.1007/s10182-020-00371-
8.

Shi:2020:BAM

[501] Lei Shi. Bayesian analysis of multivariate ordered probit model with individual heterogeneity. AStA. Advances in Statistical Analysis, 104(4):649–665, December 2020. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

URL https://link.springer.com/
article/10.1007/s10182-020-00369-
2.

Kuroki:2020:VFE

[502] Manabu Kuroki and Hisayoshi Nanmo.

Variance formulas for estimated mean response and predicted response with external intervention based on the backdoor criterion in linear structural equation models. AStA. Advances in Statistical Analysis, 104(4):667–685, December 2020. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

URL https://link.springer.com/
article/10.1007/s10182-020-00372-
7.

Rauber:2020:ITI

[503] Cristine Rauber, Francisco Cribari-Neto, and Fábio M. Bayer. Improved testing inferences for beta regressions with parametric mean link function. AStA. Advances in Statistical Analysis, 104(4):687–717, December 2020. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

URL https://link.springer.com/
article/10.1007/s10182-020-00376-
3.

Abid:2021:PET

[504] Rahma Abid, Célestin C. Kokonendji, and Afif Masmoudi. On Poissonexponential-Tweedie models for ultraoverdispersed count data. AStA. Advances in Statistical Analysis, 105(1):1– 23, March 2021. CODEN ???? ISSN 1863-8171 (print), 1863-818X (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10182-020-
00375-4.

Yuan:2021:ELI

[505] Xiaohui Yuan, Huixian Li, and Tianqing Liu. Empirical likelihood inference for rank regression with doubly truncated data. AStA. Advances in Statistical Analysis, 105(1):25–73, March 2021. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

URL https://link.springer.com/
article/10.1007/s10182-020-00374-
5.

Bohning:2021:COM

[506] Dankmar Böhning and Patarawan Sangnawakij. Count outcome meta-analysis for comparing treatments by fusing mixed data sources: comparing interventions using across report information. AStA. Advances in Statistical Analysis, 105(1):75–85, March 2021. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

URL https://link.springer.com/
article/10.1007/s10182-020-00370-
9.

Zhao:2021:UAR

[507] Yang Zhao and Meng Liu. Unified approach for regression models with nonmonotone missing at random data. AStA. Advances in Statistical Analysis, 105(1):87–101, March 2021. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

URL https://link.springer.com/
article/10.1007/s10182-020-00389-
y.

Yoshida:2021:AME

[508] Takuma Yoshida. Additive models for extremal quantile regression with Pareto-type distributions. AStA. Advances in Statistical Analysis, 105 (1):103–134, March 2021. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL https:

//link.springer.com/article/10.
1007/s10182-020-00386-1.

Perez-Fernandez:2021:VDR

[509] Sonia Pérez-Fernández, Pablo Martínez-Camblor, and Norberto Corral. Visualizing the decision rules behind the ROC curves: understanding the classification process. AStA. Advances in Statistical Analysis, 105(1):135–161, March 2021. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

URL https://link.springer.com/
article/10.1007/s10182-020-00385-
2.

Bahari:2021:GFT

[510] Fayyaz Bahari, Safar Parsi, and Mojtaba Ganjali. Goodness of fit test for general linear model with nonignorable missing on response variable. AStA. Advances in Statistical Analysis, 105(1):163–196, March 2021. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

URL https://link.springer.com/
article/10.1007/s10182-020-00367-
4.

Kazemi:2021:MRE

[511] Iraj Kazemi and Fatemeh Hassanzadeh. Marginalized random-effects models for clustered binomial data through innovative link functions. AStA. Advances in Statistical Analysis, 105(2):197–228, June 2021. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

URL https://link.springer.com/
article/10.1007/s10182-021-00400-
0.

Bailey:2021:EWC

[512] Joseph D. Bailey and Edward A. Codling. Emergence of the wrapped Cauchy distribution in mixed directional data. AStA. Advances in Statistical Analysis, 105(2):229–246, June 2021. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

URL https://link.springer.com/
article/10.1007/s10182-020-00380-

Alvarez:2021:VSP

[513] Agustín Alvarez and Marcela Svarc. A variable selection procedure for depth measures. AStA. Advances in Statistical Analysis, 105(2):247–271, June 2021. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

URL https://link.springer.com/
article/10.1007/s10182-021-00391-
y.

Kulkarni:2021:UIS

[514] H. V. Kulkarni and S. M. Patil.

Uniformly implementable small sample integrated likelihood ratio test for one-way and two-way ANOVA under heteroscedasticity and normality. AStA. Advances in Statistical Analysis, 105(2):273–305, June 2021. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

URL https://link.springer.com/
article/10.1007/s10182-021-00404-
w.

Barbiero:2021:IDV

[515] Alessandro Barbiero. Inducing a desired value of correlation between two point-scale variables: a two-step procedure using copulas. AStA. Advances in Statistical Analysis, 105(2):307–334, June 2021. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

URL https://link.springer.com/
article/10.1007/s10182-021-00405-
9.

Tan:2021:PPT

[516] Wai Hong Tan and Feng Chen.

Predicting the popularity of tweets using internal and external knowledge: an empirical Bayes type approach. AStA. Advances in Statistical Analysis, 105(2):335–352, June 2021. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

URL https://link.springer.com/
article/10.1007/s10182-021-00390-
z.

Mayrink:2021:SEM

[517] Vinícius Diniz Mayrink, Renato Valladares Panaro, and Marcelo Azevedo Costa. Structural equation modeling with time dependence: an application comparing Brazilian energy distributors. AStA. Advances in Statistical Analysis, 105(2):353–383, June 2021. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

URL https://link.springer.com/
article/10.1007/s10182-020-00377-
2.

Shen:2021:EED

[518] Cencheng Shen and Joshua T. Vogelstein. The exact equivalence of distance and kernel methods in hypothesis testing. AStA. Advances in Statistical Analysis, 105(3):385–403, September 2021. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

URL https://link.springer.com/
article/10.1007/s10182-020-00378-
1.

DAmbra:2021:CRO

[519] Antonello D’Ambra, Pietro Amenta, and Eric J. Beh. Confidence regions and other tools for an extension of correspondence analysis based on cumulative frequencies. AStA. Advances in Statistical Analysis, 105(3):405–429, September 2021. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

URL https://link.springer.com/

article/10.1007/s10182-020-00382-
5.

Cavicchioli:2021:OEM

[520] Maddalena Cavicchioli. OLS estimation of Markov switching VAR models: asymptotics and application to energy use. AStA. Advances in Statistical Analysis, 105(3):431–449, September 2021. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

URL https://link.springer.com/
article/10.1007/s10182-020-00383-
4.

Hoseinzadeh:2021:HNR

[521] Akram Hoseinzadeh, Mohsen Maleki, and Zahra Khodadadi. Heteroscedastic nonlinear regression models using asymmetric and heavy tailed two-piece distributions. AStA. Advances in Statistical Analysis, 105(3):451–467, September 2021. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

URL https://link.springer.com/
article/10.1007/s10182-020-00384-
3.

Giudici:2021:CRO

[522] Paolo Giudici and Emanuela Raffinetti.

Cyber risk ordering with rank-based statistical models. AStA. Advances in Statistical Analysis, 105(3):469–484, September 2021. CODEN ???? ISSN 1863-8171 (print), 1863-818X (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10182-020-
00387-0.

Zhang:2021:ODH

[523] Min-Jue Zhang and Rong-Xian Yue. Optimal designs for homoscedastic functional polynomial measurement error models. AStA. Advances in Statistical Analysis, 105(3):485–501, September 2021. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

URL https://link.springer.com/
article/10.1007/s10182-021-00399-
4.

Zaim:2021:BRP

[524] Yasser Al Zaim and Mohammad Reza Faridrohani. Bayesian random projectionbased signal detection for Gaussian scale space random fields. AStA. Advances in Statistical Analysis, 105(3):503–532, September 2021. CODEN ???? ISSN 1863-8171 (print), 1863-818X (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10182-021-
00408-6.

Yang:2021:RCI

[525] Kai Yang, Han Li, and Chenhui Zhang. Random coefficients integervalued threshold autoregressive processes driven by logistic regression. AStA. Advances in Statistical Analysis, 105(4):533–557, December 2021. CO-DEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL https:

//link.springer.com/article/10.
1007/s10182-020-00379-0.

Orozco:2021:NMF

[526] Daniel L. R. Orozco, Lucas O. F. Sales, and André L. S. Pinho. A new mixed first-order integer-valued autoregressive process with Poisson innovations. AStA. Advances in Statistical Analysis, 105(4):559–580, December 2021. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

URL https://link.springer.com/
article/10.1007/s10182-020-00381-
6.

Martinez-Camblor:2021:OCS

[527] Pablo Martínez-Camblor, Sonia Pérez-Fernández, and Susana Díaz-Coto. Optimal classification scores based on multivariate marker transformations. AStA. Advances in Statistical Analysis, 105(4):581–599, December 2021. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

URL https://link.springer.com/
article/10.1007/s10182-020-00388-
z.

Cho:2021:PDM

[528] Seonghun Cho, Shota Katayama, and Young-Geun Choi. Positive-definite modification of a covariance matrix by minimizing the matrix ℓ_{∞} norm with applications to portfolio optimization. AStA. Advances in Statistical Analysis, 105(4):601–627, December 2021. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

URL https://link.springer.com/
article/10.1007/s10182-021-00396-
7.

MacDonald:2021:ERN

[529] Iain L. MacDonald. Is EM really necessary here? Examples where it seems simpler not to use EM. AStA. Advances in Statistical Analysis, 105(4):629–647, December 2021. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

URL https://link.springer.com/
article/10.1007/s10182-021-00392-
x.

Cascos:2021:SMO

[530] Ignacio Cascos. Simultaneous monitoring of origin and scale in left-bounded processes via depth. AStA. Advances in Statistical Analysis, 105(4):649–673, December 2021. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10182-021-00401-
                                        z.

Vencalek:2021:RCN

[531] Ondrej Vencalek and Olusola Samuel Makinde. RR-classifier: a nonparametric classification procedure in multidimensional space based on relative ranks. AStA. Advances in Statistical Analysis, 105(4):675–693, December 2021. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10182-021-00423-
                                        7.

Lu:2022:LIA

[532] Jun Lu, Wen Gan, and Lei Shi. Local influence analysis for GMM estimation. AStA. Advances in Statistical Analysis, 106(1):1–23, March 2022. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10182-021-00398-
                                        5.

Greco:2022:RFM

[533] Luca Greco. Robust fitting of mixtures of GLMs by weighted likelihood. AStA. Advances in Statistical Analysis, 106(1):25–48, March 2022. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10182-021-00402-
                                        y.

Maddanu:2022:HWF

[534] Federico Maddanu. A harmonically weighted filter for cyclical long memory processes. AStA. Advances in Statistical Analysis, 106(1):49–78, March 2022. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

URL https://link.springer.com/
article/10.1007/s10182-021-00394-
9.

Chen:2022:VSC

[535] Bingzhen Chen, Wenjuan Zhai, and Lingchen Kong. Variable selection and collinearity processing for multivariate data via row-elastic-net regularization. AStA. Advances in Statistical Analysis, 106(1):79–96, March 2022. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

URL https://link.springer.com/
article/10.1007/s10182-021-00403-
x.

Bauer:2022:SDN

[536] Verena Bauer, Dietmar Harhoff, and Göran Kauermann. A smooth dynamic network model for patent collaboration data. AStA. Advances in Statistical Analysis, 106(1):97–116, March 2022. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

URL https://link.springer.com/
article/10.1007/s10182-021-00393-
w.

Weinand:2022:MSP

[537] Sebastian Weinand. Measuring spatial price differentials at the basic heading level: a comparison of stochastic index number methods. AStA. Advances in Statistical Analysis, 106(1):117–143, March 2022. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

URL https://link.springer.com/
article/10.1007/s10182-021-00409-
5.

Calcagni:2022:MRN

[538] Antonio Calcagnì and Luigi Lombardi. Modeling random and non-random decision uncertainty in ratings data: a fuzzy beta model. AStA. Advances in Statistical Analysis, 106(1):145–173, March 2022. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10182-021-00407-
                                        7.

Fernandez-Piana:2022:ILD

[539] Lucas Fernandez-Piana and Marcela Svarc. An integrated local depth measure. AStA. Advances in Statistical Analysis, 106(2):175–197, June 2022. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10182-021-00424-
                                        6.

Wand:2022:DEB

[540] M. P. Wand and J. C. F. Yu. Density estimation via Bayesian inference engines. AStA. Advances in Statistical Analysis, 106(2):199–216, June 2022. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10182-021-00422-
                                        8.

Al-Labadi:2022:BNM

[541] Luai Al-Labadi, Forough Fazeli Asl, and Zahra Saberi. A Bayesian nonparametric multi-sample test in any dimension. AStA. Advances in Statistical Analysis, 106(2):217–242, June 2022. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

                                        URL https://link.springer.com/

article/10.1007/s10182-021-00419-
3.

Chen:2022:NCI

[542] Huaping Chen, Qi Li, and Fukang Zhu. A new class of integer-valued GARCH models for time series of bounded counts with extra-binomial variation. AStA. Advances in Statistical Analysis, 106(2):243–270, June 2022. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

URL https://link.springer.com/
article/10.1007/s10182-021-00414-
8.

Zhao:2022:DCM

[543] Yang Zhao. Diagnostic checking of multiple imputation models. AStA. Advances in Statistical Analysis, 106(2): 271–286, June 2022. CODEN ???? ISSN 1863-8171 (print), 1863-818X (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10182-021-
00429-1.

Burgard:2022:SAE

[544] Jan Pablo Burgard, Domingo Morales, and Anna-Lena Wölwer. Small area estimation of socioeconomic indicators for sampled and unsampled domains. AStA. Advances in Statistical Analysis, 106(2):287–314, June 2022. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

URL https://link.springer.com/
article/10.1007/s10182-021-00426-
4.

Tomarchio:2022:MBC

[545] Salvatore D. Tomarchio, Luca Bagnato, and Antonio Punzo. Modelbased clustering via new parsimonious mixtures of heavy-tailed distributions. AStA. Advances in Statistical Analysis, 106(2):315–347, June 2022. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

URL https://link.springer.com/
article/10.1007/s10182-021-00430-
8.

Jahn:2022:RDS

[546] Beate Jahn, Sarah Friedrich, and Tim Friede. On the role of data, statistics and decisions in a pandemic. AStA. Advances in Statistical Analysis, 106(3):349–382, September 2022. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

URL https://link.springer.com/
article/10.1007/s10182-022-00439-

7. See comments [547, 548, 549] and response [551].

Hohle:2022:CRD

[547] Michael Höhle. Comment “On the role of data, statistics and decisions in a pandemic” by Jahn et al. AStA. Advances in Statistical Analysis, 106(3):383–386, September 2022. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

URL https://link.springer.com/
article/10.1007/s10182-022-00451-

x. See [546, 551].

Berger:2022:DER

[548] Ursula Berger, Göran Kauermann, and Helmut Küchenhoff. Discussion on On the role of data, statistics and decisions in a pandemic. AStA. Advances in Statistical Analysis, 106(3):387–390, September 2022. CODEN ???? ISSN 1863-8171 (print), 1863-818X (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10182-022-

00450-y. See [546, 551].

Radermacher:2022:CRD

[549] Walter J. Radermacher. Comment on:

On the role of data, statistics and decisions in a pandemic statistics for climate protection and health-dare (more) progress! AStA. Advances in Statistical Analysis, 106(3):391–397, September 2022. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

URL https://link.springer.com/
article/10.1007/s10182-022-00447-

7. See [546, 551].

Contreras:2022:DLW

[550] Sebastian Contreras, Jonas Dehning, and Viola Priesemann. Describing a landscape we are yet discovering. AStA. Advances in Statistical Analysis, 106(3):399–402, September 2022. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

URL https://link.springer.com/
article/10.1007/s10182-022-00449-
5.

Jahn:2022:ARR

[551] Beate Jahn, Sarah Friedrich, and Tim Friede. Authors’ response: On the role of data, statistics and decisions in a pandemic. AStA. Advances in Statistical Analysis, 106(3):403–405, September 2022. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

URL https://link.springer.com/
article/10.1007/s10182-022-00460-
w.

Nicola:2022:RNF

[552] Giacomo De Nicola, Marc Schneble, and Ursula Berger. Regional now- and forecasting for data reported with delay: toward surveillance of COVID-19 infections. AStA. Advances in Statistical Analysis, 106(3):407–426, September 2022. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

URL https://link.springer.com/
article/10.1007/s10182-021-00433-
5.

Peterson:2022:RSC

[553] Ryan A. Peterson and Joseph E. Cavanaugh. Ranked sparsity: a cogent regularization framework for selecting and estimating feature interactions and polynomials. AStA. Advances in Statistical Analysis, 106(3):427–454, September 2022. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

URL https://link.springer.com/
article/10.1007/s10182-021-00431-
7.

Arisido:2022:ICT

[554] Maeregu W. Arisido, Fulvia Mecatti, and Paola Rebora. Improving the causal treatment effect estimation with propensity scores by the bootstrap. AStA. Advances in Statistical Analysis, 106(3):455–471, September 2022. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

URL https://link.springer.com/
article/10.1007/s10182-021-00427-
3.

Toledo:2022:FMN

[555] Douglas Toledo, Cristiane Akemi Umetsu, and Idemauro Antonio Rodrigues de Lara. Flexible models for non-equidispersed count data: comparative performance of parametric models to deal with underdispersion. AStA. Advances in Statistical Analysis, 106(3):473–497, September 2022. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

URL https://link.springer.com/
article/10.1007/s10182-021-00432-
6.

Caballero:2022:SRT

[556] Yolanda Caballero, Ramón Giraldo, and Jorge Mateu. A spatial randomness test based on the box-counting dimension. AStA. Advances in Statistical Analysis, 106(3):499–524, September 2022. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

URL https://link.springer.com/
article/10.1007/s10182-021-00434-
4.

Magrini:2022:CAA

[557] Alessandro Magrini. Correction to:

Assessment of agricultural sustainability in European Union countries: a group-based multivariate trajectory approach. AStA. Advances in Statistical Analysis, 106(3):525–526, September 2022. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

URL https://link.springer.com/
article/10.1007/s10182-022-00438-

8. See [563].

Gandy:2022:SPE

[558] Axel Gandy, Kaushik Jana, and Almut E. D. Veraart. Scoring predictions at extreme quantiles. AStA. Advances in Statistical Analysis, 106(4):527–544, December 2022. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

URL https://link.springer.com/
article/10.1007/s10182-021-00421-
9.

Davino:2022:QRP

[559] Cristina Davino, Tormod Næs, Rosaria Romano, and Domenico Vistocco. A quantile regression perspective on external preference mapping. AStA. Advances in Statistical Analysis, 106 (4):545–571, December 2022. CO-DEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL https:

//link.springer.com/article/10.
1007/s10182-022-00440-0.

Borroni:2022:SMK

[560] Claudio Giovanni Borroni and Lucio De Capitani. Some measures of kurtosis and their inference on large datasets. AStA. Advances in Statistical Analysis, 106(4):573–607, December 2022. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

URL https://link.springer.com/
article/10.1007/s10182-022-00442-
y.

Hassairi:2022:GRR

[561] Abdelhamid Hassairi, Fatma Ktari, and Raoudha Zine. On the Gaussian representation of the Riesz probability distribution on symmetric matrices. AStA. Advances in Statistical Analysis, 106(4):609–632, December 2022. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

URL https://link.springer.com/
article/10.1007/s10182-022-00436-
w.

DAngelo:2022:LSL

[562] Nicoletta D’Angelo, Marianna Siino, Antonino D’Alessandro, and Giada Adelfio. Local spatial log-Gaussian Cox processes for seismic data. AStA. Advances in Statistical Analysis, 106(4):633–671, December 2022. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

URL https://link.springer.com/
article/10.1007/s10182-022-00444-

w. See correction [599].

Magrini:2022:AAS

[563] Alessandro Magrini. Assessment of agricultural sustainability in European Union countries: a group-based multivariate trajectory approach. AStA. Advances in Statistical Analysis, 106 (4):673–703, December 2022. CO-DEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL https:

//link.springer.com/article/10.
1007/s10182-022-00437-9. See correc-

tion [557].

Zhou:2022:IBE

[564] Xiaoshuang Zhou, Peixin Zhao, and Yujie Gai. Imputation-based empirical likelihood inferences for partially nonlinear quantile regression models with missing responses. AStA. Advances in Statistical Analysis, 106(4):705–722, December 2022. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

URL https://link.springer.com/
article/10.1007/s10182-022-00441-
z.

Groll:2023:ESI

[565] Andreas Groll and Dominik Liebl.

Editorial special issue: Statistics in sports. AStA. Advances in Statistical Analysis, 107(1–2):1–7, March 2023. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

URL https://link.springer.com/
article/10.1007/s10182-022-00453-
9.

Otting:2023:CBM

[566] Marius Ötting, Roland Langrock, and Antonello Maruotti. A copula-based multivariate hidden Markov model for modelling momentum in football. AStA. Advances in Statistical Analysis, 107(1–2):9–27, March 2023. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

URL https://link.springer.com/
article/10.1007/s10182-021-00395-
8.

Dick:2023:ARM

[567] Uwe Dick and Ulf Brefeld. Action rate models for predicting actions in soccer. AStA. Advances in Statistical Analysis, 107(1–2):29–49, March 2023. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

URL https://link.springer.com/
article/10.1007/s10182-022-00435-
x.

Fadel:2023:CMM

[568] Samuel G. Fadel, Sebastian Mair, Ricardo da Silva Torres, and Ulf Brefeld. Contextual movement models based on normalizing flows. AStA. Advances in Statistical Analysis, 107(1–2):51–72, March 2023. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

URL https://link.springer.com/
article/10.1007/s10182-021-00412-
w.

Porro:2023:DSS

[569] Francesco Porro and Mariangela Zenga. Decompositions by sources and by subpopulations of the Pietra index: two applications to professional football teams in Italy. AStA. Advances in Statistical Analysis, 107(1–2):73–100, March 2023. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

URL https://link.springer.com/

article/10.1007/s10182-021-00397-
6.

Zumeta-Olaskoaga:2023:PSI

[570] Lore Zumeta-Olaskoaga, Maximilian Weigert, Jon Larruskain, Eder Bikandi, Igor Setuain, Josean Lekue, Helmut Küchenhoff, and Dae-Jin Lee. Prediction of sports injuries in football: a recurrent time-to-event approach using regularized Cox models. AStA. Advances in Statistical Analysis, 107(1–2):101–126, March 2023. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

URL https://link.springer.com/
article/10.1007/s10182-021-00428-
2.

vanderWurp:2023:ILT

[571] Hendrik van der Wurp and Andreas Groll. Introducing LASSO-type penalisation to generalised joint regression modelling for count data. AStA. Advances in Statistical Analysis, 107 (1–2):127–151, March 2023. CO-DEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL https:

//link.springer.com/article/10.
1007/s10182-021-00425-5.

Ievoli:2023:RPN

[572] Riccardo Ievoli, Aldo Gardini, and Lucio Palazzo. The role of passing network indicators in modeling football outcomes: an application using Bayesian hierarchical models. AStA. Advances in Statistical Analysis, 107(1–2):153–175, March 2023. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

URL https://link.springer.com/
article/10.1007/s10182-021-00411-

Hanck:2023:HBM

[573] Christoph Hanck and Martin C. Arnold. Hierarchical Bayes modelling of penalty conversion rates of Bundesliga players. AStA. Advances in Statistical Analysis, 107(1–2):177–204, March 2023. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10182-021-00420-
                                        w.

Benz:2023:ECS

[574] Luke S. Benz and Michael J. Lopez. Estimating the change in soccer’s home advantage during the Covid-19 pandemic using bivariate Poisson regression. AStA. Advances in Statistical Analysis, 107(1–2):205–232, March 2023. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10182-021-00413-
                                        9.

Gorgi:2023:EFS

[575] P. Gorgi, S. J. Koopman, and R. Lit. Estimation of final standings in football competitions with a premature ending: the case of COVID-19. AStA. Advances in Statistical Analysis, 107(1–2):233–250, March 2023. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10182-021-00415-
                                        7.

Eetvelde:2023:PFS

[576] Hans Van Eetvelde, Lars Magnus Hvattum, and Christophe Ley. The Probabilistic Final Standing Calculator: a fair stochastic tool to handle abruptly stopped football seasons. AStA. Advances in Statistical Analysis, 107(1–2):251–269, March 2023. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

URL https://link.springer.com/
article/10.1007/s10182-021-00416-
6.

Migliorati:2023:IMB

[577] Manlio Migliorati, Marica Manisera, and Paola Zuccolotto. Integration of model-based recursive partitioning with bias reduction estimation: a case study assessing the impact of Oliver’s four factors on the probability of winning a basketball game. AStA. Advances in Statistical Analysis, 107(1–2):271–293, March 2023. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

URL https://link.springer.com/
article/10.1007/s10182-022-00456-
6.

Ekstrom:2023:HBE

[578] Claus Thorn Ekstrøm and Andreas Kryger Jensen. Having a ball: evaluating scoring streaks and game excitement using in-match trend estimation. AStA. Advances in Statistical Analysis, 107(1–2):295–311, March 2023. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

URL https://link.springer.com/
article/10.1007/s10182-022-00452-
w.

Mews:2023:CTS

[579] Sina Mews and Marius Ötting. Continuoustime state-space modelling of the hot hand in basketball. AStA. Advances in Statistical Analysis, 107(1–2):313–326, March 2023. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10182-021-00410-
                                        y.

Reyers:2023:QEN

[580] Matthew Reyers and Tim B. Swartz. Quarterback evaluation in the National Football League using tracking data. AStA. Advances in Statistical Analysis, 107(1–2):327–342, March 2023. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10182-021-00406-
                                        8.

Fabbricatore:2023:CBS

[581] Rosa Fabbricatore, Maria Iannario, Rosaria Romano, and Domenico Vistocco. Component-based structural equation modeling for the assessment of psycho-social aspects and performance of athletes. AStA. Advances in Statistical Analysis, 107(1–2):343–367, March 2023. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10182-021-00417-
                                        5.

Pataky:2023:SIF

[582] Todd Colin Pataky, Konrad Abramowicz, Dominik Liebl, Alessia Pini, Sara Sjöstedt de Luna, and Lina Schelin. Simultaneous inference for functional data in sports biomechanics. AStA. Advances in Statistical Analysis, 107(1–2):369–392, March 2023. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10182-021-00418-

Haupt:2023:E

[583] Harry Haupt, Thomas Kneib, and Yarema Okhrin. Editorial. AStA. Advances in Statistical Analysis, 107 (3):393–396, September 2023. CO-DEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL https:

//link.springer.com/article/10.
1007/s10182-023-00480-0.

Okhrin:2023:DPC

[584] Ostap Okhrin, Michael Rockinger, and Manuel Schmid. Distributional properties of continuous time processes: from CIR to Bates. AStA. Advances in Statistical Analysis, 107(3):397–419, September 2023. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

URL https://link.springer.com/
article/10.1007/s10182-022-00459-
3.

Shang:2023:SBM

[585] Han Lin Shang. Sieve bootstrapping the memory parameter in long-range dependent stationary functional time series. AStA. Advances in Statistical Analysis, 107(3):421–441, September 2023. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

URL https://link.springer.com/
article/10.1007/s10182-022-00463-
7.

Angelov:2023:TSD

[586] Angel G. Angelov and Magnus Ekström. Tests of stochastic dominance with repeated measurements data. AStA. Advances in Statistical Analysis, 107(3):443–467, September 2023. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

URL https://link.springer.com/

                                        article/10.1007/s10182-022-00446-
                                        8.

Xie:2023:GSR

[587] Wanling Xie and Hu Yang. Group sparse recovery via group squareroot elastic net and the iterative multivariate thresholding-based algorithm. AStA. Advances in Statistical Analysis, 107(3):469–507, September 2023. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10182-022-00443-
                                        x.

Saldanha:2023:DUP

[588] Matheus Henrique Junqueira Saldanha and Adriano Kamimura Suzuki. On dealing with the unknown population minimum in parametric inference. AStA. Advances in Statistical Analysis, 107(3):509–535, September 2023. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10182-022-00445-
                                        9.

Cavicchia:2023:HDP

[589] Carlo Cavicchia, Maurizio Vichi, and Giorgia Zaccaria. Hierarchical disjoint principal component analysis. AStA. Advances in Statistical Analysis, 107 (3):537–574, September 2023. CO-DEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10182-022-00458-4.

Metulini:2023:HCM

[590] Rodolfo Metulini, Giorgio Gnecco, Francesco Biancalani, and Massimo Riccaboni. Hierarchical clustering and matrix completion for the reconstruction of world input-output tables. AStA. Advances in Statistical Analysis, 107(3):575–620, September 2023. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

URL https://link.springer.com/
article/10.1007/s10182-022-00448-
6.

Faliva:2023:NPI

[591] Mario Faliva, Consuelo Rubina Nava, and Maria Grazia Zoia. A new price index for multi-period and multilateral comparisons. AStA. Advances in Statistical Analysis, 107(4):621–640, December 2023. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

URL https://link.springer.com/
article/10.1007/s10182-022-00457-
5.

Conti:2023:TPT

[592] Pier Luigi Conti and Livia De Giovanni. Testing for the presence of treatment effect under selection on observables. AStA. Advances in Statistical Analysis, 107(4):641–669, December 2023. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

URL https://link.springer.com/
article/10.1007/s10182-022-00454-
8.

Hammon:2023:MIO

[593] Angelina Hammon. Multiple imputation of ordinal missing not at random data. AStA. Advances in Statistical Analysis, 107(4):671–692, December 2023. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

URL https://link.springer.com/
article/10.1007/s10182-022-00461-
9.

Golosnoy:2023:CCM

[594] Vasyl Golosnoy, Benno Hildebrandt, Steffen Köhler, Wolfgang Schmid, and Miriam Isabel Seifert. Control charts for measurement error models. AStA. Advances in Statistical Analysis, 107(4):693–712, December 2023. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10182-022-00462-
                                        8.

Filippou:2023:EIM

[595] Panagiota Filippou, Giampiero Marra, Rosalba Radice, and David Zimmer. Estimating the impact of medical care usage on work absenteeism by a trivariate probit model with two binary endogenous explanatory variables. AStA. Advances in Statistical Analysis, 107(4):713–731, December 2023. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10182-022-00464-
                                        6.

Crescenzi:2023:HPM

[596] Federico Crescenzi. Hedonic pricing modelling with unstructured predictors: an application to Italian Fashion Industry. AStA. Advances in Statistical Analysis, 107(4):733–753, December 2023. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10182-022-00465-
                                        5.

Michimae:2023:BRR

[597] Hirofumi Michimae and Takeshi Emura. Bayesian ridge regression for survival data based on a vine copula-based prior. AStA. Advances in Statistical Analysis, 107(4):755–784, December 2023. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

URL https://link.springer.com/
article/10.1007/s10182-022-00466-
4.

Osorio:2023:ANN

[598] Felipe Osorio, Manuel Galea, Claudio Henríquez, and Reinaldo Arellano-Valle. Addressing non-normality in multivariate analysis using the tdistribution. AStA. Advances in Statistical Analysis, 107(4):785–813, December 2023. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

URL https://link.springer.com/
article/10.1007/s10182-022-00468-
2.

DAngelo:2023:CLS

[599] Nicoletta D’Angelo, Marianna Siino, Antonino D’Alessandro, and Giada Adelfio. Correction to: Local spatial log-Gaussian Cox processes for seismic data. AStA. Advances in Statistical Analysis, 107(4):815–819, December 2023. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

URL https://link.springer.com/
article/10.1007/s10182-022-00455-

7. See [562].

Bilancia:2024:DCM

[600] Massimo Bilancia, Domenico Vitale, Fabio Manca, Paola Perchinunno, and Luigi Santacroce. A dynamic causal modeling of the second outbreak of COVID-19 in Italy. AStA. Advances in Statistical Analysis, 108(1):1–30, March 2024. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10182-023-00469-
                                        9.

Weissbach:2024:LTH

[601] Rafael Weißbach, Achim Dörre, Dominik Wied, Gabriele Doblhammer, and Anne Fink. Left-truncated health insurance claims data: theoretical review and empirical application. AStA. Advances in Statistical Analysis, 108(1):31–68, March 2024. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10182-023-00471-
                                        1.

Freo:2024:LBV

[602] Marzia Freo and Alessandra Luati. Lasso-based variable selection methods in text regression: the case of short texts. AStA. Advances in Statistical Analysis, 108(1):69–99, March 2024. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10182-023-00472-
                                        0.

Ferreira:2024:CEV

[603] Marta Ferreira. Clustering of extreme values: estimation and application. AStA. Advances in Statistical Analysis, 108(1):101–125, March 2024. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10182-023-00474-
                                        y.

Saraceno:2024:REF

[604] Giovanni Saraceno, Abhik Ghosh, Ayanendranath Basu, and Claudio Agostinelli. Robust estimation of fixed effect parameters and variances of linear mixed models: the minimum density power divergence approach. AStA. Advances in Statistical Analysis, 108(1):127–157, March 2024. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

URL https://link.springer.com/
article/10.1007/s10182-023-00473-
z.

Spagnolo:2024:SSQ

[605] Francesco Schirripa Spagnolo, Riccardo Borgoni, Antonella Carcagnì, Alessandra Michelangeli, and Nicola Salvati. A spatial semiparametric Mquantile regression for hedonic price modelling. AStA. Advances in Statistical Analysis, 108(1):159–183, March 2024. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

URL https://link.springer.com/
article/10.1007/s10182-023-00476-
w.

Halaj:2024:CTG

[606] Katarina Halaj, Bojana Milošević, Marko Obradović, and M. Dolores Jiménez-Gamero. Correlation-type goodness-of-fit tests based on independence characterizations. AStA. Advances in Statistical Analysis, 108 (1):185–207, March 2024. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic). URL https:

//link.springer.com/article/10.
1007/s10182-023-00475-x.

Noia:2024:FCN

[607] Antonio Di Noia, Marzia Marcheselli, Caterina Pisani, and Luca Pratelli. A family of consistent normally distributed tests for Poissonity. AStA. Advances in Statistical Analysis, 108(1):209–223, March 2024. CODEN ???? ISSN 1863-8171 (print), 1863-818X (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10182-023-00478-
                                        8.